### v1

# A Comprehensive Evaluation of Parameter-Efficient Fine-Tuning on Method-Level Code Smell Detection

Abstract.
Code smells, which are suboptimal coding practices that can potentially lead to defects or maintenance issues, can negatively impact the quality of software systems. Most existing code smell detection methods rely on heuristics-based or machine learning (ML) and deep learning (DL)-based techniques. However, these techniques have several drawbacks (e.g., unsatisfactory performance). Large Language Models (LLMs) have garnered significant attention in the software engineering (SE) field, achieving state-of-the-art performance across a wide range of SE tasks. Parameter-Efficient Fine-Tuning (PEFT) methods, which are commonly used to adapt LLMs to specific tasks with fewer parameters and reduced computational resources, have emerged as a promising approach for enhancing the performance of LLMs in various SE tasks. However, LLMs have not yet been explored for code smell detection, and their effectiveness for this task remains unclear. Furthermore, no comprehensive investigation has been conducted on the efficiency of PEFT methods for method-level code smell detection. In this regard, we systematically evaluate the effectiveness of state-of-the-art PEFT methods on both small and large Language Models (LMs) for method-level code smell detection. To begin, we constructed high-quality Java code smell datasets sourced from GitHub. We then fine-tuned four small LMs and six LLMs using various PEFT techniques, including prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$, for code smell detection. Our comparison against full fine-tuning revealed that PEFT methods not only achieve comparable or better effectiveness but also consume less peak GPU memory. Our analysis further explored the performance of small LMs versus LLMs in the context of code smell detection. Surprisingly, we found that LLMs did not outperform small LMs in this specific task, suggesting that smaller models may be more suited for method-level code smell detection. We also investigated the impact of varying hyper-parameter settings and the availability of training data. Our results show that increasing the number of trainable parameters does not necessarily improve the effectiveness of PEFT methods for code smell detection. In contrast, we observed significant performance gains when the training dataset size was increased. Finally, our results show that LMs combined with PEFT methods outperform traditional heuristic-based detectors, advanced DL-based approaches, and GPT-3.5-turbo in method-level code smell detection. Specifically, MCC values improve by 13.24% to 50.59% for Complex Conditional and 11.22% to 54.32% for Complex Method, demonstrating a more effective solution for identifying these code smells. In conclusion, our work provides a comprehensive evaluation of PEFT methods on both small and large LMs for method-level code smell detection. The results highlight the efficiency and effectiveness of PEFT methods in this domain, offering valuable insights for optimizing LMs for method-level code smell detection. We suggest that the choice of PEFT method for code smell detection should be guided by the model used, data availability, and computational resources. Our replication package is available at (replicationPackage, ), enabling others to reproduce and extend our findings.
Code Smell Detection, Large Language Model, Parameter-Efficient Fine-Tuning
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: TOSEMjournalvolume: 0journalnumber: 0article: 0publicationmonth: 0ccs: Software and its engineering Software maintenance tools
1. Introduction
Code smells are indicators of potential issues in code or design (fowler1999refactoring, ), which could adversely affect software quality, particularly in terms of maintainability (palomba2018maintainability, ), code readability (abbes2011empirical, ), and testability (chatzigeorgiou2010investigating, ). Detecting code smells can be used to identify software design issues that lead to potential problems in future development and maintenance (fowler1999refactoring, ; lanza2007object, ; moha2010decor, ). Manual identification of code smells can be time-consuming and costly (prokic2024prescriptive, ), prompting many researchers to explore automated methods for more efficient code smell detection. Existing automated code smell detection methods are typically heuristics-based (marinescu2004detection, ; morales2017on, ; moha2010decor, ). Traditional heuristics-based detectors for code smells rely on a set of manually curated features and thresholds to discriminate between smelly and non-smelly classes (azeem2019machine, ). While it has been shown that these static detectors provide reasonably accurate results, they are proven to have limitations that hinder their practical application (fernandes2016review, ; zhang2011code, ), particularly because their performance is highly dependent on the selection of thresholds. Besides, code smells identified by such heuristics-based methods can be subjectively interpreted by developers, leading to a significant gap between detection results and developers’ agreement (azeem2019machine, ; fontana2016antipattern, ; mantyla2006subjective, ; schumacher2010building, ).
To address these issues, many researchers have incorporated machine learning (ML) and deep learning (DL) algorithms to mitigate the performance instability in code smell detection that arises from the manual threshold-setting process in traditional detectors (fontana2013code, ; arcelli2016comparing, ; sharma2021code, ; liu2021deep, ). By learning to discern between smelly and non-smelly code elements, ML and DL techniques offer a solution to the subjectivity problem inherent in heuristics-based methods (nucci2018detecting, ). However, despite these advancements, ML and DL methods still face significant challenges in achieving considerable performance for code smell detection (pereira2022code, ; alazba2023deep, ).
In recent years, in the field of Natural Language Processing (NLP), increased computational power, advanced machine learning techniques, and access to very large-scale data have led to the emergence of Large Language Models (LLMs). The exceptional performance of LLMs in NLP tasks (e.g., text classification, generation, and summarization) has garnered significant attention and interest from the Software Engineering (SE) domain. Researchers have begun employed LLMs, particularly code-specific LLMs like CodeBERT-Base (feng2020codebert, ) and CodeT5 (wang2021codet5, ), using fine-tuning techniques to tackle various SE tasks, such as code review (lu2023llamareviewer, ) and vulnerability repair (le2024study, ). Among all the fine-tuning methods, Parameter-Efficient Fine-Tuning (PEFT) has emerged as a promising approach, offering an efficient alternative to full fine-tuning, particularly when computational resources are limited. PEFT methods have shown potential in adapting LLMs to various downstream tasks with fewer parameters and reduced resource requirements, thus expanding the practical applicability of LLMs in diverse SE contexts. Previous studies have explored the efficiency of PEFT methods for Language Models (LMs) on automated program repair (li2024comprehensive, ), code generation (liu2023empirical, ), code search (wang2023one, ), and etc. However, to the best of our knowledge, there is currently no comprehensive research exploring the use of LLMs for code smell detection, nor any in-depth studies analyzing the application of PEFT methods in this area.
Moreover, utilizing PEFT methods for code smell detection can facilitate just-in-time code smell detection. Traditional heuristic-based methods for code smell detection are often applied after the software systems have been completed (designitejava, ; pmd, ; marinescu2005iplasma, ). As a result, code smells are typically discovered at later stages of development, requiring refactoring that involves modifying all relevant parts of the system, thereby increasing the cost and complexity of addressing these code smells. In contrast, incorporating LLMs with PEFT techniques allows for real-time detection of code smells, enabling developers to identify and address potential code smells as they write code. For instance, upon completing a function, developers can immediately receive feedback on whether it contains a code smell. This proactive approach contributes to maintaining high code quality throughout the development life cycle, leading to more efficient and effective code maintenance and improvement (ardimento2021temporal, ; ardimento2021transfer, ).
To this end, we conduct a series of experiments on fine-tuning both small and large LMs with PEFT methods on the code smell detection task. In this paper, we define small LMs as those with fewer than 1B parameters, while LMs with 1B parameters or more are considered large LMs. Specifically, we construct high-quality Java code smell datasets comprising two types of widely studied and sophisticated method-level code smells, i.e., Complex Conditional (CC) and Complex Method (CM), as no existing datasets are available. We then apply various PEFT methods, including prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$, to fine-tune four popular small LMs and six cutting-edge LLMs with parameter sizes ranging from 125M to 7B. The effectiveness of these PEFT methods is assessed for code smell detection, and their performance is compared to that of full fine-tuning in terms of both effectiveness and peak GPU memory usage. Additionally, we investigate the performance of PEFT techniques under different hyper-parameter settings and evaluate their effectiveness in low-resource scenarios with varying numbers of training samples on the method-level code smell detection task. Finally, we employ a heuristics-based static analysis tool alongside two state-of-the-art DL-based code smell detection tools to compare the performance of PEFT methods with existing approaches.
The main contributions of this work are as follows:
  • We mine Java repositories from GitHub and employ a traditional heuristic-based detector, along with manual validation, to create high-quality datasets for two types of method-level code smells, i.e., Complex Conditional and Complex Method.
  • We conduct a comprehensive evaluation of multiple PEFT methods on both small and large LMs for code smell detection. Our results indicate that LLMs, when fine-tuned, do not always outperform small LMs. Additionally, PEFT methods achieve better or comparable effectiveness compared to full fine-tuning.
  • We explore the performance of PEFT methods under different hyper-parameter settings and in low-resource scenarios. Our results reveal that adjusting hyper-parameters to increase trainable parameters does not necessarily lead to better effectiveness in detecting code smells. However, the performance of PEFT methods improves with more training samples.
  • We compare a heuristic-based detector and two best-in-class DL-based code smell detection tools with both small and large LMs using PEFT methods on code smell detection. Our results demonstrate that both small and large LMs with PEFT methods consistently outperform the existing heuristic- and DL-based approaches in detecting the method-level code smells, highlighting the superiority of PEFT methods.
  • We provide actionable suggestions and insights for both developers and researchers for using PEFT methods on method-level code smell detection.
The remainder of this paper is organized as follows: Section 2 presents the background information. Section 3 outlines the research questions and the process of dataset construction, while the study results are provided in Section 4. Section 5 discusses the insights for practitioners. Potential threats to validity are addressed in Section 6, with related work reviewed in Section 7. Finally, Section 8 concludes the paper and suggests future directions for research.
2. Background
2.1. The Need of Applying Language Models on Code Smell Detection
Existing heuristic-based code smell detectors rely on pre-defined rules and metrics, while ML- or DL-based approaches depend on algorithms to identify the optimal threshold for code smell classification. However, these tools often fail to consider the inherent context of the code, which is crucial for accurately detecting and understanding code smells (sharma2021code, ). For example, the Java method shown in Figure 1 contains a switch-case statement with seven cases. From the perspective of a Java developer, this method simply maps the numbers 1 through 7 to the days of the week, from Monday to Sunday. It is straight-forward and easy to understand, as all the case statements follow the same structure. However, when scanned by an existing heuristic-based code smell detector, such as DesigniteJava (designitejava, ), the tool flags the method with the Complex Method code smell. This decision is based solely on the method’s cyclomatic complexity, which exceeds a pre-defined threshold, without considering the method’s context.
In contrast, Language Models (LMs), such as ChatGPT (achiam2023gpt, ), have the advantage of considering the inherent context of the code (wang2021codet5, ; roziere2024codellama, ). The contextual understanding allows the LM to accurately determine that the method does not exhibit Complex Method, as it can perform a role similar to a skilled developer assessing the code. The LM can evaluate not only traditional metrics, such as cyclomatic complexity, but also analyze additional factors, including readability, modularity, and overall design. Therefore, compared to traditional heuristic-based static analysis tools and certain ML- or DL-based detectors, LMs have the potential to detect code smells more accurately by considering both structural and contextual factors, rather than relying solely on numerical thresholds or pre-defined rules.
Figure 1. An example of a Java method that is considered to contain Complex Method code smell by the heuristic-based tool
2.2. Full Fine-tuning
Language Models (LMs) which are pre-trained on massive general-purpose corpus via self-supervised method, can be applied to various downstream NLP tasks by adapting their rich possession of learned knowledge through full fine-tuning (qiu2020pre, ). Full fine-tuning involves updating all parameters of an LM, including multi-head attention (MHA), feed-forward neural network (FFN), layer normalization, and so on. However, as LMs scale up, full fine-tuning a large-sized LM, i.e., LLM, can be excessively costly and ultimately make it practically unfeasible (hu2024survey, ; ding2023parameter, ).
2.3. Parameter-Efficient Fine-Tuning (PEFT) Methods
Instead of full fine-tuning, pre-trained LMs could be tailored to downstream tasks by Parameter-Efficient Fine-Tuning (PEFT) methods. PEFT methods involve optimizing only a subset of model parameters or incorporating external modules for new tasks while keeping the majority of parameters frozen. Compared to full fine-tuning, PEFT methods can reduce the substantial computational and data demands of LLMs and achieve comparable or even better results (han2024parameter, ). The right side of Figure 2 illustrates several state-of-the-art PEFT methods employed in this study, including prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$.
Prompt tuning (lester2021power, ) directly appends learnable vectors, known as soft prompts, to the initial input embedding layer. When fine-tuning PLMs with prompt tuning, only the soft prompts are updated, while the remaining model weights and architecture are kept fixed (not updated).
Prefix tuning (li2021prefix, ) introduces adjustable prefix (i.e., a sequence of task-specific vectors) across all Transformer layers. Specifically, these trainable prefix vectors are prepended to the key $K$ and value $V$ of all the MHA mechanisms. Prefix tuning adopts a reparameterization strategy, utilizing a multilayer perceptron (MLP) layer to generate the added vectors instead of optimizing them in a straightforward manner (han2024parameter, ).
LoRA (Low-Rank Adaptation) (hu2022lora, ) injects trainable low-rank decomposition matrices into the query $W_{q}$, key $W_{k}$, and value $W_{v}$ projection matrices of the MHA. For a pre-trained weight matrix $W_{0}\in\{W_{q},W_{k},W_{v}\}$, an additive operation $W_{0}+\Delta W=W_{0}+BA$ is performed, where the rank of matrices $B$ and $A$ is considerably smaller than that of $W_{0}$, consequently effectively reducing the number of trainable parameters while preserving the essential characteristics of the original matrices.
$\mathbf{(IA)^{3}}$ (liu2022few, ) (Infused Adapter by Inhibiting and Amplifying Inner Activation) incorporates three learnable rescaling vectors, $l_{k}$, $l_{v}$, and $l_{ff}$, to rescale the key $K$ and value $V$ in MHA and the inner activations in FFN. The operations conducted within the MHA can be outlined as follows:
[…]
where $\odot$ denotes the element-wise multiplication and $d_{k}$ refers to the dimension of key $K$.
The modifications in the FFN can be defined as:
[…]
where $\sigma$ is the activation function within FFN, and $W_{up}$ and $W_{down}$ indicate the weight matrices of the FFN.
3. Methodology
Figure 2. Left: architecture of our proposed method for method-level code smell detection. Right: a detailed illustration of the PEFT module structure for the four PEFT methods.
3.1. Research Goal
Figure 2 illustrates the architecture of our LLM-based code smell detector using PEFT methods. We frame the method-level code smell detection task as a binary classification problem, where the input is a piece of code. A sequence classification head is added on top of the LLM to handle the classification task. For model fine-tuning, we use four distinct PEFT methods. The model then outputs either 00 or $1$, representing the absence or presence of the specific type of code smell.
3.2. Research Questions
This study aims to evaluate the effectiveness of state-of-the-art PEFT methods (prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$) in small LMs and LLMs for method-level code smell detection. To achieve this goal, we conducted a set of experiments to answer the following Research Questions (RQs):
RQ1: What is the effectiveness of different PEFT methods and full fine-tuning on various small LMs and LLMs? Previous studies have shown remarkable results using PEFT methods to fine-tune small LMs for software engineering tasks like code clone detection and defect detection (liu2023empirical, ; liu2024delving, ; li2024comprehensive, ; wang2023one, ; weyssow2023exploring, ). However, these studies have not paid attention to the effectiveness of PEFT methods for code smell detection. With this RQ, we aim to compare and validate the effectiveness of PEFT methods across multiple LMs of varying parameter sizes (see Section 3.4) specifically for code smell detection.
RQ2: What is the effectiveness of PEFT methods with different hyper-parameters? When fine-tuning LMs for downstream tasks, the configuration of hyper-parameters plays a crucial role in determining the final results by affecting the number of trainable parameters. In prompt tuning and prefix tuning, the number of virtual tokens serves as a critical hyper-parameter (lester2021power, ; li2021prefix, ). For LoRA, the key hyper-parameter is the rank $r$ (hu2022lora, ), while in $(IA)^{3}$, it is the selection of activated modules (liu2022few, ). Due to our limited computational resources and the popularity of LoRA compared to other PEFT methods (gupta2024beyond, ; zhang2024star, ), we decided to employ LoRA to fine-tune a small LM and an LLM with different $r$ values in a range of {8, 32, 64, 640, 1,280, 2,560} to explore the effect of hyper-parameter settings on PEFT methods for code smell detection.
RQ3: What is the effectiveness of PEFT methods in low-resource scenarios? Low-resource scenarios, characterized by constraints on the amount of training data, pose significant challenges to fine-tuning LMs for specific tasks. Due to the scarcity of real-world datasets (zakeri2023systematic, ), these scenarios are common in tasks such as code smell detection. Therefore, it is crucial to investigate how the number of training samples influences the effectiveness of PEFT methods for code smell detection.
RQ4: What is the effectiveness of PEFT methods compared to state-of-the-art code smell detection approaches? Existing code smell detectors include traditional heuristic-based and DL-based methods. However, heuristic-based approaches for code smell detection depend on rules and heuristics derived from software metrics (moha2010decor, ; palomba2015mining, ; palomba2016a, ; fokaefs2007jdeodorant, ), which may suffer from several limitations (e.g. different tools may produce inconsistent results, complicating developers’ refactoring decisions (fontana2012automatic, )). As a result, researchers have been exploring the application of DL techniques as alternatives for code smell detection due to the drawbacks of heuristic-based methods (pecorelli2022on, ), and state-of-the-art DL-based code smell detectors have demonstrated strong effectiveness in identifying code smells (alazba2023deep, ; sharma2021code, ). With this RQ, we aim to assess how LMs utilizing PEFT methods compare against heuristic-based code smell detectors and established DL techniques, including general-purpose LLMs, in terms of their ability to detect and classify code smells accurately.
3.3. Dataset Construction
In this section, we elaborate on the process with the following steps of constructing the high-quality datasets for two types of method-level code smells used to fine-tune both small and large LMs for code smell detection. We created our own datasets because most existing datasets only offer code metrics related to the code smells instead of code snippets, which are not usable for our purposes (sharma2021qscored, ; binh2022mlcodesmell, ; madeyski2023detecting, ). While some code smell datasets provide valid code samples, they do not include Complex Conditional and Complex Method (madeyski2020mlcq, ), two types of method-level code smells. Additionally, the source projects from these existing datasets are quite outdated (fabio2015landfill, ; palomba2018a, ; fontana2016comparing, ; amorim2015experience, ). Considering this, we aim to construct new datasets sourced from the latest trending GitHub projects. Figure 3 illustrates the overview process of dataset construction.
Figure 3. Overview of dataset construction
3.3.1. Selected Code Smells
In this study, we intend to use LLMs to detect two types of method-level code smells, i.e., Complex Conditional and Complex Method. The definitions of the two types of code smells are provided below:
  • Complex Conditional (CC): This code smell occurs when a conditional statement is excessively intricate (sharma2017house, ).
  • Complex Method (CM): This code smell occurs when a method exhibits high cyclomatic complexity (sharma2017house, ).
We chose CC and CM because they have been widely studied in research on DL-based code smell detection (e.g., (sharma2021code, ; ho2023fusion, ; zhang2024code, ; alazba2023deep, )). Another reason for selecting CC and CM is the availability of detection tools, such as DesigniteJava (designitejava, ), which can assist us in collecting candidate code samples that may exhibit these two types of code smells. We excluded very simple smells (e.g., Long Method), which can be easily identified by basic techniques. In contrast, detecting the two selected method-level code smells requires LLMs to spot specific features, pay attention to cohesion and coupling aspects, and analyze the entire method for its structural properties (sharma2021code, ), thereby expanding the ambition and enhance the applicability of this study.
3.3.2. Java Repository Selection
We used GitHub REST API [^1] https://docs.github.com/en/rest?apiVersion=2022-11-28 to search Java repositories from GitHub. From the extensive array of available Java projects hosted on GitHub, we set the following criteria:
  • The repositories should be created between January 1, 2020 and April 30, 2024.
  • The repositories should exclude forked ones to avoid duplicates.
  • The repositories should have a minimum of 1,000 stars and 1,000 lines of code.
These criteria ensured that the selected Java repositories were recent, popular, active, and non-trivial Java repositories. By employing the stringent filtering criteria, we obtained a list of 294 GitHub repositories coded in Java.
3.3.3. Potential Code Smell Detection
DesigniteJava is an open-source tool that supports detection of 18 design and 10 implementation smells, which have been extensively used in many studies as a mature code smell detector (sharma2021qscored, ; shah2023mining, ; barbosa2020revealing, ). In this paper, we had developed DesigniteJava to detect potential CC and CM in the Java repositories we collected in Section 3.3.2. After this step, we got a list of code smells, consisting of the project name, package name, class name, and method name where each detected code smell occurs. It is important to note that the code smell detection results from DesigniteJava may not be entirely accurate, as the detection results can contain both false positives and false negatives. False positives occur when the tool incorrectly flags code as exhibiting CC or CM, while false negatives happen when actual code smells are missed. These inaccuracies stem from the rule-based and heuristic nature of DesigniteJava, which relies on predefined criteria that may not capture all nuances of the code.
3.3.4. Data Preparation
CodeSplitJava (codesplitjava, ) is a utility program designed to partition methods from Java source code into distinct files. We used CodeSplitJava to parse the 294 Java repositories collected in Section 3.3.2 into code files. Each of these code files corresponds to a method in the 294 repositories, and its path reflects the hierarchical structure of the method (i.e., namespaces and packages are converted into folders). By correlating the list of detected code smells collected in Section 3.3.3, with the paths of the code files generated by CodeSplitJava, we identified the method-level code snippets with code smells and those without (i.e., Clean Code).
We removed any duplicate samples within each type of code smell to ensure uniqueness. Additionally, we limited the maximum token length of the samples to 1,024, filtering out any that exceeded this limit. Table 1 provides the number of potential positive samples after each step for the two types of code smells. In total, we collected 4,266 potential positive code samples containing CC and 6,479 containing CM. For both CC and CM, we introduced an equal number of potential negative samples (i.e., code samples without code smells as identified by DesigniteJava) for the subsequent steps. We also verified that all negative samples adhered to the maximum token limit and were free of duplicates. However, the quality of these datasets may be compromised as they could contain false positives and false negatives, requiring further refinement.
Table 1. Number of potential positive code samples after each step of data preparation (#Initial Samples shows the initially collected potential positive samples, #After Deduplication reflects the number of samples after duplicate removal, and #After Token Limit indicates the count of samples that pass the token length filter, i.e., no more than 1,024 tokens)
Code Smell #Initial Samples #After Deduplication #After Token Limit
Complex Conditional 19,754 4,589 4,266
Complex Method 14,979 9,781 6,479
3.3.5. Manual Validation
After the above steps, we obtained an equal number of potential positive and negative samples for both CC ($4,266\times 2=8,532$) and CM ($6,479\times 2=12,958$). To ensure the quality of the datasets for these two types of code smells and to exclude false positives and false negatives, a manual validation was conducted by the first and fourth authors. Initially, a pilot validation was performed at the 95% confidence level, with a margin of error of less than 5%, resulting in the random selection of 368 samples from CC and 374 from CM. The first and the fourth author annotated the selected samples independently, and the pilot annotations were subsequently compared. Cohen’s Kappa coefficient (jacob1960coefficient, ) was used to measure the inter-rater agreement for both CC and CM. The resulting Kappa coefficient for CC was 0.85, while for CM it was 0.81, both exceeding 0.7, indicating a strong level of agreement between the two annotators. Following the pilot validation, the two authors manually annotated the remaining code samples. Throughout the entire validation process, including the pilot phase, any disagreements between the first and fourth authors were resolved through discussions involving the second author, ensuring consistency in the annotations. Overall, we identified 2,708 false positives and 29 false negatives for CC, and 2,376 false positives and 20 false negatives for CM. The numbers of both positive and negative samples for CC and CM after manual validation are presented in Table 2.
Table 2. Number of positive and negative samples after data annotation
Code Smell #Positive #Negative Total
Complex Conditional 1,587 6,945 8,532
Complex Method 4,123 8,835 12,958
3.3.6. Data Split
After manually reviewing the entire dataset, we split the code samples for both CC and CM into training, validation, and test sets following an 8:1:1 ratio, as employed in previous studies (yang2024federated, ; liu2023benchmarking, ). Specifically, the positive and negative samples for each method-level code smell were distributed according to the same 8:1:1 proportion. This approach ensures a consistent distribution of positive and negative cases, providing a solid foundation for model training, validation, and evaluation. The number of code samples in the datasets of CC and CM after this step is presented in Table 3. We used these clean and high-quality datasets as our final datasets for a series of experiments.
Table 3. Number of positive and negative samples after data split
Set Complex Conditional Complex Method
#Positive #Negative #Total #Positive #Negative #Total
Train 1,269 5,557 6,826 3,299 7,067 10,366
Valid 159 694 853 412 884 1,296
Test 159 694 853 412 884 1,296
Total 1,587 6,945 8,532 4,123 8,835 12,958
3.4. Base Model Selection
In this work, we aim to explore the effectiveness of PEFT methods on method-level code smell detection task. To ensure a thorough analysis, we have chosen to use both small LMs and LLMs as our base models. Note that we deem models with 1B or more parameters as LLMs, while referring to those with fewer parameters as small language models.
First, the selected LMs should be open-source as PEFT methods require access to their parameters. Second, the models should support classification tasks as we frame code smell detection as a binary classification problem. Third, the base models used in this study must be able to be fine-tuned and tested on a single 48GB GPU (our maximum computational resource) without encountering memory overflow. Fourth, the selected models should either be popular small LMs or recent advanced LLMs to ensure they represent state-of-the-art capabilities.
Based on the above criteria, we employed 4 small LMs (i.e., CodeBERT (feng2020codebert, ), GraphCodeBERT (guo2021graphcodebert, ), UnixCoder (guo2022unixcoder, ), and CodeT5 (wang2021codet5, )) and 6 LLMs (DeepSeek-Coder-1.3B (guo2024deepseekcoder, ), DeepSeek-Coder-6.7B (guo2024deepseekcoder, ), StarCoderBase-1B (li2023starcoder, ), StarCoderBase-3B (li2023starcoder, ), StarCoderBase-7B (li2023starcoder, ), and CodeLlama-7B (roziere2024codellama, )) for our experiments. These models feature a range of transformer architectures, including encoder-only, encoder-decoder, and decoder-only modes. All the models we selected have demonstrated encouraging results across various code understanding tasks like code clone detection, defect detection, etc. (liu2024vul, ; karmakar2021what, ; lu2021codexglue, ; li2022automating, ; tang2023csgvd, ; zeng2022an, ; chen2023diversevul, ; wang2023one, ; fu2022vulrepair, ; nguyen2024empirical, ; su2024distilled, ). Table 4 presents the details of the 10 base models, including their parameter sizes and architecture.
Table 4. Selected LMs and their corresponding parameter size and architecture
Model #Params Architecture
Small LMs CodeBERT 125M Encoder-only
GraphCodeBERT 125M
UnixCoder 126M
CodeT5 220M Encoder-decoder
LLMs DeepSeek-Coder-1.3B 1.3B Decoder-only
DeepSeek-Coder-6.7B 6.7B
StarCoderBase-1B 1.1B
StarCoderBase-3B 3B
StarCoderBase-7B 7B
CodeLlama-7B 7B
3.5. Metrics
We use metrics that align with previous research (sharma2021code, ; ho2023fusion, ; pecorelli2019comparing, ) to maintain consistency in evaluating the effectiveness of PEFT methods for code smell detection on LLMs. Specifically, we utilize precision, recall, F1-score, and MCC (Matthews Correlation Coefficient) for assessing the performance of detecting CC and CM. We include MCC because, unlike other metrics like F1-score, accounts for true negative instances (yao2020assessing, ). These metrics are calculated using the following formulas:
(3) $\text{Precision}=\frac{TP}{TP+FP}$
(4) $\text{Recall}=\frac{TP}{TP+FN}$
[…]
(6) $\text{MCC}=\frac{TP\times TN-FP\times FN}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}}$
where $TP$, $TN$, $FP$, $FN$ represent true positives, true negatives, false positives, and false negatives, respectively.
3.6. Implementation Details
The code for our experiments is implemented by the PEFT library [^2] https://huggingface.co/docs/peft/index from Hugging Face. After reviewing prior studies (liu2023empirical, ; liu2024delving, ; li2024comprehensive, ; wang2023one, ; weyssow2023exploring, ) and observing the results of pilot experiments, we set the learning rate to $3e^{-4}$ for all PEFT methods and $1e^{-5}$ for full fine-tuning the small LMs. Throughout all the fine-tuning experiments, we utilized the AdamW optimizer (loshchilov2017decoupled, ) and fixed the Transformer input window size at 1,024. All the models were trained for 10 epochs, and we selected the epochs at which models exhibit minimum loss on the validation sets for inference. We executed all models on a single Linux GPU server equipped with a 48G NVIDIA RTX 6000 Ada Generation GPU. Additionally, we have released our datasets and source code online for replication purpose (replicationPackage, ) where more details, including hyper-parameter values, can be found.
4. Results
In this section, we first illustrate the experiment setup for each RQ formulated in Section 3, and we then present the results and findings for the four RQs.
4.1. RQ1: What is the effectiveness of different PEFT methods and full fine-tuning on various small LMs and LLMs?
Since our objective is to demonstrate the effectiveness of PEFT methods on both small and large LMs for code smell detection, we conducted a series of experiments involving four PEFT methods (i.e., prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$) and a group of models with different parameter sizes. Each PEFT method was employed to fine-tune the 9 base models (see Section 3.4). Note that we only applied LoRA and $(IA)^{3}$ to Code-T5 as the current PEFT library does not support fine-tuning encoder-decoder models with prompt tuning and prefix tuning. We full fine-tuned all the small models due to resource constraints, as the computational demands of tuning all parameters in LLMs would surpass the available 48GB of GPU memory. Additionally, we reported and analyzed the computing resource consumption of these fine-tuning methods, and the results are shown in Figures 4, 5, and 6. Table 5 and Table 6 present the comprehensive comparison of small LMs and LLMs tuning with various PEFT techniques for detecting CC and CM, respectively.
Table 5. Results of small LMs and LLMs with different PEFT methods and full fine-tuning on detecting CC (bold: highest values per model, underline: highest values among all models)
Model Method Precision Recall F1 MCC
Small LMs CodeBERT Full FT 90.68% 50.00% 44.86% 0.00%
Prompt Tuning 72.81% 69.55% 70.92% 42.23%
Prefix Tuning 75.82% 73.23% 74.38% 48.98%
LoRA 80.57% 69.03% 72.49% 48.24%
$(IA)^{3}$ 81.22% 68.86% 72.45% 48.53%
GraphCodeBERT Full FT 70.45% 75.65% 72.25% 45.80%
Prompt Tuning 73.88% 71.85% 72.77% 45.69%
Prefix Tuning 81.05% 73.00% 75.95% 53.44%
LoRA 84.72% 69.67% 73.82% 52.27%
$(IA)^{3}$ 83.33% 74.30% 77.58% 56.91%
UnixCoder Full FT 67.67% 60.01% 61.69% 26.60%
Prompt Tuning 73.93% 70.64% 72.03% 44.44%
Prefix Tuning 78.73% 73.53% 75.64% 52.00%
LoRA 82.12% 69.00% 72.74% 49.40%
$(IA)^{3}$ 74.33% 69.26% 71.21% 43.29%
CodeT5 Full FT 76.94% 78.09% 77.49% 55.02%
LoRA 63.08% 67.65% 46.82% 30.38%
$(IA)^{3}$ 62.42% 67.82% 63.11% 29.76%
LLMs DeepSeek-Coder-1.3B Prompt Tuning 75.25% 70.27% 72.23% 45.25%
Prefix Tuning 75.07% 75.68% 75.37% 50.75%
LoRA 78.35% 70.51% 73.28% 48.22%
$(IA)^{3}$ 75.74% 70.66% 72.65% 46.11%
DeepSeek-Coder-6.7B Prompt Tuning 74.96% 68.91% 71.13% 43.46%
Prefix Tuning 79.36% 70.72% 73.70% 49.33%
LoRA 75.55% 71.63% 73.26% 47.02%
$(IA)^{3}$ 75.79% 67.29% 69.99% 42.23%
StarCoderBase-1B Prompt Tuning 78.44% 71.06% 73.74% 48.95%
Prefix Tuning 76.84% 70.95% 73.20% 47.42%
LoRA 83.53% 54.82% 54.21% 25.41%
$(IA)^{3}$ 80.02% 70.31% 73.51% 49.39%
StarCoderBase-3B Prompt Tuning 73.67% 71.78% 72.64% 45.41%
Prefix Tuning 74.18% 70.47% 72.01% 44.49%
LoRA 69.39% 54.89% 54.72% 19.49%
$(IA)^{3}$ 79.58% 69.12% 72.39% 47.57%
StarCoderBase-7B Prompt Tuning 73.31% 68.97% 70.69% 42.05%
Prefix Tuning 74.73% 72.62% 73.58% 47.30%
LoRA 69.67% 65.46% 67.03% 34.88%
$(IA)^{3}$ 75.27% 76.48% 75.84% 51.73%
CodeLlama-7B Prompt Tuning 74.10% 71.92% 72.90% 45.97%
Prefix Tuning 71.88% 71.13% 71.49% 43.00%
LoRA 76.57% 73.20% 74.66% 49.66%
$(IA)^{3}$ 71.26% 71.57% 71.41% 42.83%
Table 6. Results of small LMs and LLMs with different PEFT methods and full fine-tuning on detecting CM (bold: highest values per model, underline: highest values among all models)
Model Method Precision Recall F1 MCC
Small LMs CodeBERT Full FT 78.55% 53.84% 48.59% 20.95%
Prompt Tuning 84.14% 88.18% 85.20% 72.21%
Prefix Tuning 85.98% 88.42% 86.95% 74.36%
LoRA 85.07% 86.07% 85.53% 71.13%
$(IA)^{3}$ 86.17% 87.32% 86.70% 73.48%
GraphCodeBERT Full FT 78.34% 82.63% 77.01% 60.81%
Prompt Tuning 84.75% 87.83% 85.84% 72.52%
Prefix Tuning 85.54% 86.79% 86.11% 72.33%
LoRA 84.92% 86.97% 85.77% 71.86%
$(IA)^{3}$ 84.15% 86.40% 85.05% 70.51%
UnixCoder Full FT 76.24% 74.36% 75.13% 50.57%
Prompt Tuning 84.64% 87.26% 85.64% 71.85%
Prefix Tuning 86.89% 88.79% 87.70% 75.66%
LoRA 86.99% 90.66% 88.20% 77.56%
$(IA)^{3}$ 86.26% 87.82% 86.95% 74.07%
CodeT5 Full FT 85.66% 89.44% 86.83% 75.00%
LoRA 78.08% 52.26% 45.44% 15.92%
$(IA)^{3}$ 42.00% 43.21% 42.36% -14.73%
LLMs DeepSeek-Coder-1.3B Prompt Tuning 85.83% 88.05% 86.74% 73.85%
Prefix Tuning 86.40% 89.81% 87.57% 76.13%
LoRA 85.14% 87.79% 86.15% 72.88%
$(IA)^{3}$ 85.23% 87.59% 86.17% 72.78%
DeepSeek-Coder-6.7B Prompt Tuning 86.22% 88.92% 87.26% 75.08%
Prefix Tuning 85.73% 89.96% 86.86% 75.57%
LoRA 84.98% 87.48% 85.95% 72.42%
$(IA)^{3}$ 85.58% 83.14% 84.18% 68.68%
StarCoderBase-1B Prompt Tuning 85.68% 87.74% 86.54% 73.39%
Prefix Tuning 84.82% 87.05% 85.72% 71.83%
LoRA 84.81% 87.57% 85.84% 72.33%
$(IA)^{3}$ 86.30% 89.30% 87.40% 75.53%
StarCoderBase-3B Prompt Tuning 84.53% 86.24% 85.26% 70.74%
Prefix Tuning 85.10% 86.38% 85.68% 71.48%
LoRA 84.83% 89.44% 85.77% 74.13%
$(IA)^{3}$ 86.46% 88.38% 87.28% 74.81%
StarCoderBase-7B Prompt Tuning 83.64% 88.07% 84.53% 71.57%
Prefix Tuning 85.35% 87.26% 86.16% 72.59%
LoRA 85.64% 88.45% 86.70% 74.04%
$(IA)^{3}$ 85.72% 88.38% 86.75% 74.05%
CodeLlama-7B Prompt Tuning 74.05% 87.68% 86.00% 72.60%
Prefix Tuning 84.10% 84.36% 84.23% 68.46%
LoRA 85.39% 88.29% 86.46% 73.62%
$(IA)^{3}$ 84.07% 88.40% 85.05% 72.34%
4.1.1. Small LMs
Generally, GraphCodeBERT is the most effective in identifying CC, UnixCoder excels at detecting CM, and CodeT5 with PEFT methods performs the worst in classifying both types of method-level code smells. For CC and CM, CodeBERT demonstrates comparable results to GraphCodeBERT. For example, the F1-score and MCC for CodeBERT with prompt tuning in detecting CM are 85.20% and 72.21%, respectively, while the F1-score and MCC for GraphCodeBERT with prompt tuning for CM are 85.84% and 72.52%. These results are consistent with previous studies (arshad2022codebert, ; rabbani2022comparative, ; lu2021codexglue, ) that CodeBERT and GraphCodeBERT are powerful in code semantics learning, showing outstanding performance in Java code clone detection. GraphCodeBERT is built upon CodeBERT and additionally considers the inherent structure of code (i.e., data flow) in the pre-training stage (guo2021graphcodebert, ). Therefore, it is reasonable that GraphCodeBERT manifests a bit stronger capability than CodeBERT in code smell detection. UnixCoder also shows similar performance to GraphCodeBERT for CC, while it stands out as the best small model for CM, achieving an F1-score of 88.20% and an MCC of 77.56% with LoRA. For all the small LMs, CodeT5 with PEFT methods significantly lags behind the other three models with PEFT methods for both CC and CM, with MCC values well below 50%, and even a negative value when tuned with $(IA)^{3}$ for CM. This suggests that CodeT5 with PEFT methods struggles to perform effectively in code smell detection compared to the other models, highlighting the limitations of its PEFT adaptation. In contrast, fully fine-tuned CodeT5 outperforms all other fully fine-tuned small models in detecting both CC and CM. The stark performance gap between CodeT5 with full fine-tuning and PEFT methods could stem from CodeT5’s architecture, which may depend more on full fine-tuning for optimal adaptation to the method-level code smell detection task. Compared to full fine-tuning, PEFT methods like LoRA and $(IA)^{3}$ might not provide sufficient adjustments to effectively adapt CodeT5’s encoder-decoder architecture.
4.1.2. LLMs
DeepSeek-Coder-1.3B stands out as the top-performing LLM, consistently achieving the highest or on-par F1-scores and MCC values across all fine-tuning methods when compared to other LLMs. While StarCoderBase-7B with $(IA)^{3}$ gains higher F1-score and MCC value than DeepSeek-Coder-1.3B in detecting CC, it significantly underperforms when fine-tuned with LoRA. Specifically, the F1-score and MCC for DeepSeek-Coder-1.3B with LoRA are 73.28% and 48.22%, respectively, compared to 67.03% and 34.88% for StarCoderBase-7B. Notably, when fine-tuned with PEFT techniques, models with larger parameter sizes within the same family do not surpass, and in some cases, even slightly underperform their smaller counterparts. For example, when detecting CC, DeepSeekCoder-1.3B gains higher F1-scores and MCC values than DeepSeek-Coder-6.7B across the four PEFT methods, and StarCoderBase-1B with prompt tuning and prefix tuning also outperforms StarCoderBase-7B. This observation aligns with Fan et al.’s research (fan2024exploring, ), which shows that larger models do not always have better performance on code-related tasks.
Figure 4 illustrates the peak GPU memory consumption for the three largest models — DeepSeek-Coder-6.7B, StarCoderBase-7B, and CodeLlama-7B — using different PEFT techniques for code smell detection. It is evident that, for both CodeLlama-7B and DeepSeek-Coder-7B, $(IA)^{3}$ requires the highest computational resources, with LoRA ranking second. This observation is partly consistent with Li et al.’s work (li2024comprehensive, ) that $(IA)^{3}$ tends to exhibit the highest memory usage. However, Li et al. (li2024comprehensive, ) also found that LoRA consumes less memory than prefix tuning, which is not the case in our results. Unlike DeepSeek-Coder-6.7B and CodeLlama-7B, StarCoderBase-7B shows that LoRA demands more GPU memory than the other PEFT techniques.
Figure 4. Peak GPU memory of DeepSeek-Coder-6.7B, StarCoderBase-7B, and CodeLlama-7B with different PEFT methods on detecting CC and CM
4.1.3. Small LMs vs. LLMs
From Table 5 and Table 6, it can be observed that the best-performing models for detecting CC and CM are GraphCodeBERT with $(IA)^{3}$ and UnixCoder with LoRA, both of which are small LMs. More specifically, small LMs with PEFT techniques, with the exception of CodeT5, generally outperform or achieve comparable performance to LLMs. While Weyssow et al.’s work (weyssow2023exploring, ) highlights that LLMs significantly overtake small LMs in code generation tasks, our findings suggest that, for the binary classification task which identifies method-level code smells, small LMs with fewer number of trainable parameters excel beyond LLMs. This discrepancy might arise because the large capacity of LLMs introduces unnecessary complexity, which can obscure the essential patterns needed for the code smell detection task. In contrast, small LMs appear to possess an optimal capacity for capturing these crucial patterns without being encumbered by excessive complexity (tay2023scaling, ). Another possible explanation for the discrepancy is that small LMs, being encoder-only models, are better suited to classification tasks like code smell detection. Encoder-only models are typically more efficient at tasks that require pattern recognition, such as binary classification, while LLMs are often optimized for generative tasks. This indicates that the advantages of LLMs observed in code generation tasks do not necessarily translate to all types of software engineering tasks, and larger models are not always superior when fine-tuned for different downstream applications.
4.1.4. Best PEFT Method
Overall, for both small LMs and LLMs in the DeepSeek-Coder family, prefix tuning emerges as the most effective technique for method-level code smell detection among the four PEFT methods investigated in this study. In contrast, for the StarCoderBase family, $(IA)^{3}$ tends to be the most effective technique. Meanwhile, for CodeLlama-7B, LoRA proves to be the best PEFT method. For small models, except for CodeT5, as well as LLMs like DeepSeek-Coder-1.3B and DeepSeek-Coder-6.7B, prefix tuning consistently yields the highest F1-scores and MCC values when detecting CC and CM, demonstrating its generally strong effectiveness in fine-tuning these models for code smell detection. In the StarCoderBase family, LLMs fine-tuned with $(IA)^{3}$ outperform those fine-tuned with other PEFT techniques. For example, when detecting CM, StarCoderBase-1B, StarCoderBase-3B, and StarCoderBase-7B all achieve their highest F1-scores and MCC values using $(IA)^{3}$. However, this does not apply to CodeLlama-7B, where using $(IA)^{3}$ results in the worst performance for detecting CC and the second-worst performance for detecting CM. On the other hand, CodeLlama-7B always achieves its best results on this task with LoRA. These findings contrast with those of previous studies (li2024comprehensive, ; weyssow2023exploring, ), which suggested that a single PEFT method could be the best across all models for a specific downstream task. For method-level code smell detection, the optimal PEFT method varies depending on the model. Notably, prompt tuning is the least effective PEFT method for adapting LMs to detect CC and CM, as most models tuned with prefix tuning show worse performance compared to those tuned with other PEFT techniques.
Figure 5. Peak GPU memory of small LMs with different PEFT methods and full fine-tuning on detecting CC
Figure 6. Peak GPU memory of small LMs with different PEFT methods and full fine-tuning on detecting CM
4.1.5. PEFT methods vs. Full Fine-tuning
According to Table 5 and Table 6, we can find that all the four PEFT methods achieve better or at least comparable effectiveness than full fine-tuning on most small models for method-level code smell detection. For small LMs, excluding CodeT5, the F1-scores and MCC values of prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$ all show varying degrees of improvement over full fine-tuning. Our results align with Liu et al.’s prior study (liu2023empirical, ), confirming that PEFT methods can match or surpass the performance of standard full fine-tuning for code understanding tasks while adjusting fewer number of trainable parameters. One possible reason for this could be that full fine-tuning adjusts all of the model’s parameters, which may lead to overfitting, especially when the training data is limited. In contrast, PEFT methods, by focusing on specific parts of the model, allow for more efficient learning without the risk of overfitting to the task at hand.
PEFT methods focus on enhancing the effectiveness of LMs to levels similar to full fine-tuning, while aiming to minimize the use of computational resources, such as peak GPU memory consumption. Figure 5 and Figure 6 depict the peak GPU memory usage and corresponding MCC values of fine-tuning the four small models. It can be observed that prefix tuning and prompt tuning require significantly less GPU memory than other fine-tuning methods, yet they can still achieve considerable effectiveness. Notably, prefix tuning attains the highest or second-highest MCC value while using minimal GPU memory, proving that prefix tuning is the most efficient PEFT method for small models on code smell detection. When juxtaposed with full fine-tuning, all of the four PEFT methods lead to a substantial reduction in peak GPU memory usage while achieving better or similar results. This further validates the efficiency of PEFT methods for code smell detection relative to full fine-tuning.
Answer to RQ1: Small LMs, with the exception of CodeT5, generally achieve better effectiveness than LLMs when using PEFT methods for code smell detection. The most effective PEFT method varies across different models. All PEFT methods outperform or match the performance of full fine-tuning while requiring less peak GPU memory, highlighting the efficiency of PEFT techniques compared to full fine-tuning for detecting CC and CM.
4.2. RQ2: What is the effectiveness of PEFT methods with different hyper-parameters?
Given the restrictions of our computational resources and popularity of LoRA, we focused on a single PEFT method, LoRA, to investigate how small LMs and LLMs perform under different hyper-parameter settings. Specifically, we used LoRA to tune a small LM (GraphCodeBERT) and an LLM (DeepSeek-Coder-1.3B), adapting the value of LoRA rank $r$ in the range of {16, 32, 64, 640, 1,280, 2,560}. At the same time, we kept other hyper-parameters of LoRA constant, with the LoRA alpha set to 16 and the LoRA dropout set to 0.1. The experiment results are presented in Figure 7.
Figure 7. F1-scores and MCC of GraphCodeBERT and DeepSeek-Coder-1.3B using LoRA on detecting CC and CM with different LoRA rank values
The results in Figure 7 show that both CodeBERT and DeepSeek-Coder-1.3B achieve their best performance in detecting CC when fine-tuned with LoRA at an $r$ value of 32. For CM, GraphCodeBERT performs best at a LoRA $r$ value of 32, while DeepSeek-Coder-1.3B reaches its peak performance at $r=64$. As the $r$ value increases from 16 to 2,560, the effectiveness of code smell detection initially increases, then decreases, and finally rises again at $r=2,560$, but it does not exceed the level at $r=16$.
This pattern can be explained by the increase in trainable parameters as $r$ rises, which initially enhances the performance by allowing the models to learn more complex data patterns. However, beyond a certain threshold, the challenge of tuning these additional parameters effectively, especially with relatively limited amount of data, leads to diminishing returns. Without pre-training for these extra parameters, the models are prone to overfitting, where they may learn noise or irrelevant patterns, limiting their generalization capability. Consequently, as the number of trainable parameters grows beyond what the available training data can support, the models’ performance may plateau or even decline. At $r=2,560$, the models appear to utilize the expanded parameter space somewhat more effectively, leading to a slight performance recovery. However, this does not surpass the performance at $r=16$, suggesting that the added parameters introduce more complexity than the data can justify. These results highlight that increasing the number of trainable parameters does not always lead to linear performance gains.
Answer to RQ2: Increasing the number of trainable parameters by adjusting hyper-parameters does not necessarily ensure effectiveness enhancement for code smell detection task. In fact, the effectiveness often declines as the number of trainable parameters continues to increase.
4.3. RQ3: What is the effectiveness of PEFT methods in low-resource scenarios?
To simulate the limited availability of data in low-resource scenarios, following previous works (wang2023one, ; liu2023empirical, ), we randomly selected 100, 200, 500, and 1,000 samples from our datasets while maintaining the proportions of the original data distribution. In this RQ, we employed the four PEFT methods and full fine-tuning on a small LM (GraphCodeBERT) and an LLM (DeepSeek-Coder-1.3B) to conduct the experiments. The experiment results are shown in Table 7.
Table 7. Results of GrahphCodeBERT and DeepSeek-Coder-1.3B with different PEFT methods and full fine-tuning on detecting CC and CM in low-resource scenarios (bold: highest values per scenario for a model, underline: highest values among all scenarios for a model)
Model #Training Samples Method CC CM
F1 MCC F1 MCC
GraphCodeBERT 100 Full FT 44.86% 0.00% 83.87% 69.52%
Prompt Tuning 47.18% 6.84% 80.25% 61.85%
Prefix Tuning 63.94% 32.85% 76.33% 52.67%
LoRA 57.34% 25.28% 82.44% 64.96%
$(IA)^{3}$ 69.63% 39.47% 81.32% 65.84%
200 Full FT 52.48% 21.18% 86.26% 73.93%
Prompt Tuning 67.05% 34.12% 83.26% 67.26%
Prefix Tuning 72.47% 46.00% 81.15% 62.49%
LoRA 66.61% 41.44% 85.20% 71.60%
$(IA)^{3}$ 71.92% 44.12% 84.05% 68.31%
500 Full FT 67.78% 43.34% 86.69% 74.50%
Prompt Tuning 62.44% 27.96% 80.20% 60.40%
Prefix Tuning 69.92% 41.74% 83.26% 66.72%
LoRA 71.08% 45.30% 84.92% 69.84%
$(IA)^{3}$ 67.87% 39.54% 84.25% 69.17%
1,000 Full FT 75.22% 50.81% 85.49% 73.83%
Prompt Tuning 69.63% 39.74% 82.58% 65.30%
Prefix Tuning 73.44% 47.02% 84.16% 68.84%
LoRA 74.76% 49.54% 86.67% 74.16%
$(IA)^{3}$ 63.95% 35.41% 84.07% 68.13%
DeepSeekCoder-1.3B 100 Full FT 44.86% 0.00% 81.48% 68.21%
Prompt Tuning 47.43% 14.34% 78.45% 56.93%
Prefix Tuning 59.82% 20.70% 82.35% 67.39%
LoRA 54.90% 16.98% 83.48% 67.47%
$(IA)^{3}$ 50.93% 8.63% 74.02% 48.52%
200 Full FT 72.76% 50.97% 85.25% 72.17%
Prompt Tuning 60.80% 30.66% 77.63% 55.49%
Prefix Tuning 69.20% 39.65% 82.52% 67.05%
LoRA 66.65% 36.72% 83.18% 66.49%
$(IA)^{3}$ 64.97% 32.56% 77.91% 55.91%
500 Full FT 66.34% 40.34% 83.32% 69.66%
Prompt Tuning 62.13% 30.85% 80.12% 60.25%
Prefix Tuning 71.96% 44.14% 83.91% 68.70%
LoRA 72.94% 45.89% 83.79% 69.70%
$(IA)^{3}$ 63.07% 34.91% 79.17% 58.36%
1,000 Full FT 72.80% 45.63% 82.50% 65.25%
Prompt Tuning 70.24% 42.31% 82.04% 64.10%
Prefix Tuning 70.26% 42.97% 84.43% 69.58%
LoRA 71.84% 44.27% 86.13% 73.00%
$(IA)^{3}$ 73.33% 46.68% 80.33% 60.65%
According to the results reported in Table 7, it is evident that when the data is limited to 100 or 200 samples, the effectiveness of both GraphCodeBERT and DeepSeek-Coder-1.3B with PEFT methods in detecting method-level code smells drops dramatically. This significant decline in performance highlights the challenges small and large LMs face in low-resource scenarios. However, when the number of training samples of CC or CM increases to 1,000, both models show improved performance compared to when only 100 samples are available. This improvement underscores that a larger dataset enables both small and large LMs using PEFT methods to achieve substantially better performance in code smell detection. Moreover, with 1,000 samples, GraphCodeBERT and DeepSeekCoder-1.3B with PEFT methods are able to yield reasonably good results compared to training with the original datasets for both types of method-level code smells. For example, the F1-scores of GraphCodeBERT training with 1,000 samples only degrade by no more than 3.26%, and the MCC decrease by no more than 7.22% compared to training with sufficient data for detecting CM, despite the dataset size being reduced by 90.35% ((10,366 - 1,000) / 10,366). This highlights the benefit of PEFT in resource-limited settings, as it enables the reduction of not only training parameters but also the amount of required training data, while still maintaining strong effectiveness on code smell detection. Notably, when the number of training samples increases to 1,000, GraphCodeBERT and DeepSeek-Coder-1.3B with PEFT methods demonstrate effectiveness in detecting CC and CM that is on par with, or even exceeds, that of full fine-tuning, further highlighting the superiority of PEFT methods in low-resource scenarios.
Secondly, compared to the results in Table 5 and Table 6, we can find that GraphCodeBERT with full fine-tuning performs better in low-resource scenarios than when there is sufficient data. This improvement can be attributed to the fact that, with ample data, full fine-tuning may lead to overfitting, causing the model to memorize rather than generalize the data. In contrast, in low-resource scenarios, the model is forced to learn more generalizable patterns, thus improving its performance.
Finally, in low-resource scenarios, LoRA tends to be the most effective PEFT method. As indicated by the results of RQ1 (see Section 4.1.4), for both GraphCodeBERT and DeepSeek-Coder-1.3B, fine-tuning with prefix tuning enables them to achieve their best performance compared to the other three PEFT methods. However, when data is scarce, GraphCodeBERT and DeepSeek-Coder-1.3B often attain their highest F1-scores or MCC values through LoRA. It suggests that under different resource settings, the most effective PEFT method may vary.
Answer to RQ3: In low-resource scenarios, the effectiveness of PEFT methods for code smell detection diminishes considerably when data is limited to 100 or 200 samples. However, as the number of training samples increases, the performance of PEFT methods shows significant improvement, and in some cases, even surpasses that of full fine-tuning.
4.4. RQ4: What is the effectiveness of PEFT methods compared to state-of-the-art code smell detection approaches?
Prompt Template for Detecting Complex Conditional with GPT-3.5-turbo You are a professional Java programmer. Given a piece of Java code, determine whether it contains “Complex Conditional” code smell. If it does, return only “Yes”, else return “No”. Definition of “Complex Conditional”: This code smell occurs when a conditional statement is excessively intricate. ### Input: [Code Sample] ### Output: [Yes/No]
Prompt Template for Detecting Complex Method with GPT-3.5-turbo You are a professional Java programmer. Given a piece of Java code, determine whether it contains “Complex Method” code smell. If it does, return only “Yes”, else return “No”. Definition of “Complex Method”: This code smell occurs when a method exhibits high cyclomatic complexity. ### Input: [Code Sample] ### Output: [Yes/No]
Traditional heuristic-based tools for code smell detection have been widely used in academia (sharma2021qscored, ; shah2023mining, ; barbosa2020revealing, ), while DL has recently gained widespread adoption in research on code smell detection, achieving impressive performance (zhang2024code, ). In this RQ, we established both heuristic-based (i.e., DesigniteJava) and best-in-class DL detectors (i.e., DeepSmells (ho2023fusion, ) and AE-Dense (sharma2021code, )), as well as a general-purpose LLM (i.e., GPT-3.5-turobo [^3] https://platform.openai.com/docs/models/o1) which has not been specifically tuned for method-level code smell detection, as baselines. To obtain valid responses from GPT-3.5-turbo in a zero-shot setting (kojima2024large, ), we designed the above prompt templates that include both instructions and definitions of CC or CM to guide ChatGPT in annotating the test sets. GraphCodeBERT with $(IA)^{3}$ for CC and UnixCoder with LoRA for CM are selected for comparison. The results of this evaluation are shown in Table 8.
From the results reported in Table 8, we can find that GraphCodeBERT with $(IA)^{3}$ and UnixCoder with LoRA exhibit significantly better performance than the state-of-the-art code smell detection approaches and GPT-3.5-turbo in detecting CC and CM, respectively. Specifically, for CC, GraphCodeBERT with $(IA)^{3}$ significantly surpasses the effectiveness of the other four baseline methods with an F1-score disparity ranging from 23.41% to 47.94% and an MCC difference ranging from 13.24% to 50.59%. For CM, UnixCoder with LoRA achieves an F1-score improvement ranging from 11.22% to 54.32% and an MCC improvement ranging from 10.28% to 72.91%.
Table 8. Results of GraphCodeBERT with $(IA)^{3}$ and UnixCoder with LoRA compared to state-of-the-art code smell detection approaches on detecting CC and CM (bold: highest values per code smell)
Code Smell Method Precision Recall F1 MCC
CC DesigniteJava 34.58% 98.74% 51.22% 43.67%
DeepSmells 21.61% 47.17% 29.64% 6.32%
AE-Dense 42.90% 53.82% 47.74% 15.20%
GPT-3.5-turbo 42.86% 73.58% 54.17% 42.66%
GraphCodeBERT with $(IA)^{3}$ 83.33% 74.30% 77.58% 56.91%
CM DesigniteJava 61.88% 99.27% 76.23% 65.92%
DeepSmells 33.82% 56.80% 42.39% 4.65%
AE-Dense 24.33% 55.76% 33.88% 9.78%
GPT-3.5-turbo 81.35% 73.06% 76.98% 67.28%
UnixCoder with LoRA 86.99% 90.66% 88.20% 77.56%
Answer to RQ4: In comparison with existing code smell detection approaches, LMs tuned with PEFT methods exhibit notable performance improvements on identifying CC and CM, highlighting the effectiveness of PEFT methods for code smell detection.
5. Discussion
Approaches with improved effectiveness for CC code smell detection are needed: According to Fowler (fowler1999refactoring, ), four factors contribute to making programs difficult to work with, one of which is Complex Conditional (CC) logic that complicates modification. The presence of CC in source code hinders testing and maintenance (fowler1999refactoring, ; liu2014automated, ; alhefdhi2022framework, ). Thus, identifying CC code smells in software can significantly improve code quality (liu2014automated, ). However, the effectiveness of existing methods for detecting this type of code smell is considerably limited. As shown in Table 8, existing methods — whether heuristic-based (e.g., DesigniteJava), DL-based (e.g., DeepSmells and AE-Dense), or GPT-3.5-turbo — demonstrate limited success in detecting CC, with all MCC values below 45%, and some even as low as 6.32%. Although our proposed method, GraphCodeBERT with $(IA)^{3}$, improved the performance of CC code smell detection significantly, reaching an F1-score of 77.58% and an MCC value of 56.91%, it still falls short of expectations due to challenges such as the considerable existence of false positives and negatives, which may limit its practical applicability. Accurate identification of CC code smells can help developers better refactor the code, ultimately improving the maintainability of the software system. Despite the progress made, the detection of CC remains a challenging task that requires further improvements.
Model size is not the decisive factor contributing to the effectiveness of code smell detection: Previous research has demonstrated that larger models with more parameters tend to deliver better performance on code-related tasks (nijkamp2022codegen, ; roziere2024codellama, ; weyssow2023exploring, ). For example, Weyssow et al. (weyssow2023exploring, ) suggested a potential shift in the software engineering domain, moving away from the long-established practice of fine-tuning smaller LMs like CodeT5 over the past 3-4 years, and towards embracing larger LLMs. However, the results of RQ1 indicate that, for the method-level code smell detection task, model size is not the sole determinant of performance. In fact, all the small models, with the exception of CodeT5, generally perform as well as or better than larger LLMs in detecting CC and CM (see Section 4.1.3). Furthermore, within the same model family, larger models do not always outperform its smaller-sized counterparts (see Section 4.1.2). In contrast, the alignment between the model and the specific task at hand seems to play a crucial role in determining effectiveness for code smell detection. While CodeT5, when fine-tuned with PEFT methods, has demonstrated strong performance in tasks such as code summarization and code change prediction (liu2023empirical, ; fan2024exploring, ), it surprisingly underperforms in the task of code smell detection. This suggests that models may exhibit varying performance across different downstream tasks, emphasizing the importance of aligning a model’s design and fine-tuning with the specific requirements of each target task. Therefore, it is essential to focus on selecting models that are well-suited to the method-level code smell detection task, rather than relying solely on model size as a performance indicator.
Considering the model used and the availability of resources when selecting the appropriate PEFT method for code smell detection: Based on the results of RQ1 and RQ3, we can find that for method-level code smell detection, the most effective PEFT method can differ significantly depending on the models and the amount of available data. From Table 5 and Table 6, it is obvious that each model responds to different PEFT techniques in unique ways, suggesting that the optimal PEFT method is rather model-dependent. The effectiveness of PEFT methods can change even for the same model under different resource scenarios. For instance, in low-resource settings with limited data, the best PEFT method for models like GraphCodeBERT and DeepSeek-Coder-1.3B shifts from prefix tuning to LoRA. Moreover, peak GPU consumption can vary across different models (see Section 4.1.2). Generally, $(IA)^{3}$ requires more GPU resources than LoRA for code smell detection, whereas the reverse is true for StarCoderBase-7B. These variations underscore that there is no universally best PEFT method for method-level code smell detection; instead, the most suitable method depends on the specific model and resource constraints. Therefore, we recommend that practitioners consider both the model’s characteristics and the availability of resources, including data volume and GPU memory, when selecting the appropriate PEFT method for a given task.
Start model selection for code smell detection with a smaller dataset to optimize resource usage: In low-resource scenarios, when the dataset size was reduced by more than 90%, the models’ effectiveness did not degrade significantly, and they still yielded promising results, as shown in the results of RQ3 (see Section 4.3). This finding indicates that, even with a significantly smaller dataset, both small and large LMs can maintain robust performance for the code smell detection task. Consequently, beginning fine-tuning experiments with a smaller-size dataset can be an effective strategy to optimize resource usage. This approach allows researchers and practitioners to initially evaluate a wide range of LMs in a shorter time and wither fewer computational resources, saving substantial effort. It enables them to identify the most suitable model for the specific task through an initial evaluation before committing to the more resource-intensive process of training on larger datasets. Additionally, it helps in assessing the trade-offs between resource consumption and performance, providing insight into the models’ scalability. By leveraging smaller datasets for initial evaluation, one can make informed decisions about which models are likely to perform well with limited resources, and further fine-tune or scale up as necessary, without wasting valuable time or computational power.
6. Threats to Validity
Construct Validity: One primary threat to the construct validity of this study is the construction of the datasets. We first used DesigniteJava (designitejava, ) to detect potential positive and negative samples, then manually reviewed all the candidate data to validate the datasets, ensuring there were no false positives or negatives.
First, relying solely on DesigniteJava for initial sample detection may introduce bias. However, DesigniteJava is a well-established and widely adopted tool in Java code smell detection (chaniotaki2021architecture, ; sharma2021qscored, ; shah2023mining, ), providing a robust foundation for the identification of potential code smells. We also acknowledge that future studies could benefit from incorporating additional tools in the initial code smell detection process, which may help mitigate the threat to construct validity by providing more diverse perspectives on the data.
Second, the manual validation process, while aimed at ensuring dataset accuracy, could introduce some subjectivity due to its reliance on human judgment. To minimize this threat, we conducted separate pilot validations for CC and CM before formally annotating all the data samples. This pilot phase helped ensure that the two annotators used the same criteria for validation, thus improving consistency and reducing potential bias in the dataset labeling.
External Validity: In this study, we evaluated the effectiveness of state-of-the-art PEFT methods using 4 small LMs and 6 LLMs. However, due to limited computational resources, we did not include larger models with over 7B parameters, which may impact the generalizability of our findings. Additionally, we only employed four popular and typical PEFT techniques, i.e., prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$. Other PEFT methods, such as adapter tuning (houlsby2019parameter, ), could be explored in our future work. Moreover, our study only applied PEFT methods to detect Java code smells. Expanding this research to include code smells in other programming languages, such as C# and Python, could provide further insights into the effectiveness and adaptability of PEFT methods across diverse language-specific code smell patterns.
Internal Validity: The settings of hyper-parameters for the four PEFT methods and full fine-tuning play a crucial role in the performance of the models. Variations in these settings could lead to different results, potentially affecting the reliability of our findings. To mitigate this threat, we referred to previous works and conducted pilot experiments to determine the appropriate hyper-parameter values.
Additionally, data leakage represents another potential threat to internal validity. First, the poor zero-shot performance of GPT-3.5-turbo in detecting CC and CM code smells suggests that the model was not exposed to the test data (see Section 4.4), as we would expect stronger zero-shot effectiveness if GPT-3.5-turbo had memorized the data. Second, the significant improvements in performance for both small and large LMs following the application of PEFT techniques further suggest that memorization is unlikely to be an issue. Such substantial performance gains would not occur if the models had already memorized the test sets. Furthermore, Cao et al. (cao2024concerned, ) highlight that code LMs do not consistently exhibit reduced performance on data extending beyond their training cut-off dates. This careful management of data, along with the observed performance improvements after fine-tuning, ensures that data leakage was not a major threat, thereby maintaining the internal validity of our findings.
7. Related Work
7.1. Code Smell Detection
Existing approaches to code smell detection predominantly rely on heuristic-based methods, which utilize predefined rules to identify problematic code patterns (pmd, ; designitejava, ; campbell2013sonarqube, ; fokaefs2007jdeodorant, ; vidal2015jspirit, ; moha2010decor, ). These methods aim to flag potential smells by analyzing the source code against established standards. For instance, PMD (pmd, ), a widely used static code analysis tool, supports 16 programming languages and applies rule sets to detect code smells. These rule sets can be tailored to suit specific project requirements or design standards, making the tool adaptable for different contexts. Similarly, Sharma (designitejava, ) introduced DesigniteJava, an advanced tool capable of identifying code smells at multiple levels — ranging from architecture and design to implementation.
Traditional heuristic-based code smell detectors face inherent challenges, such as difficulty in defining effective thresholds for detection. In response, Machine Learning (ML) and Deep Learning (DL) techniques have gained traction as promising alternatives. Maiga et al. (maiga2012support, ) implemented SVMDetect based on an ML technique — support vector machines. Fontana et al. (arcelli2016comparing, ) evaluated 16 ML algorithms across four types of code smells using 74 software systems, reporting that J48 and Random Forest delivered the highest performance.
Liu et al. (liu2021deep, ) utilized deep neural networks and advanced DL techniques to perform useful feature selection from source code and map these features to labels (smelly or not). Sharma et al. (sharma2021code, ) trained Convolution Neural Network (CNN), Recurrent Neural Network (RNN), and autoencoder DL models in different configurations to detect code smells. Their work confirms the feasibility of applying transfer-learning for code smell detection. Ho et al. (ho2023fusion, ) developed DeepSmells based on a CNN and long short-term memory (LSTM) networks for detecting code smells. They conducted an empirical study to compare DeepSmells with other well-established detection tools. Their results demonstrate that DeepSmells outperforms the baselines, achieving superior performance. They also anticipated that leveraging CodeBERT would enhance the effectiveness of code smell detection. Recently, Wu et al. (wu2024ismell, ) trained a Mixture of Experts (MoE) model, leveraging specialized expert toolsets for code smell detection. They subsequently utilized the detection results to prompt LLMs for refactoring.
Most prevailing ML- or DL-based approaches for code smell detection train models from scratch. Compared to their works, we harness the power of LLMs using PEFT methods to detect two types of complex method-level code smells (Complex Conditional and Complex Method), aiming to enhance the models’ ability to capture the nuanced relationships between different types of method-level code smells.
7.2. PEFT Methods on Software Engineering Tasks
Several studies have evaluated PEFT methods on various SE tasks. Wang et al. (wang2023one, ) fine-tuned UniXcoder and CodeT5 by inserting the parameter-efficient structure adapter for code search and summarization. According to their results, adapter tuning has shown effectiveness in cross-lingual and low-resource scenarios, surpassing full fine-tuning and effectively mitigating catastrophic forgetting. Liu et al. (liu2023empirical, ) carried out experiments by fine-tuning code models with adapter, prefix tuning, LoRA, and MHM (he2022towards, ) on both the code understanding task and code generation task. Their findings indicate that adapter tuning and LoRA can achieve promising performance on these two SE tasks with reduced training costs and memory requirements. Liu et al. (liu2024delving, ) examined whether adapter tuning and LoRA can perform well on code-change-related tasks, specially just-in-time defect prediction and commit message generation. Their study demonstrates that PEFT methods can serve as a powerful alternative to full fine-tuning for dynamic code comprehension. Weyssow et al. (weyssow2023exploring, ) compared PEFT methods, including prompt tuning and others, with In-Context Learning (ICL) in terms of code generation. Their investigation reveals the superiority of PEFT over ICL and highlights the potential opportunities of tuning LLMs in diverse software engineering scenarios. Li et al. (li2024comprehensive, ) created an instruction dataset of Automated Program Repair (APR) using prompt engineering. They employed this dataset to fine-tune four LLMs using four different PEFT methods. The results show that $(IA)^{3}$ achieves the highest fixing ability compared to prompt tuning, prefix tuning, and LoRA.
These studies focus on evaluating PEFT methods across a range of software engineering tasks, such as code summarization and code generation. Compared to their works, our work pays specific attention to a binary classification task, aiming to assess the effectiveness of LLMs with PEFT methods in identifying different types of code smells.
8. Conclusions and Future Work
In this work, we investigated the effectiveness of various PEFT methods on different LMs for code smell detection. We also evaluated how the performance of these PEFT methods varies under different conditions, such as varying hyper-parameter settings and the availability of training data. Our analysis reveals the efficiency of PEFT methods compared to full fine-tuning, underscoring the potential of PEFT methods to enhance method-level code smell detection and offer valuable insights for optimizing LMs for other software engineering tasks. Additionally, we have constructed high-quality datasets for CC and CM code smells, and we provided the replication package of this work online (replicationPackage, ) to facilitate future research and ensure the reproducibility of our study results.
In the next step, we plan to further explore how PEFT performs in cross-language scenarios for code smell detection. Moreover, we plan to investigate the application of PEFT methods for fine-tuning LMs to assist in code smell refactoring, thereby improving overall code quality and maintainability by providing just-in-time code smell detection and refactoring suggestions.
Data Availability
The replication package for this work has been made available at https://github.com/MabelQi/PEFT4CSD.
Acknowledgements.
This work has been partially supported by the National Natural Science Foundation of China (NSFC) with Grant No. 62172311.

### v2

@@
-A Comprehensive Evaluation of Parameter-Efficient Fine-Tuning on Method-Level Code Smell Detection
+A Comprehensive Evaluation of Parameter-Efficient Fine-Tuning on Code Smell Detection
 Abstract.
@@
 However, LLMs have not yet been explored for code smell detection, and their effectiveness for this task remains unclear.
-Furthermore, no comprehensive investigation has been conducted on the efficiency of PEFT methods for method-level code smell detection.
-In this regard, we systematically evaluate the effectiveness of state-of-the-art PEFT methods on both small and large Language Models (LMs) for method-level code smell detection.
+Furthermore, no comprehensive investigation has been conducted on the efficiency of PEFT methods for code smell detection.
+In this regard, we systematically evaluate the effectiveness of state-of-the-art PEFT methods on both small and large Language Models (LMs) for code smell detection.
 To begin, we constructed high-quality Java code smell datasets sourced from GitHub.
-We then fine-tuned four small LMs and six LLMs using various PEFT techniques, including prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$, for code smell detection.
+We then fine-tuned four small LMs (SLMs) and five LLMs using various PEFT techniques, including prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$, for code smell detection.
 Our comparison against full fine-tuning revealed that PEFT methods not only achieve comparable or better effectiveness but also consume less peak GPU memory.
-Our analysis further explored the performance of small LMs versus LLMs in the context of code smell detection.
-Surprisingly, we found that LLMs did not outperform small LMs in this specific task, suggesting that smaller models may be more suited for method-level code smell detection.
+Our analysis further explored the performance of SLMs versus LLMs in the context of code smell detection.
+Generally, LLMs outperform SLMs on detecting Complex Conditional, Complex Method, and Feature Envy.
+However, SLMs demonstrate superior performance on detecting Data Class.
 We also investigated the impact of varying hyper-parameter settings and the availability of training data.
@@
 In contrast, we observed significant performance gains when the training dataset size was increased.
-Finally, our results show that LMs combined with PEFT methods outperform traditional heuristic-based detectors, advanced DL-based approaches, and GPT-3.5-turbo in method-level code smell detection.
-Specifically, MCC values improve by 13.24% to 50.59% for Complex Conditional and 11.22% to 54.32% for Complex Method, demonstrating a more effective solution for identifying these code smells.
-In conclusion, our work provides a comprehensive evaluation of PEFT methods on both small and large LMs for method-level code smell detection.
-The results highlight the efficiency and effectiveness of PEFT methods in this domain, offering valuable insights for optimizing LMs for method-level code smell detection.
+Finally, our results show that LMs with PEFT methods outperform traditional heuristic-based detectors, advanced DL-based approaches, and general-purpose LLMs under In-Context Learning in the task of code smell detection.
+Specifically, we observe improvements in MCC ranging from 4.11% to 19.25% across the four types of code smells studied, demonstrating a more effective solution for this task.
+In conclusion, our work provides a comprehensive evaluation of PEFT methods on both small and large LMs for code smell detection.
+The results highlight the efficiency and effectiveness of PEFT methods in this domain, offering valuable insights for optimizing LMs for code smell detection.
 We suggest that the choice of PEFT method for code smell detection should be guided by the model used, data availability, and computational resources.
@@
 Code Smell Detection, Large Language Model, Parameter-Efficient Fine-Tuning
-copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: TOSEMjournalvolume: 0journalnumber: 0article: 0publicationmonth: 0ccs: Software and its engineering Software maintenance tools
+copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: TOSEMjournalvolume: 0journalnumber: 0article: 0publicationmonth: 0ccs: Software and its engineering Software maintenance tools
 1. Introduction
@@
 Besides, code smells identified by such heuristics-based methods can be subjectively interpreted by developers, leading to a significant gap between detection results and developers’ agreement (azeem2019machine, ; fontana2016antipattern, ; mantyla2006subjective, ; schumacher2010building, ).
-To address these issues, many researchers have incorporated machine learning (ML) and deep learning (DL) algorithms to mitigate the performance instability in code smell detection that arises from the manual threshold-setting process in traditional detectors (fontana2013code, ; arcelli2016comparing, ; sharma2021code, ; liu2021deep, ).
+To address these issues, many researchers have incorporated machine learning (ML) and deep learning (DL) algorithms to mitigate the performance instability in code smell detection that arises from the manual threshold-setting process in traditional detectors (fontana2013code, ; fontana2016comparing, ; sharma2021code, ; liu2021deep, ).
 By learning to discern between smelly and non-smelly code elements, ML and DL techniques offer a solution to the subjectivity problem inherent in heuristics-based methods (nucci2018detecting, ).
@@
 The exceptional performance of LLMs in NLP tasks (e.g., text classification, generation, and summarization) has garnered significant attention and interest from the Software Engineering (SE) domain.
-Researchers have begun employed LLMs, particularly code-specific LLMs like CodeBERT-Base (feng2020codebert, ) and CodeT5 (wang2021codet5, ), using fine-tuning techniques to tackle various SE tasks, such as code review (lu2023llamareviewer, ) and vulnerability repair (le2024study, ).
+Researchers have begun employed LLMs, particularly code-specific LLMs like CodeBERT-Base (feng2020codebert, ) and CodeT5 (wang2021codet5, ), to tackle various SE tasks, such as code review (lu2023llamareviewer, ) and vulnerability repair (le2024study, ).
+However, directly using LLMs without any task-specific fine-tuning for code smell detection yields suboptimal results (mesbah2025leveraging, ; silva2024detecting, ; sadik2025benchmarking, ), as these models lack adaptation to the unique demands of the task.
+Fine-tuning enables the adaptation of Language Models (LMs) to specific downstream tasks by updating model parameters, allowing them to better capture task-relevant features and improve performance.
 Among all the fine-tuning methods, Parameter-Efficient Fine-Tuning (PEFT) has emerged as a promising approach, offering an efficient alternative to full fine-tuning, particularly when computational resources are limited.
 PEFT methods have shown potential in adapting LLMs to various downstream tasks with fewer parameters and reduced resource requirements, thus expanding the practical applicability of LLMs in diverse SE contexts.
-Previous studies have explored the efficiency of PEFT methods for Language Models (LMs) on automated program repair (li2024comprehensive, ), code generation (liu2023empirical, ), code search (wang2023one, ), and etc.
-However, to the best of our knowledge, there is currently no comprehensive research exploring the use of LLMs for code smell detection, nor any in-depth studies analyzing the application of PEFT methods in this area.
-Moreover, utilizing PEFT methods for code smell detection can facilitate just-in-time code smell detection.
+Previous studies have explored the efficiency of PEFT methods for LMs on automated program repair (li2024comprehensive, ), code generation (liu2023empirical, ), code search (wang2023one, ), etc.
+However, to the best of our knowledge, no prior research has focused on improving the performance of LLMs for code smell detection, nor has any in-depth study been conducted analyzing the application of PEFT methods in this area.
+Additionally, identifying the code smells selected in this study requires both method- and class-level code understanding.
+While existing research on PEFT methods has predominantly focused on code-related tasks at the method level (li2024comprehensive, ; liu2023empirical, ; wang2023one, ; liu2024delving, ; weyssow2023exploring, ), a significant gap remains in assessing their effectiveness at the class level.
+By leveraging the task of code smell detection, our study aims to fill this gap by systematically exploring the capacities of LMs with PEFT methods in class-level code understanding.
+Furthermore, utilizing PEFT methods for code smell detection can facilitate just-in-time code smell detection.
 Traditional heuristic-based methods for code smell detection are often applied after the software systems have been completed (designitejava, ; pmd, ; marinescu2005iplasma, ).
@@
 To this end, we conduct a series of experiments on fine-tuning both small and large LMs with PEFT methods on the code smell detection task.
-In this paper, we define small LMs as those with fewer than 1B parameters, while LMs with 1B parameters or more are considered large LMs.
-Specifically, we construct high-quality Java code smell datasets comprising two types of widely studied and sophisticated method-level code smells, i.e., Complex Conditional (CC) and Complex Method (CM), as no existing datasets are available.
-We then apply various PEFT methods, including prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$, to fine-tune four popular small LMs and six cutting-edge LLMs with parameter sizes ranging from 125M to 7B.
+In this paper, we define small LMs (SLMs) as those with fewer than 1B parameters, while LMs with 1B parameters or more are considered LLMs.
+Specifically, we construct high-quality Java code smell datasets comprising four types of widely studied and sophisticated code smells at both the method and class level, i.e., Complex Conditional (CC), Complex Method (CM), Feature Envy (FE), and Data Class (DC).
+These datasets are developed to address the lack of suitable existing resources, as most available datasets either do not conform to the required data formats or are based on outdated codebases.
+We then apply various PEFT methods, including prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$, to fine-tune four popular SLMs and five cutting-edge LLMs with parameter sizes ranging from 125M to 7B.
 The effectiveness of these PEFT methods is assessed for code smell detection, and their performance is compared to that of full fine-tuning in terms of both effectiveness and peak GPU memory usage.
-Additionally, we investigate the performance of PEFT techniques under different hyper-parameter settings and evaluate their effectiveness in low-resource scenarios with varying numbers of training samples on the method-level code smell detection task.
+Additionally, we investigate the performance of PEFT techniques under different hyper-parameter settings and evaluate their effectiveness in low-resource scenarios with varying numbers of training samples on the code smell detection task.
 Finally, we employ a heuristics-based static analysis tool alongside two state-of-the-art DL-based code smell detection tools to compare the performance of PEFT methods with existing approaches.
 The main contributions of this work are as follows:
-  • We mine Java repositories from GitHub and employ a traditional heuristic-based detector, along with manual validation, to create high-quality datasets for two types of method-level code smells, i.e., Complex Conditional and Complex Method.
+  • We mine Java repositories from GitHub and employ traditional heuristic-based detectors, along with a two-stage manual review, to create high-quality and balanced datasets for four types of code smells, i.e., Complex Conditional, Complex Method, Feature Envy, and Data Class.
   • We conduct a comprehensive evaluation of multiple PEFT methods on both small and large LMs for code smell detection.
-Our results indicate that LLMs, when fine-tuned, do not always outperform small LMs.
-Additionally, PEFT methods achieve better or comparable effectiveness compared to full fine-tuning.
+Our results indicate that LLMs, when fine-tuned, do not always outperform SLMs.
+Moreover, PEFT methods achieve better or comparable effectiveness compared to full fine-tuning.
+Overall, LMs with PEFT methods exhibit promising capabilities in method- and class-level code understanding.
   • We explore the performance of PEFT methods under different hyper-parameter settings and in low-resource scenarios.
@@
 However, the performance of PEFT methods improves with more training samples.
-  • We compare a heuristic-based detector and two best-in-class DL-based code smell detection tools with both small and large LMs using PEFT methods on code smell detection.
-Our results demonstrate that both small and large LMs with PEFT methods consistently outperform the existing heuristic- and DL-based approaches in detecting the method-level code smells, highlighting the superiority of PEFT methods.
-  • We provide actionable suggestions and insights for both developers and researchers for using PEFT methods on method-level code smell detection.
+  • We compare a heuristic-based detector, two best-in-class DL-based detection tools, and two general-purpose LLMs using In-Context Learning (ICL) against our PEFT-based method on code smell detection.
+Our results demonstrate that LMs with PEFT methods consistently outperform all baselines across multiple code smell types, highlighting the superiority of PEFT methods.
+  • We provide actionable suggestions and insights for both developers and researchers for using PEFT methods on code smell detection.
 The remainder of this paper is organized as follows: Section 2 presents the background information.
@@
 The right side of Figure 2 illustrates several state-of-the-art PEFT methods employed in this study, including prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$.
+These methods were selected based on several considerations.
+First, they are all supported by the widely used Hugging Face PEFT library [^1] https://github.com/huggingface/peft, which provides a standardized and efficient framework for applying PEFT to LMs.
+Second, according to the PEFT library’s documentation [^2] https://huggingface.co/docs/peft/index, PEFT methods are categorized into three main types: Prompt-based methods, LoRA methods, and $(IA)^{3}$.
+The four selected methods span these categories, making them representative of the major approaches currently available.
+Finally, these methods are not only recent but also widely adopted in contemporary research (weyssow2023exploring, ; li2024comprehensive, ; mannisto2025comparative, ; zhou2024empirical, ; taylor2022galactica, ; esmaeili2024empirical, ; wistuba2024choice, ).
+Their demonstrated effectiveness across various tasks and their popularity in both the NLP and SE communities make them ideal candidates for this study.
+By selecting these four methods, we ensure that our experiments cover a range of the most promising and commonly adopted techniques in PEFT.
 Prompt tuning (lester2021power, ) directly appends learnable vectors, known as soft prompts, to the initial input embedding layer.
@@
 3. Methodology
-Figure 2. Left: architecture of our proposed method for method-level code smell detection. Right: a detailed illustration of the PEFT module structure for the four PEFT methods.
+Figure 2. Left: architecture of our proposed method for code smell detection. Right: a detailed illustration of the PEFT module structure for the four PEFT methods.
 3.1. Research Goal
 Figure 2 illustrates the architecture of our LLM-based code smell detector using PEFT methods.
-We frame the method-level code smell detection task as a binary classification problem, where the input is a piece of code.
+We frame the code smell detection task as a binary classification problem, where the input is a piece of code.
 A sequence classification head is added on top of the LLM to handle the classification task.
@@
 3.2. Research Questions
-This study aims to evaluate the effectiveness of state-of-the-art PEFT methods (prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$) in small LMs and LLMs for method-level code smell detection.
+This study aims to evaluate the effectiveness of state-of-the-art PEFT methods (prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$) in SLMs and LLMs for code smell detection.
 To achieve this goal, we conducted a set of experiments to answer the following Research Questions (RQs):
-RQ1: What is the effectiveness of different PEFT methods and full fine-tuning on various small LMs and LLMs?
-Previous studies have shown remarkable results using PEFT methods to fine-tune small LMs for software engineering tasks like code clone detection and defect detection (liu2023empirical, ; liu2024delving, ; li2024comprehensive, ; wang2023one, ; weyssow2023exploring, ).
+RQ1: What is the effectiveness of different PEFT methods and full fine-tuning on various SLMs and LLMs?
+Previous studies have shown remarkable results using PEFT methods to fine-tune SLMs for software engineering tasks like code clone detection and defect detection (liu2023empirical, ; liu2024delving, ; li2024comprehensive, ; wang2023one, ; weyssow2023exploring, ).
 However, these studies have not paid attention to the effectiveness of PEFT methods for code smell detection.
@@
 For LoRA, the key hyper-parameter is the rank $r$ (hu2022lora, ), while in $(IA)^{3}$, it is the selection of activated modules (liu2022few, ).
-Due to our limited computational resources and the popularity of LoRA compared to other PEFT methods (gupta2024beyond, ; zhang2024star, ), we decided to employ LoRA to fine-tune a small LM and an LLM with different $r$ values in a range of {8, 32, 64, 640, 1,280, 2,560} to explore the effect of hyper-parameter settings on PEFT methods for code smell detection.
+Due to our limited computational resources and the popularity of LoRA compared to other PEFT methods (gupta2024beyond, ; zhang2024star, ), we decided to employ LoRA to fine-tune a SLM and an LLM with different $r$ values in a range of {8, 32, 64, 640, 1,280, 2,560} to explore the effect of hyper-parameter settings on PEFT methods for code smell detection.
 RQ3: What is the effectiveness of PEFT methods in low-resource scenarios?
@@
 Existing code smell detectors include traditional heuristic-based and DL-based methods.
-However, heuristic-based approaches for code smell detection depend on rules and heuristics derived from software metrics (moha2010decor, ; palomba2015mining, ; palomba2016a, ; fokaefs2007jdeodorant, ), which may suffer from several limitations (e.g.
-different tools may produce inconsistent results, complicating developers’ refactoring decisions (fontana2012automatic, )).
-As a result, researchers have been exploring the application of DL techniques as alternatives for code smell detection due to the drawbacks of heuristic-based methods (pecorelli2022on, ), and state-of-the-art DL-based code smell detectors have demonstrated strong effectiveness in identifying code smells (alazba2023deep, ; sharma2021code, ).
-With this RQ, we aim to assess how LMs utilizing PEFT methods compare against heuristic-based code smell detectors and established DL techniques, including general-purpose LLMs, in terms of their ability to detect and classify code smells accurately.
+Additionally, the use of general-purpose LLMs with In-Context Learning (ICL) has become increasingly popular in SE tasks (zhou2024large, ; shin2023prompt, ; li2025large, ; xu2024unilog, ), and such models can also be leveraged for code smell detection.
+With this RQ, we aim to assess how LMs utilizing PEFT methods compare against heuristic-based code smell detectors, established DL techniques, and general-purpose LLMs under zero- and few-shot settings for the task of code smell detection.
 3.3. Dataset Construction
-In this section, we elaborate on the process with the following steps of constructing the high-quality datasets for two types of method-level code smells used to fine-tune both small and large LMs for code smell detection.
+In this section, we elaborate on the process with the following steps of constructing the high-quality and balanced datasets for four types of code smells used to fine-tune both small and large LMs for code smell detection.
 We created our own datasets because most existing datasets only offer code metrics related to the code smells instead of code snippets, which are not usable for our purposes (sharma2021qscored, ; binh2022mlcodesmell, ; madeyski2023detecting, ).
-While some code smell datasets provide valid code samples, they do not include Complex Conditional and Complex Method (madeyski2020mlcq, ), two types of method-level code smells.
-Additionally, the source projects from these existing datasets are quite outdated (fabio2015landfill, ; palomba2018a, ; fontana2016comparing, ; amorim2015experience, ).
-Considering this, we aim to construct new datasets sourced from the latest trending GitHub projects.
-Figure 3 illustrates the overview process of dataset construction.
+While some code smell datasets provide valid code samples, they do not cover the specific types of code smells we focus on (e.g., Complex Conditional and Complex Method) (madeyski2020mlcq, ).
+To better align with current software development practices, we constructed new datasets based on trending GitHub projects created within the past six years.
+Furthermore, since data imbalance may negatively impact model performance on code smell detection (khleel2024improving, ; nandini2025improving, ; fontana2016comparing, ; nunes2024tuning, ), we ensured that our constructed datasets contain an equal number of positive and negative samples.
+Figure 3 illustrates the overview of the dataset construction process.
 Figure 3. Overview of dataset construction
 3.3.1. Selected Code Smells
-In this study, we intend to use LLMs to detect two types of method-level code smells, i.e., Complex Conditional and Complex Method. The definitions of the two types of code smells are provided below:
-  • Complex Conditional (CC): This code smell occurs when a conditional statement is excessively intricate (sharma2017house, ).
-  • Complex Method (CM): This code smell occurs when a method exhibits high cyclomatic complexity (sharma2017house, ).
-We chose CC and CM because they have been widely studied in research on DL-based code smell detection (e.g., (sharma2021code, ; ho2023fusion, ; zhang2024code, ; alazba2023deep, )).
-Another reason for selecting CC and CM is the availability of detection tools, such as DesigniteJava (designitejava, ), which can assist us in collecting candidate code samples that may exhibit these two types of code smells.
-We excluded very simple smells (e.g., Long Method), which can be easily identified by basic techniques.
-In contrast, detecting the two selected method-level code smells requires LLMs to spot specific features, pay attention to cohesion and coupling aspects, and analyze the entire method for its structural properties (sharma2021code, ), thereby expanding the ambition and enhance the applicability of this study.
+In this study, we intend to use LMs with PEFT methods to detect four types of code smells, comprising three at the method level (i.e., Complex Conditional, Complex Method, and Feature Envy) and one at the class level (i.e., Data Class).
+The definitions of these four code smell types are provided below:
+We selected CC, CM, FE, and DC because they are among the most extensively studied code smells in prior research on ML- and DL-based code smell detection (e.g., (sharma2021code, ; ho2023fusion, ; zhang2024code, ; alazba2023deep, ; fontana2016comparing, ; barbez2020machine, ; kim2017finding, ; liu2023deep, ; liu2018deep, ; skipina2024automatic, ; yang2024feature, ; fontana2013code, ; madeyski2020mlcq, )).
+Moreover, these code smells are frequently observed in real-world software projects and are known to significantly impact code quality, maintainability, and the likelihood of introducing defects (fowler1999refactoring, ).
+For example, CC and CM can lead to convoluted control flows that hinder code readability and increase cognitive load on developers, while FE and DC often indicate bad modularization and poor class design, which complicate maintenance and evolution (fowler1999refactoring, ; sjoberg2012questioning, ; khomh2009exploratory, ).
+Another reason for choosing these specific smell types is the availability of detection tools, such as DesigniteJava (designitejava, ) and PMD (pmd, ), which facilitates the collection of candidate code samples potentially exhibiting these smells.
+We excluded overly simple smells (e.g., Long Method) that can be easily identified using basic heuristics.
+In contrast, detecting the four selected code smells requires LLMs to capture subtle features, consider aspects such as cohesion and coupling, and analyze the overall structure of a code sample (sharma2021code, ), which expands the ambition and enhances the applicability of this study.
+At the same time, identifying these selected code smells allows us to assess the effectiveness of LMs fine-tuned with PEFT methods in both method-level and class-level code understanding.
+CC, CM, and FE are generally considered method-level code smells, while DC represents a class-level code smell.
+However, detecting FE requires comprehension of not only the focal method but also its surrounding class context and inter-class dependencies (fowler1999refactoring, ; tsantalis2009indentification, ; liu2023deep, ; alkharabsheh2021analysing, ).
+Therefore, it also demands class-level code understanding.
+The selection of these four code smells allows our study to address a key gap in existing research related to PEFT methods (li2024comprehensive, ; liu2023empirical, ; wang2023one, ; liu2024delving, ; weyssow2023exploring, ), which has primarily concentrated on the effectiveness of PEFT techniques on method-level code-related tasks.
 3.3.2. Java Repository Selection
-We used GitHub REST API [^1] https://docs.github.com/en/rest?apiVersion=2022-11-28 to search Java repositories from GitHub.
-From the extensive array of available Java projects hosted on GitHub, we set the following criteria:
-  • The repositories should be created between January 1, 2020 and April 30, 2024.
+To construct a code smell dataset that accurately represents modern coding practices, we selected Java repositories created in the past six years from GitHub.
+We used the GitHub REST API [^3] https://docs.github.com/en/rest?apiVersion=2022-11-28 to gather available Java projects that meet the following criteria:
+  • The repositories should be created between January 1, 2019 and April 30, 2024.
   • The repositories should exclude forked ones to avoid duplicates.
   • The repositories should have a minimum of 1,000 stars and 1,000 lines of code.
-These criteria ensured that the selected Java repositories were recent, popular, active, and non-trivial Java repositories.
-By employing the stringent filtering criteria, we obtained a list of 294 GitHub repositories coded in Java.
+These criteria ensured that the selected Java repositories were recent, popular, active, and non-trivial.
+By employing these stringent filtering criteria, we obtained a list of 401 GitHub repositories coded in Java.
 3.3.3. Potential Code Smell Detection
-DesigniteJava is an open-source tool that supports detection of 18 design and 10 implementation smells, which have been extensively used in many studies as a mature code smell detector (sharma2021qscored, ; shah2023mining, ; barbosa2020revealing, ).
-In this paper, we had developed DesigniteJava to detect potential CC and CM in the Java repositories we collected in Section 3.3.2.
-After this step, we got a list of code smells, consisting of the project name, package name, class name, and method name where each detected code smell occurs.
-It is important to note that the code smell detection results from DesigniteJava may not be entirely accurate, as the detection results can contain both false positives and false negatives.
-False positives occur when the tool incorrectly flags code as exhibiting CC or CM, while false negatives happen when actual code smells are missed.
-These inaccuracies stem from the rule-based and heuristic nature of DesigniteJava, which relies on predefined criteria that may not capture all nuances of the code.
+DesigniteJava is an open-source tool capable of detecting 18 design smells and 10 implementation smells, whereas PMD is a rule-based static analysis tool that identifies common programming issues, including code smells.
+Both of them have been extensively used in prior studies (sharma2021qscored, ; shah2023mining, ; barbosa2020revealing, ; fernandes2016review, ; aniche2018code, ; wu2024ismell, ).
+In this paper, we leveraged DesigniteJava to detect potential instances of CC, CM, and FE, and used PMD to detect potential instances of DC.
+The output of this step is a list of detected code smells, each annotated with the corresponding project, package, class, and, where applicable, method name.
+It is worth noting that the code smell detection results from DesigniteJava and PMD may not be entirely accurate, as the detection results can contain both false positives and false negatives.
+False positives occur when the tool incorrectly flags code as exhibiting certain type of code smell, while false negatives happen when actual code smells are missed.
+These inaccuracies stem from the rule-based and heuristic nature of DesigniteJava and PMD, which relies on predefined criteria that may not capture all nuances of the code.
 3.3.4. Data Preparation
-CodeSplitJava (codesplitjava, ) is a utility program designed to partition methods from Java source code into distinct files.
-We used CodeSplitJava to parse the 294 Java repositories collected in Section 3.3.2 into code files.
-Each of these code files corresponds to a method in the 294 repositories, and its path reflects the hierarchical structure of the method (i.e., namespaces and packages are converted into folders).
-By correlating the list of detected code smells collected in Section 3.3.3, with the paths of the code files generated by CodeSplitJava, we identified the method-level code snippets with code smells and those without (i.e., Clean Code).
-We removed any duplicate samples within each type of code smell to ensure uniqueness.
-Additionally, we limited the maximum token length of the samples to 1,024, filtering out any that exceeded this limit.
-Table 1 provides the number of potential positive samples after each step for the two types of code smells.
-In total, we collected 4,266 potential positive code samples containing CC and 6,479 containing CM.
-For both CC and CM, we introduced an equal number of potential negative samples (i.e., code samples without code smells as identified by DesigniteJava) for the subsequent steps.
-We also verified that all negative samples adhered to the maximum token limit and were free of duplicates.
-However, the quality of these datasets may be compromised as they could contain false positives and false negatives, requiring further refinement.
+CodeSplitJava (codesplitjava, ) is a utility program designed to extract methods or classes from Java source code and organize them into separate files.
+We employed CodeSplitJava to parse the 401 Java repositories collected in Section 3.3.2, generating individual code files for each method or class.
+The file paths reflect the hierarchical structure of the original code (i.e., namespaces and packages are converted into nested folders).
+By aligning the list of detected code smells obtained in Section 3.3.3 with the paths of the code files generated by CodeSplitJava, we were able to collect code snippets that were labeled by heuristic-based detectors as either containing code smells or being free of them.
 Table 1.
@@
 Code Smell #Initial Samples #After Deduplication #After Token Limit
-Complex Conditional 19,754 4,589 4,266
-Complex Method 14,979 9,781 6,479
-3.3.5. Manual Validation
-After the above steps, we obtained an equal number of potential positive and negative samples for both CC ($4,266\times 2=8,532$) and CM ($6,479\times 2=12,958$).
-To ensure the quality of the datasets for these two types of code smells and to exclude false positives and false negatives, a manual validation was conducted by the first and fourth authors.
-Initially, a pilot validation was performed at the 95% confidence level, with a margin of error of less than 5%, resulting in the random selection of 368 samples from CC and 374 from CM.
-The first and the fourth author annotated the selected samples independently, and the pilot annotations were subsequently compared.
-Cohen’s Kappa coefficient (jacob1960coefficient, ) was used to measure the inter-rater agreement for both CC and CM.
-The resulting Kappa coefficient for CC was 0.85, while for CM it was 0.81, both exceeding 0.7, indicating a strong level of agreement between the two annotators.
-Following the pilot validation, the two authors manually annotated the remaining code samples.
-Throughout the entire validation process, including the pilot phase, any disagreements between the first and fourth authors were resolved through discussions involving the second author, ensuring consistency in the annotations.
-Overall, we identified 2,708 false positives and 29 false negatives for CC, and 2,376 false positives and 20 false negatives for CM.
-The numbers of both positive and negative samples for CC and CM after manual validation are presented in Table 2.
-Table 2. Number of positive and negative samples after data annotation
+Complex Conditional 23,984 7,768 6,901
+Complex Method 20,171 14,687 9,819
+Feature Envy 2,152 1,113 337
+Data Class 9,672 8,501 7,121
+We removed any duplicate samples within each type of code smell to ensure uniqueness.
+Additionally, we limited the maximum token length of the samples to 1,024, filtering out any that exceeded this limit.
+This decision was made to accommodate the input constraints of the models used in our experiments (see Section 3.6).
+Specifically, the input window size for the LMs was set to 1,024 tokens to align with GPU memory limitations.
+Samples exceeding 1,024 tokens undergo truncation, which may result in the loss of critical contextual information.
+For instance, key sections of code relevant to the detection of code smells could be omitted, potentially impacting the model’s performance.
+Therefore, we set a hard limit of 1,024 tokens to avoid such issues, ensuring that all samples in our dataset can be fully utilized by the models without sacrificing the integrity of their content.
+Table 1 provides the number of potential positive samples after each step for the four types of code smells.
+Note that although FE is a method-level code smell, we retained the entire class that contains the method as the sample.
+This decision aligns with the definition of FE (see Section 3.3.1), which emphasizes the method’s interaction with its enclosing class.
+Providing the full class context ensures that the code sample contains sufficient structural and semantic information for FE detection.
+3.3.5. Manual Review
+After the above steps, we obtained two sets of samples for each code smell: a potential positive set and a potential negative set, with the latter being significantly larger in size.
+To construct a high-quality and balanced dataset, the first and the fourth authors carried out a two-stage manual review process, as illustrated in Figure 4.
+In the first stage, we reviewed all potential positive samples manually.
+Each sample was carefully examined to identify and eliminate false positives.
+If a sample in the potential positive set had been incorrectly labeled as positive by DesigniteJava or PMD, it was reclassified as negative and moved to the negative set; otherwise, it was transferred into the positive set.
+This stage yielded a verified positive set and a smaller verified negative set.
+In the second stage, in order to enlarge the negative set to achieve dataset balance, we randomly selected samples from the potential negative set for manual validation.
+Similar to the first stage, samples confirmed as true negatives were added to the negative set, while those identified as false negatives were reclassified into the positive set.
+This process was repeated iteratively until an equal number of manually reviewed positive and negative samples was achieved.
+During this stage, all randomly selected potential negative samples were required to be unique, adhere to the predefined maximum token limit, and be representative.
+To ensure the representativeness of negative samples across the four types of code smells, we established the following criteria: for CC, negative samples must contain at least one conditional statement; for CM, negative samples must contain at least 15 lines of code; and for FE and DC, negative samples must contain at least 30 lines of code, as we extract the corresponding class for these two smells.
+These constraints help filter out trivial or uninformative code snippets and ensure that the negative samples are sufficiently complex and realistic, thereby contributing to a more robust and reliable dataset.
+Through this two-stage review, we ensured that the final dataset for each code smell type was both balanced and of high quality.
+Figure 4. The process of two-stage manual review in dataset construction
+Table 2. Number of positive and negative samples after manual review
 Code Smell #Positive #Negative Total
-Complex Conditional 1,587 6,945 8,532
-Complex Method 4,123 8,835 12,958
+Complex Conditional 4,604 4,604 9,208
+Complex Method 5,748 5,748 11,496
+Feature Envy 307 307 614
+Data Class 6,922 6,922 13,844
+In order to ensure the consistency of data annotation, a pilot review was performed by the first and the fourth authors before conducting the formal manual review process.
+For each code smell type (i.e., CC, CM, FE, and DC), 30 potential positive samples and 30 potential negative samples labeled by heuristic-based tools were randomly selected.
+The two authors independently examined these samples and then compared their labeling results.
+Cohen’s Kappa coefficient (jacob1960coefficient, ) was used to assess inter-rater agreement between the two authors.
+The Kappa coefficient for CC was 0.83, for CM was 0.80, for FE was 0.77, and for DC was 0.87.
+All values exceeded the threshold of 0.7, indicating a strong level of agreement between the annotators.
+Following the pilot review, the two authors held a discussion to reconcile discrepancies and formalize the labeling criteria below:
+  • CC: Labeled as positive if the method contains at least one conditional statement that either (i) has a nesting depth of three or more, making the method hard to read, or (ii) includes compound boolean expressions involving three or more operands, exhibiting complex logic that makes the condition challenging to evaluate and understand.
+  • CM: Labeled as positive if the method simultaneously satisfies both (i) three or more execution paths with complex control flow, and (ii) an implementation that is difficult to comprehend or follow.
+  • FE: Labeled as positive if the method inside the input class satisfies both (i) over 50% of its data access operations, such as method calls and field accesses, target members of external classes rather than its own class, and (ii) relocating the method to one of these external classes would improve encapsulation and cohesion.
+  • DC: Labeled as positive if the class contains (i) primarily fields accompanied by trivial getters/setters (at least 80% of its fields), and (ii) lacks methods that implement significant logic or domain behavior, indicating its role as a passive data holder.
+These criteria were grounded in established definitions from prior studies (sharma2017house, ; fowler1999refactoring, ; fontana2016comparing, ; pietrzak2006leveraging, ) and refined based on insights from our pilot study.
+During the formal manual review, the two authors collaboratively labeled all samples (i.e., each sample was assigned to one of the two authors), which took approximately 300 person-hours in total.
+Whenever questions or ambiguities arose, the second author, who is an expert in code smell detection, was consulted.
+In such cases, all three authors discussed the sample until consensus was reached, ensuring annotation reliability and consistency.
+The final count of positive and negative samples, following the manual review, is presented in Table 2.
 3.3.6. Data Split
-After manually reviewing the entire dataset, we split the code samples for both CC and CM into training, validation, and test sets following an 8:1:1 ratio, as employed in previous studies (yang2024federated, ; liu2023benchmarking, ).
-Specifically, the positive and negative samples for each method-level code smell were distributed according to the same 8:1:1 proportion.
-This approach ensures a consistent distribution of positive and negative cases, providing a solid foundation for model training, validation, and evaluation.
-The number of code samples in the datasets of CC and CM after this step is presented in Table 3.
-We used these clean and high-quality datasets as our final datasets for a series of experiments.
+In this step, we split the code samples of CC, CM, FE, and DC into training, validation, and test sets following an 8:1:1 ratio, as employed in previous studies (yang2024federated, ; liu2023benchmarking, ).
+For each type of code smell, we ensured that the number of positive and negative samples remains equal across the training, validation, and test sets, thereby maintaining class balance throughout the entire dataset.
+Table 3 outlines the number of code samples for each code smell type after the data split.
+We then used these clean and high-quality datasets as our final datasets for a series of experiments.
 Table 3. Number of positive and negative samples after data split
-Set Complex Conditional Complex Method
-#Positive #Negative #Total #Positive #Negative #Total
-Train 1,269 5,557 6,826 3,299 7,067 10,366
-Valid 159 694 853 412 884 1,296
-Test 159 694 853 412 884 1,296
-Total 1,587 6,945 8,532 4,123 8,835 12,958
+Code Smell Train Valid Test Total
+Complex Conditional #Positive 3,684 460 460 4,604
+#Negative 3,684 460 460 4,604
+Complex Method #Positive 4,600 574 574 5,748
+#Negative 4,600 574 574 5,748
+Feature Envy #Positive 245 31 31 307
+#Negative 245 31 31 307
+Data Class #Positive 5,538 692 692 6,922
+#Negative 5,538 692 692 6,922
 3.4. Base Model Selection
-In this work, we aim to explore the effectiveness of PEFT methods on method-level code smell detection task.
-To ensure a thorough analysis, we have chosen to use both small LMs and LLMs as our base models.
-Note that we deem models with 1B or more parameters as LLMs, while referring to those with fewer parameters as small language models.
-First, the selected LMs should be open-source as PEFT methods require access to their parameters.
-Second, the models should support classification tasks as we frame code smell detection as a binary classification problem.
-Third, the base models used in this study must be able to be fine-tuned and tested on a single 48GB GPU (our maximum computational resource) without encountering memory overflow.
-Fourth, the selected models should either be popular small LMs or recent advanced LLMs to ensure they represent state-of-the-art capabilities.
-Based on the above criteria, we employed 4 small LMs (i.e., CodeBERT (feng2020codebert, ), GraphCodeBERT (guo2021graphcodebert, ), UnixCoder (guo2022unixcoder, ), and CodeT5 (wang2021codet5, )) and 6 LLMs (DeepSeek-Coder-1.3B (guo2024deepseekcoder, ), DeepSeek-Coder-6.7B (guo2024deepseekcoder, ), StarCoderBase-1B (li2023starcoder, ), StarCoderBase-3B (li2023starcoder, ), StarCoderBase-7B (li2023starcoder, ), and CodeLlama-7B (roziere2024codellama, )) for our experiments.
-These models feature a range of transformer architectures, including encoder-only, encoder-decoder, and decoder-only modes.
-All the models we selected have demonstrated encouraging results across various code understanding tasks like code clone detection, defect detection, etc.
-(liu2024vul, ; karmakar2021what, ; lu2021codexglue, ; li2022automating, ; tang2023csgvd, ; zeng2022an, ; chen2023diversevul, ; wang2023one, ; fu2022vulrepair, ; nguyen2024empirical, ; su2024distilled, ).
-Table 4 presents the details of the 10 base models, including their parameter sizes and architecture.
+In this work, we aim to explore the effectiveness of PEFT methods for the task of code smell detection.
+To ensure a thorough analysis, we included both small and large LMs as our base models.
+Consistent with the definition adopted by Weyssow et al.
+(weyssow2023exploring, ), we classified models with 1B or more parameters as LLMs, and those with fewer parameters as SLMs.
 Table 4. Selected LMs and their corresponding parameter size and architecture
 Model #Params Architecture
-Small LMs CodeBERT 125M Encoder-only
+SLMs CodeBERT 125M Encoder-only
 GraphCodeBERT 125M
@@
 CodeT5 220M Encoder-decoder
-LLMs DeepSeek-Coder-1.3B 1.3B Decoder-only
+LLMs StarCoderBase-1B 1.1B Decoder-only
+StarCoderBase-3B 3B
+DeepSeek-Coder-1.3B 1.3B
 DeepSeek-Coder-6.7B 6.7B
-StarCoderBase-1B 1.1B
-StarCoderBase-3B 3B
-StarCoderBase-7B 7B
 CodeLlama-7B 7B
+First, the selected LMs should be open-source as PEFT methods require access to their parameters.
+Second, the models should support classification tasks as we frame code smell detection as a binary classification problem.
+Third, the base models used in this study must be able to be fine-tuned and tested on a single 40GB GPU (our maximum computational resource) without encountering memory overflow.
+Fourth, the selected models should either be popular SLMs or recent advanced LLMs to ensure they represent state-of-the-art capabilities.
+Based on the above criteria, we employed 4 SLMs (i.e., CodeBERT (feng2020codebert, ), GraphCodeBERT (guo2021graphcodebert, ), UnixCoder (guo2022unixcoder, ), and CodeT5 (wang2021codet5, )) and 5 LLMs (StarCoderBase-1B (li2023starcoder, ), StarCoderBase-3B (li2023starcoder, ), DeepSeek-Coder-1.3B (guo2024deepseekcoder, ), DeepSeek-Coder-6.7B (guo2024deepseekcoder, ), and CodeLlama-7B (roziere2024codellama, )) for our experiments.
+To ensure a comprehensive analysis, we deliberately included models representing all three major Transformer architectures: encoder-only, encoder-decoder, and decoder-only.
+Although decoder-only models were originally designed for generative tasks (vaswani2017attention, ; radford2019language, ), prior studies have shown their competitive performance in sequence classification tasks as well (wolf2020transformers, ; kariyappa2024progressive, ).
+Therefore, we also included decoder-only models as base models in our study.
+Additionally, all the models we selected have demonstrated encouraging results across various code understanding tasks like code clone detection, defect detection, etc.
+(liu2024vul, ; karmakar2021what, ; lu2021codexglue, ; li2022automating, ; tang2023csgvd, ; zeng2022an, ; chen2023diversevul, ; wang2023one, ; fu2022vulrepair, ; nguyen2024empirical, ; su2024distilled, ).
+Table 4 presents the details of the 9 base models, including their parameter sizes and architectural types.
 3.5. Metrics
@@
 Specifically, we utilize precision, recall, F1-score, and MCC (Matthews Correlation Coefficient) for assessing the performance of detecting CC and CM.
-We include MCC because, unlike other metrics like F1-score, accounts for true negative instances (yao2020assessing, ).
+We include MCC because, unlike other metrics like F1-score, MCC accounts for true negative instances (yao2020assessing, ).
 These metrics are calculated using the following formulas:
@@
 The code for our experiments is implemented by the PEFT library [^4] https://huggingface.co/docs/peft/index from Hugging Face.
-After reviewing prior studies (liu2023empirical, ; liu2024delving, ; li2024comprehensive, ; wang2023one, ; weyssow2023exploring, ) and observing the results of pilot experiments, we set the learning rate to $3e^{-4}$ for all PEFT methods and $1e^{-5}$ for full fine-tuning the small LMs.
+After reviewing prior studies (liu2023empirical, ; liu2024delving, ; li2024comprehensive, ; wang2023one, ; weyssow2023exploring, ) and observing the results of pilot experiments, we set the learning rate to $3e^{-4}$ for all PEFT methods and $1e^{-5}$ for full fine-tuning the SLMs.
 Throughout all the fine-tuning experiments, we utilized the AdamW optimizer (loshchilov2017decoupled, ) and fixed the Transformer input window size at 1,024.
+The choice to limit the input length to 1,024 tokens was driven by hardware constraints.
+As the GPU memory consumption of transformer-based models grows roughly quadratically with input length (tay2022efficient, ), this setting enables all selected base models to train within available GPU memory capacity.
 All the models were trained for 10 epochs, and we selected the epochs at which models exhibit minimum loss on the validation sets for inference.
-We executed all models on a single Linux GPU server equipped with a 48G NVIDIA RTX 6000 Ada Generation GPU.
-Additionally, we have released our datasets and source code online for replication purpose (replicationPackage, ) where more details, including hyper-parameter values, can be found.
+Table 5 shows the hyperparameter settings of different PEFT methods.
+We adopted a batch size of 1 for all experiments to avoid GPU out-of-memory issues.
+We used a fixed random seed (42) across all runs to ensure experimental reproducibility.
+For prompt tuning and prefix tuning, we used 20 virtual tokens, which are learnable vectors prepended to the input and optimized during training.
+This setting is widely used in the literature and offers a balance between representational capacity and computational cost (li2021prefix, ; weyssow2023exploring, ).
+As for LoRA, we set the rank $r$ to 8, the scaling factor $\alpha$ to 16, and the dropout rate to 0.1 to prevent overfitting, following configurations reported in (liu2024delving, ).
+For both LoRA and ${IA}^{3}$, we applied parameter-efficient adaptation to the core components of the Transformer architecture, including the query ($Q$), key ($K$), and value ($V$) projections in the attention layers.
+In the case of $(\text{IA})^{3}$, we also included the Feed-Forward Network (FFN) for adaptation (see Section 2.3).
+We executed all models on a single Linux GPU server equipped with a 40GB NVIDIA A100 GPU.
+Additionally, we have released our datasets and source code online for replication purpose (replicationPackage, ) where more details can be found.
+Table 5. Hyperparameter settings of different PEFT methods and full fine-tuning
+PEFT Method Hyperparameter Setting Batch Size Seed
+Prompt Tuning Number of virtual tokens: 20 1 42
+Prefix Tuning
+LoRA Rank $r$: 8
+$alpha$: 16
+$dropout$: 0.1
+Target modules: $Q$, $K$, and $V$
+$(IA)^{3}$ Target modules: $Q$, $K$, $V$, and FFN
+Full Fine-tuing None
 4. Results
 In this section, we first illustrate the experiment setup for each RQ formulated in Section 3, and we then present the results and findings for the four RQs.
-4.1. RQ1: What is the effectiveness of different PEFT methods and full fine-tuning on various small LMs and LLMs?
+4.1. RQ1: What is the effectiveness of different PEFT methods and full fine-tuning on various SLMs and LLMs?
+4.1.1. Experiment Setup
 Since our objective is to demonstrate the effectiveness of PEFT methods on both small and large LMs for code smell detection, we conducted a series of experiments involving four PEFT methods (i.e., prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$) and a group of models with different parameter sizes.
@@
 Note that we only applied LoRA and $(IA)^{3}$ to Code-T5 as the current PEFT library does not support fine-tuning encoder-decoder models with prompt tuning and prefix tuning.
-We full fine-tuned all the small models due to resource constraints, as the computational demands of tuning all parameters in LLMs would surpass the available 48GB of GPU memory.
-Additionally, we reported and analyzed the computing resource consumption of these fine-tuning methods, and the results are shown in Figures 4, 5, and 6.
-Table 5 and Table 6 present the comprehensive comparison of small LMs and LLMs tuning with various PEFT techniques for detecting CC and CM, respectively.
-Table 5. Results of small LMs and LLMs with different PEFT methods and full fine-tuning on detecting CC (bold: highest values per model, underline: highest values among all models)
-Model Method Precision Recall F1 MCC
-Small LMs CodeBERT Full FT 90.68% 50.00% 44.86% 0.00%
-Prompt Tuning 72.81% 69.55% 70.92% 42.23%
-Prefix Tuning 75.82% 73.23% 74.38% 48.98%
-LoRA 80.57% 69.03% 72.49% 48.24%
-$(IA)^{3}$ 81.22% 68.86% 72.45% 48.53%
-GraphCodeBERT Full FT 70.45% 75.65% 72.25% 45.80%
-Prompt Tuning 73.88% 71.85% 72.77% 45.69%
-Prefix Tuning 81.05% 73.00% 75.95% 53.44%
-LoRA 84.72% 69.67% 73.82% 52.27%
-$(IA)^{3}$ 83.33% 74.30% 77.58% 56.91%
-UnixCoder Full FT 67.67% 60.01% 61.69% 26.60%
-Prompt Tuning 73.93% 70.64% 72.03% 44.44%
-Prefix Tuning 78.73% 73.53% 75.64% 52.00%
-LoRA 82.12% 69.00% 72.74% 49.40%
-$(IA)^{3}$ 74.33% 69.26% 71.21% 43.29%
-CodeT5 Full FT 76.94% 78.09% 77.49% 55.02%
-LoRA 63.08% 67.65% 46.82% 30.38%
-$(IA)^{3}$ 62.42% 67.82% 63.11% 29.76%
-LLMs DeepSeek-Coder-1.3B Prompt Tuning 75.25% 70.27% 72.23% 45.25%
-Prefix Tuning 75.07% 75.68% 75.37% 50.75%
-LoRA 78.35% 70.51% 73.28% 48.22%
-$(IA)^{3}$ 75.74% 70.66% 72.65% 46.11%
-DeepSeek-Coder-6.7B Prompt Tuning 74.96% 68.91% 71.13% 43.46%
-Prefix Tuning 79.36% 70.72% 73.70% 49.33%
-LoRA 75.55% 71.63% 73.26% 47.02%
-$(IA)^{3}$ 75.79% 67.29% 69.99% 42.23%
-StarCoderBase-1B Prompt Tuning 78.44% 71.06% 73.74% 48.95%
-Prefix Tuning 76.84% 70.95% 73.20% 47.42%
-LoRA 83.53% 54.82% 54.21% 25.41%
-$(IA)^{3}$ 80.02% 70.31% 73.51% 49.39%
-StarCoderBase-3B Prompt Tuning 73.67% 71.78% 72.64% 45.41%
-Prefix Tuning 74.18% 70.47% 72.01% 44.49%
-LoRA 69.39% 54.89% 54.72% 19.49%
-$(IA)^{3}$ 79.58% 69.12% 72.39% 47.57%
-StarCoderBase-7B Prompt Tuning 73.31% 68.97% 70.69% 42.05%
-Prefix Tuning 74.73% 72.62% 73.58% 47.30%
-LoRA 69.67% 65.46% 67.03% 34.88%
-$(IA)^{3}$ 75.27% 76.48% 75.84% 51.73%
-CodeLlama-7B Prompt Tuning 74.10% 71.92% 72.90% 45.97%
-Prefix Tuning 71.88% 71.13% 71.49% 43.00%
-LoRA 76.57% 73.20% 74.66% 49.66%
-$(IA)^{3}$ 71.26% 71.57% 71.41% 42.83%
-Table 6. Results of small LMs and LLMs with different PEFT methods and full fine-tuning on detecting CM (bold: highest values per model, underline: highest values among all models)
-Model Method Precision Recall F1 MCC
-Small LMs CodeBERT Full FT 78.55% 53.84% 48.59% 20.95%
-Prompt Tuning 84.14% 88.18% 85.20% 72.21%
-Prefix Tuning 85.98% 88.42% 86.95% 74.36%
-LoRA 85.07% 86.07% 85.53% 71.13%
-$(IA)^{3}$ 86.17% 87.32% 86.70% 73.48%
-GraphCodeBERT Full FT 78.34% 82.63% 77.01% 60.81%
-Prompt Tuning 84.75% 87.83% 85.84% 72.52%
-Prefix Tuning 85.54% 86.79% 86.11% 72.33%
-LoRA 84.92% 86.97% 85.77% 71.86%
-$(IA)^{3}$ 84.15% 86.40% 85.05% 70.51%
-UnixCoder Full FT 76.24% 74.36% 75.13% 50.57%
-Prompt Tuning 84.64% 87.26% 85.64% 71.85%
-Prefix Tuning 86.89% 88.79% 87.70% 75.66%
-LoRA 86.99% 90.66% 88.20% 77.56%
-$(IA)^{3}$ 86.26% 87.82% 86.95% 74.07%
-CodeT5 Full FT 85.66% 89.44% 86.83% 75.00%
-LoRA 78.08% 52.26% 45.44% 15.92%
-$(IA)^{3}$ 42.00% 43.21% 42.36% -14.73%
-LLMs DeepSeek-Coder-1.3B Prompt Tuning 85.83% 88.05% 86.74% 73.85%
-Prefix Tuning 86.40% 89.81% 87.57% 76.13%
-LoRA 85.14% 87.79% 86.15% 72.88%
-$(IA)^{3}$ 85.23% 87.59% 86.17% 72.78%
-DeepSeek-Coder-6.7B Prompt Tuning 86.22% 88.92% 87.26% 75.08%
-Prefix Tuning 85.73% 89.96% 86.86% 75.57%
-LoRA 84.98% 87.48% 85.95% 72.42%
-$(IA)^{3}$ 85.58% 83.14% 84.18% 68.68%
-StarCoderBase-1B Prompt Tuning 85.68% 87.74% 86.54% 73.39%
-Prefix Tuning 84.82% 87.05% 85.72% 71.83%
-LoRA 84.81% 87.57% 85.84% 72.33%
-$(IA)^{3}$ 86.30% 89.30% 87.40% 75.53%
-StarCoderBase-3B Prompt Tuning 84.53% 86.24% 85.26% 70.74%
-Prefix Tuning 85.10% 86.38% 85.68% 71.48%
-LoRA 84.83% 89.44% 85.77% 74.13%
-$(IA)^{3}$ 86.46% 88.38% 87.28% 74.81%
-StarCoderBase-7B Prompt Tuning 83.64% 88.07% 84.53% 71.57%
-Prefix Tuning 85.35% 87.26% 86.16% 72.59%
-LoRA 85.64% 88.45% 86.70% 74.04%
-$(IA)^{3}$ 85.72% 88.38% 86.75% 74.05%
-CodeLlama-7B Prompt Tuning 74.05% 87.68% 86.00% 72.60%
-Prefix Tuning 84.10% 84.36% 84.23% 68.46%
-LoRA 85.39% 88.29% 86.46% 73.62%
-$(IA)^{3}$ 84.07% 88.40% 85.05% 72.34%
-4.1.1. Small LMs
-Generally, GraphCodeBERT is the most effective in identifying CC, UnixCoder excels at detecting CM, and CodeT5 with PEFT methods performs the worst in classifying both types of method-level code smells.
-For CC and CM, CodeBERT demonstrates comparable results to GraphCodeBERT.
-For example, the F1-score and MCC for CodeBERT with prompt tuning in detecting CM are 85.20% and 72.21%, respectively, while the F1-score and MCC for GraphCodeBERT with prompt tuning for CM are 85.84% and 72.52%.
-These results are consistent with previous studies (arshad2022codebert, ; rabbani2022comparative, ; lu2021codexglue, ) that CodeBERT and GraphCodeBERT are powerful in code semantics learning, showing outstanding performance in Java code clone detection.
-GraphCodeBERT is built upon CodeBERT and additionally considers the inherent structure of code (i.e., data flow) in the pre-training stage (guo2021graphcodebert, ).
-Therefore, it is reasonable that GraphCodeBERT manifests a bit stronger capability than CodeBERT in code smell detection.
-UnixCoder also shows similar performance to GraphCodeBERT for CC, while it stands out as the best small model for CM, achieving an F1-score of 88.20% and an MCC of 77.56% with LoRA.
-For all the small LMs, CodeT5 with PEFT methods significantly lags behind the other three models with PEFT methods for both CC and CM, with MCC values well below 50%, and even a negative value when tuned with $(IA)^{3}$ for CM.
-This suggests that CodeT5 with PEFT methods struggles to perform effectively in code smell detection compared to the other models, highlighting the limitations of its PEFT adaptation.
-In contrast, fully fine-tuned CodeT5 outperforms all other fully fine-tuned small models in detecting both CC and CM.
+We full fine-tuned all the small models due to resource constraints, as the computational demands of tuning all parameters in LLMs would surpass the available 40GB of GPU memory.
+Table 6 and Table 7 present the comprehensive comparison of small LMs and LLMs tuning with various PEFT techniques for detecting CC, CM, FE, and DC, respectively.
+Table 6. Results of small LMs and LLMs with different PEFT methods on detecting CC and CM (bold: highest values per model, underline: highest values among all models)
+Model Method CC CM
+F1 MCC F1 MCC
+SLMs CodeBERT Full FT 70.54% 44.41% 80.57% 61.19%
+Prompt Tuning 62.38% 25.90% 76.13% 54.00%
+Prefix Tuning 63.88% 27.87% 81.33% 62.94%
+LoRA 73.04% 46.10% 71.00% 49.79%
+$(IA)^{3}$ 69.44% 40.49% 81.32% 62.96%
+GraphCodeBERT Full FT 74.01% 48.08% 80.66% 61.33%
+Prompt Tuning 65.24% 30.81% 78.90% 58.60%
+Prefix Tuning 69.82% 39.97% 79.18% 60.46%
+LoRA 73.15% 46.30% 80.03% 60.93%
+$(IA)^{3}$ 72.50% 45.00% 80.00% 60.45%
+UnixCoder Full FT 73.04% 46.10% 82.39% 64.89%
+Prompt Tuning 69.35% 38.70% 78.66% 57.95%
+Prefix Tuning 71.63% 43.26% 82.71% 65.84%
+LoRA 72.19% 46.04% 73.18% 51.52%
+$(IA)^{3}$ 73.47% 46.98% 82.23% 64.47%
+CodeT5 Full FT 73.01% 46.19% 81.13% 62.74%
+LoRA 33.33% 0.00% 83.29% -2.95%
+$(IA)^{3}$ 43.38% -4.05% 40.32% 16.70%
+LLMs StarCoderBase-1B Prompt Tuning 71.07% 42.21% 80.05% 60.11%
+Prefix Tuning 70.10% 40.23% 76.88% 54.01%
+LoRA 41.41% -1.62% 45.41% 7.01%
+$(IA)^{3}$ 74.27% 49.01% 80.90% 62.01%
+StarCoderBase-3B Prompt Tuning 69.12% 38.29% 81.87% 63.85%
+Prefix Tuning 67.81% 35.69% 78.25% 56.92%
+LoRA 40.13% 7.93% 68.41% 46.00%
+$(IA)^{3}$ 74.68% 49.96% 82.07% 64.81%
+Deepseek-Coder-1.3B Prompt Tuning 70.21% 40.44% 81.77% 63.75%
+Prefix Tuning 69.38% 39.62% 79.06% 58.36%
+LoRA 69.44% 40.49% 78.42% 57.96%
+$(IA)^{3}$ 72.50% 45.01% 81.87% 63.82%
+Deepseek-Coder-6.7B Prompt Tuning 67.93% 35.87% 81.19% 63.11%
+Prefix Tuning 68.80% 37.62% 80.09% 60.57%
+LoRA 69.98% 40.05% 81.34% 63.63%
+$(IA)^{3}$ 70.84% 42.36% 83.34% 66.88%
+CodeLlama-7B Prompt Tuning 68.68% 37.92% 81.25% 63.43%
+Prefix Tuning 53.41% 15.95% 80.77% 62.08%
+LoRA 70.12% 40.70% 80.70% 62.59%
+$(IA)^{3}$ 70.48% 42.32% 83.00% 66.17%
+Table 7. Results of small LMs and LLMs with different PEFT methods on detecting FE and DC (bold: highest values per model, underline: highest values among all models)
+Model Method FE DC
+F1 MCC F1 MCC
+SLMs CodeBERT Full FT 74.09% 48.80% 81.98% 65.60%
+Prompt Tuning 56.17% 13.07% 81.56% 64.57%
+Prefix Tuning 62.82% 25.93% 81.81% 65.43%
+LoRA 67.44% 36.17% 83.17% 68.48%
+$(IA)^{3}$ 66.49% 38.48% 83.30% 68.18%
+GraphCodeBERT Full FT 77.40% 54.95% 82.66% 67.42%
+Prompt Tuning 66.88% 37.49% 83.75% 68.95%
+Prefix Tuning 74.19% 48.39% 89.26% 79.26%
+LoRA 72.23% 46.36% 85.85% 72.40%
+$(IA)^{3}$ 69.84% 45.48% 85.58% 72.56%
+UnixCoder Full FT 77.04% 56.76% 83.51% 69.49%
+Prompt Tuning 77.32% 55.30% 83.68% 68.74%
+Prefix Tuning 75.80% 51.64% 83.89% 69.28%
+LoRA 70.97% 41.94% 83.60% 68.65%
+$(IA)^{3}$ 73.95% 49.32% 83.96% 69.44%
+CodeT5 Full FT 66.49% 38.48% 83.46% 69.12%
+LoRA 54.07% 10.02% 70.03% 49.06%
+$(IA)^{3}$ 54.65% 9.76% 46.79% -2.66%
+LLMs StarCoderBase-1B Prompt Tuning 72.23% 46.36% 84.66% 71.21%
+Prefix Tuning 75.65% 52.30% 85.44% 72.19%
+LoRA 78.58% 60.68% 79.30% 58.85%
+$(IA)^{3}$ 73.95% 49.32% 84.96% 71.71%
+StarCoderBase-3B Prompt Tuning 68.96% 39.74% 83.63% 68.47%
+Prefix Tuning 67.61% 35.78% 86.56% 74.06%
+LoRA 70.85% 42.29% 83.70% 68.64%
+$(IA)^{3}$ 69.35% 38.73% 84.90% 71.46%
+DeepSeek-Coder-1.3B Prompt Tuning 78.98% 58.34% 84.17% 69.99%
+Prefix Tuning 66.88% 37.49% 82.44% 64.90%
+LoRA 78.98% 58.34% 82.72% 67.68%
+$(IA)^{3}$ 83.87% 67.74% 85.84% 72.57%
+DeepSeek-Coder-6.7B Prompt Tuning 78.89% 58.83% 85.56% 72.72%
+Prefix Tuning 80.56% 61.81% 83.03% 68.14%
+LoRA 77.32% 55.30% 83.43% 68.64%
+$(IA)^{3}$ 77.32% 55.30% 85.34% 72.34%
+CodeLlama-7B Prompt Tuning 72.52% 45.37% 85.49% 72.55%
+Prefix Tuning 63.57% 30.67% 82.30% 64.61%
+LoRA 77.32% 55.30% 83.10% 68.31%
+$(IA)^{3}$ 74.19% 48.39% 83.54% 68.45%
+4.1.2. SLMs
+Overall, GraphCodeBERT proved the most effective in identifying CC and DC, while UnixCoder excelled at detecting CM and FE.
+In comparison, CodeT5 fine-tuned with PEFT methods performed the worst across all types of code smells.
+CodeBERT performed similarly to GraphCodeBERT, albeit slightly worse.
+For example, in CC detection, whether using full fine-tuning or PEFT methods, CodeBERT consistently achieved lower F1-scores and MCC values compared to GraphCodeBERT, with differences ranging from 0.11% to 5.93% and 0.21% to 12.10%, respectively.
+This is reasonable because GraphCodeBERT is built upon CodeBERT and further incorporates code structure information (i.e., data flow) during the pre-training phase (guo2021graphcodebert, ).
+Across all SLMs, CodeT5 with PEFT methods significantly lags behind the other three models with PEFT methods in the code smell detection task, which requires both method- and class-level code understanding.
+When fine-tuned with PEFT methods, the MCC values of CodeBERT even became negative, reaching -4.05% for LoRA on CC and -2.66% for $(IA)^{3}$ on DC.
+This suggests that CodeT5 with PEFT methods struggles to perform effectively in code smell detection, highlighting the limitations of its PEFT adaptation.
+In contrast, the fully fine-tuned CodeT5 demonstrated acceptable performance in code smell detection.
 The stark performance gap between CodeT5 with full fine-tuning and PEFT methods could stem from CodeT5’s architecture, which may depend more on full fine-tuning for optimal adaptation to the method-level code smell detection task.
@@
 DeepSeek-Coder-1.3B stands out as the top-performing LLM, consistently achieving the highest or on-par F1-scores and MCC values across all fine-tuning methods when compared to other LLMs.
-While StarCoderBase-7B with $(IA)^{3}$ gains higher F1-score and MCC value than DeepSeek-Coder-1.3B in detecting CC, it significantly underperforms when fine-tuned with LoRA.
-Specifically, the F1-score and MCC for DeepSeek-Coder-1.3B with LoRA are 73.28% and 48.22%, respectively, compared to 67.03% and 34.88% for StarCoderBase-7B.
+Although the StarCoderBase models perform similarly to the DeepSeek-Coder models in detecting FE and DC, which involve class-level code understanding, their results are notably less stable for detecting CC and CM, which require method-level understanding.
+In particular, the performance of StarCoderBase models fine-tuned with LoRA degrades significantly on the detection of CC and CM.
 Notably, when fine-tuned with PEFT techniques, models with larger parameter sizes within the same family do not surpass, and in some cases, even slightly underperform their smaller counterparts.
-For example, when detecting CC, DeepSeekCoder-1.3B gains higher F1-scores and MCC values than DeepSeek-Coder-6.7B across the four PEFT methods, and StarCoderBase-1B with prompt tuning and prefix tuning also outperforms StarCoderBase-7B.
-This observation aligns with Fan et al.’s research (fan2024exploring, ), which shows that larger models do not always have better performance on code-related tasks.
-Figure 4 illustrates the peak GPU memory consumption for the three largest models — DeepSeek-Coder-6.7B, StarCoderBase-7B, and CodeLlama-7B — using different PEFT techniques for code smell detection.
-It is evident that, for both CodeLlama-7B and DeepSeek-Coder-7B, $(IA)^{3}$ requires the highest computational resources, with LoRA ranking second.
-This observation is partly consistent with Li et al.’s work (li2024comprehensive, ) that $(IA)^{3}$ tends to exhibit the highest memory usage.
-However, Li et al.
-(li2024comprehensive, ) also found that LoRA consumes less memory than prefix tuning, which is not the case in our results.
-Unlike DeepSeek-Coder-6.7B and CodeLlama-7B, StarCoderBase-7B shows that LoRA demands more GPU memory than the other PEFT techniques.
-Figure 4. Peak GPU memory of DeepSeek-Coder-6.7B, StarCoderBase-7B, and CodeLlama-7B with different PEFT methods on detecting CC and CM
-4.1.3. Small LMs vs. LLMs
-From Table 5 and Table 6, it can be observed that the best-performing models for detecting CC and CM are GraphCodeBERT with $(IA)^{3}$ and UnixCoder with LoRA, both of which are small LMs.
-More specifically, small LMs with PEFT techniques, with the exception of CodeT5, generally outperform or achieve comparable performance to LLMs.
-While Weyssow et al.’s work (weyssow2023exploring, ) highlights that LLMs significantly overtake small LMs in code generation tasks, our findings suggest that, for the binary classification task which identifies method-level code smells, small LMs with fewer number of trainable parameters excel beyond LLMs.
-This discrepancy might arise because the large capacity of LLMs introduces unnecessary complexity, which can obscure the essential patterns needed for the code smell detection task.
-In contrast, small LMs appear to possess an optimal capacity for capturing these crucial patterns without being encumbered by excessive complexity (tay2023scaling, ).
-Another possible explanation for the discrepancy is that small LMs, being encoder-only models, are better suited to classification tasks like code smell detection.
-Encoder-only models are typically more efficient at tasks that require pattern recognition, such as binary classification, while LLMs are often optimized for generative tasks.
-This indicates that the advantages of LLMs observed in code generation tasks do not necessarily translate to all types of software engineering tasks, and larger models are not always superior when fine-tuned for different downstream applications.
+For example, in FE detection, StarCoderBase-1B and DeepSeek-Coder-1.3B generally achieve higher F1-scores and MCC values than StarCoderBase-3B and DeepSeek-Coder-6.7B.
+4.1.4. SLMs vs. LLMs
+For the detection of CC and CM, SLMs and LLMs demonstrate generally comparable performance.
+Notably, in the case of DC, GraphCodeBERT, a representative SLM, significantly outperforms all other models, including LLMs.
+These results suggest that smaller models can remain competitive, or even superior, for certain types of code smells.
+In contrast, for FE, LLMs fine-tuned with PEFT methods substantially outperform SLMs, likely due to the more complex and semantic nature of this smell type, which benefits from the richer contextual representations and broader knowledge captured by larger models.
+Figure 5. Peak GPU memory usage for detecting CC: (a) SLMs with PEFT methods and Full Fine-Tuning; (b) LLMs with PEFT methods
+Additionally, Figure 5 summarizes the peak GPU memory usage of both SLMs and LLMs with different PEFT methods on CC detection.
+Comparing Figure 5 (a) and (b), we can observe that PEFT methods impose substantially higher peak GPU memory requirements on LLMs compared to SLMs, particularly for larger-sized LLMs.
+4.1.5. Method-level vs. Class-level Code Understanding
+The detection of CC and CM requires method-level code understanding, whereas identifying FE and DC necessitates class-level analysis (see Section 3.3.1).
+By comparing the detection results of CC and CM with those of FE and DC, we observe that LMs demonstrate comparable effectiveness at both levels of code understanding.
+Notably, the models even achieve better performance on class-level code smells, as evidenced by the fact that the best-performing model yields significantly lower results on CC compared to the other smell types.
 4.1.6. Best PEFT Method
-Overall, for both small LMs and LLMs in the DeepSeek-Coder family, prefix tuning emerges as the most effective technique for method-level code smell detection among the four PEFT methods investigated in this study.
-In contrast, for the StarCoderBase family, $(IA)^{3}$ tends to be the most effective technique.
-Meanwhile, for CodeLlama-7B, LoRA proves to be the best PEFT method.
-For small models, except for CodeT5, as well as LLMs like DeepSeek-Coder-1.3B and DeepSeek-Coder-6.7B, prefix tuning consistently yields the highest F1-scores and MCC values when detecting CC and CM, demonstrating its generally strong effectiveness in fine-tuning these models for code smell detection.
-In the StarCoderBase family, LLMs fine-tuned with $(IA)^{3}$ outperform those fine-tuned with other PEFT techniques.
-For example, when detecting CM, StarCoderBase-1B, StarCoderBase-3B, and StarCoderBase-7B all achieve their highest F1-scores and MCC values using $(IA)^{3}$.
-However, this does not apply to CodeLlama-7B, where using $(IA)^{3}$ results in the worst performance for detecting CC and the second-worst performance for detecting CM.
-On the other hand, CodeLlama-7B always achieves its best results on this task with LoRA.
-These findings contrast with those of previous studies (li2024comprehensive, ; weyssow2023exploring, ), which suggested that a single PEFT method could be the best across all models for a specific downstream task.
-For method-level code smell detection, the optimal PEFT method varies depending on the model.
-Notably, prompt tuning is the least effective PEFT method for adapting LMs to detect CC and CM, as most models tuned with prefix tuning show worse performance compared to those tuned with other PEFT techniques.
-Figure 5. Peak GPU memory of small LMs with different PEFT methods and full fine-tuning on detecting CC
-Figure 6. Peak GPU memory of small LMs with different PEFT methods and full fine-tuning on detecting CM
+Generally, for SLMs, both $(IA)^{3}$ and prefix tuning demonstrate strong performance across different models.
+For LLMs, $(IA)^{3}$ consistently delivers the best results on the detection of all four types of code smells.
+To be specific, the most effective PEFT method for SLMs on the task of code smell detection varies depending on the model: for instance, CodeBERT performs optimally with $(IA)^{3}$, whereas UnixCoder tends to achieve better results with prefix tuning.
+However, for LLMs, $(IA)^{3}$ always outperforms other PEFT methods across all studied code smells.
+Whether for StarCoderBase Models, Deepseek-Coder Models, or CodeLlama-7B, they all achieve superior performance when fine-tuned using $(IA)^{3}$.
+Additionally, for LLMs, prompt tuning can be employed as an alternative option.
+Notably, despite achieving promising results on certain models and code smell types, LoRA exhibits instability in performance and can yield markedly poor effectiveness in some cases.
+An example of this is using LoRA on StarCoderBase-1B for CM detection.
+A case in point is the use of LoRA on StarCoderBase-1B for detecting CM.
+While StarCoderBase-1B achieves a minimum MCC of 54.01% with the other three PEFT methods (i.e., prompt tuning, prefix tuning, and $(IA)^{3}$), its performance drops sharply to 7.01% with LoRA, resulting in a disparity of over 40%.
 4.1.7. PEFT methods vs. Full Fine-tuning
-According to Table 5 and Table 6, we can find that all the four PEFT methods achieve better or at least comparable effectiveness than full fine-tuning on most small models for method-level code smell detection.
-For small LMs, excluding CodeT5, the F1-scores and MCC values of prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$ all show varying degrees of improvement over full fine-tuning.
-Our results align with Liu et al.’s prior study (liu2023empirical, ), confirming that PEFT methods can match or surpass the performance of standard full fine-tuning for code understanding tasks while adjusting fewer number of trainable parameters.
-One possible reason for this could be that full fine-tuning adjusts all of the model’s parameters, which may lead to overfitting, especially when the training data is limited.
-In contrast, PEFT methods, by focusing on specific parts of the model, allow for more efficient learning without the risk of overfitting to the task at hand.
+According to Table 6 and Table 7, we can find that all the four PEFT methods achieve better or at least comparable effectiveness than full fine-tuning on most small models for code smell detection.
+For small LMs (excluding CodeT5), prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$ generally achieve varying degrees of improvement in F1-scores and MCC values compared to full fine-tuning for most code smell types.
+However, for FE, all PEFT methods perform significantly worse than full fine-tuning.
+These results suggest that, in most cases, PEFT methods can match or surpass the performance of standard full fine-tuning for code smell detection while adjusting fewer number of trainable parameters.
 PEFT methods focus on enhancing the effectiveness of LMs to levels similar to full fine-tuning, while aiming to minimize the use of computational resources, such as peak GPU memory consumption.
-Figure 5 and Figure 6 depict the peak GPU memory usage and corresponding MCC values of fine-tuning the four small models.
-It can be observed that prefix tuning and prompt tuning require significantly less GPU memory than other fine-tuning methods, yet they can still achieve considerable effectiveness.
-Notably, prefix tuning attains the highest or second-highest MCC value while using minimal GPU memory, proving that prefix tuning is the most efficient PEFT method for small models on code smell detection.
-When juxtaposed with full fine-tuning, all of the four PEFT methods lead to a substantial reduction in peak GPU memory usage while achieving better or similar results.
-This further validates the efficiency of PEFT methods for code smell detection relative to full fine-tuning.
-Answer to RQ1: Small LMs, with the exception of CodeT5, generally achieve better effectiveness than LLMs when using PEFT methods for code smell detection.
-The most effective PEFT method varies across different models.
-All PEFT methods outperform or match the performance of full fine-tuning while requiring less peak GPU memory, highlighting the efficiency of PEFT techniques compared to full fine-tuning for detecting CC and CM.
+Taking CC detection as an example, Figure 5 (a) depicts the peak GPU memory consumption during the fine-tuning of the four SLMs.
+It can be observed that juxtaposed with full fine-tuning, all of the four PEFT methods lead to a substantial reduction in peak GPU memory usage, which further validates the efficiency of PEFT methods for code smell detection relative to full fine-tuning.
+Answer to RQ1: Overall, LLMs perform slightly better than SLMs in code smell detection, except for DC.
+Among different PEFT methods, both $(IA)^{3}$ and prefix tuning are most effective for SLMs, while $(IA)^{3}$ consistently yields the best results for LLMs.
+All PEFT techniques outperform or match the performance of full fine-tuning for several code smell types, while also reducing peak GPU memory usage, demonstrating their efficiency.
+However, for the detection of FE, full fine-tuning remains more effective.
 4.2. RQ2: What is the effectiveness of PEFT methods with different hyper-parameters?
-Given the restrictions of our computational resources and popularity of LoRA, we focused on a single PEFT method, LoRA, to investigate how small LMs and LLMs perform under different hyper-parameter settings.
-Specifically, we used LoRA to tune a small LM (GraphCodeBERT) and an LLM (DeepSeek-Coder-1.3B), adapting the value of LoRA rank $r$ in the range of {16, 32, 64, 640, 1,280, 2,560}.
+4.2.1. Experiment Setup
+Given the restrictions of our computational resources and popularity of LoRA, we focused on a single PEFT method, LoRA, to investigate how SLMs and LLMs perform under different hyper-parameter settings.
+Specifically, we used LoRA to tune a SLM (GraphCodeBERT) and an LLM (DeepSeek-Coder-1.3B), adapting the value of LoRA rank $r$ in the range of {8, 32, 64, 640, 1,280, 2,560}.
 At the same time, we kept other hyper-parameters of LoRA constant, with the LoRA alpha set to 16 and the LoRA dropout set to 0.1.
 The experiment results are presented in Figure 6.
-Figure 7. F1-scores and MCC of GraphCodeBERT and DeepSeek-Coder-1.3B using LoRA on detecting CC and CM with different LoRA rank values
-The results in Figure 7 show that both CodeBERT and DeepSeek-Coder-1.3B achieve their best performance in detecting CC when fine-tuned with LoRA at an $r$ value of 32.
-For CM, GraphCodeBERT performs best at a LoRA $r$ value of 32, while DeepSeek-Coder-1.3B reaches its peak performance at $r=64$.
-As the $r$ value increases from 16 to 2,560, the effectiveness of code smell detection initially increases, then decreases, and finally rises again at $r=2,560$, but it does not exceed the level at $r=16$.
-This pattern can be explained by the increase in trainable parameters as $r$ rises, which initially enhances the performance by allowing the models to learn more complex data patterns.
-However, beyond a certain threshold, the challenge of tuning these additional parameters effectively, especially with relatively limited amount of data, leads to diminishing returns.
-Without pre-training for these extra parameters, the models are prone to overfitting, where they may learn noise or irrelevant patterns, limiting their generalization capability.
-Consequently, as the number of trainable parameters grows beyond what the available training data can support, the models’ performance may plateau or even decline.
-At $r=2,560$, the models appear to utilize the expanded parameter space somewhat more effectively, leading to a slight performance recovery.
-However, this does not surpass the performance at $r=16$, suggesting that the added parameters introduce more complexity than the data can justify.
-These results highlight that increasing the number of trainable parameters does not always lead to linear performance gains.
+Figure 6. MCC results of GraphCodeBERT and DeepSeek-Coder-1.3B using LoRA for code smell detection under different rank settings: (a) CC; (b) CM; (c) FE; (d) DC
+4.2.2. Results of RQ2
+The results in Figure 6 show that for both GraphCodeBERT and DeepSeek-Coder-1.3B, the optimal $r$ value varies across different code smell types, indicating that the effectiveness of LMs on code smell detection is sensitive to the choice of LoRA configuration.
+Since the LoRA $r$ value directly controls the number of trainable parameters, this observation implies that different types of code smells may require different levels of model adaptability.
+Furthermore, increasing the LoRA $r$ value from 8 to 2,560, which corresponds to an increase in the number of trainable parameters, does not uniformly lead to improved performance.
+Although a higher $r$ value increases the number of trainable parameters, which provides the model with greater capacity to learn, the MCC scores for detecting the four code smell types do not show a monotonically increasing trend.
+These findings indicate that expanding the number of trainable parameters alone does not guarantee better effectiveness, emphasizing the importance of careful tuning of the LoRA configuration rather than relying solely on scaling the model.
 Answer to RQ2: Increasing the number of trainable parameters by adjusting hyper-parameters does not necessarily ensure effectiveness enhancement for code smell detection task.
-In fact, the effectiveness often declines as the number of trainable parameters continues to increase.
+Additionally, the effectiveness of LMs in detecting different code smells is sensitive to the configuration of PEFT methods (e.g., LoRA).
 4.3. RQ3: What is the effectiveness of PEFT methods in low-resource scenarios?
-To simulate the limited availability of data in low-resource scenarios, following previous works (wang2023one, ; liu2023empirical, ), we randomly selected 100, 200, 500, and 1,000 samples from our datasets while maintaining the proportions of the original data distribution.
-In this RQ, we employed the four PEFT methods and full fine-tuning on a small LM (GraphCodeBERT) and an LLM (DeepSeek-Coder-1.3B) to conduct the experiments.
+4.3.1. Experiment Setup
+To simulate low-resource scenarios, following previous studies (wang2023one, ; liu2023empirical, ), we randomly selected 50, 100, 250, and 500 samples for CC, CM, and DC, respectively, based on the size of our dataset.
+Note that each subset is balanced, containing an equal number of positive and negative samples.
+We excluded FE from this RQ, as the total number of available FE samples in our training set is already below 500 (see Section 3.3.6), making it inherently a low-resource task.
 The experiment results are shown in Table 8.
 Table 8.
-Results of GrahphCodeBERT and DeepSeek-Coder-1.3B with different PEFT methods and full fine-tuning on detecting CC and CM in low-resource scenarios (bold: highest values per scenario for a model, underline: highest values among all scenarios for a model)
-Model #Training Samples Method CC CM
-F1 MCC F1 MCC
-GraphCodeBERT 100 Full FT 44.86% 0.00% 83.87% 69.52%
-Prompt Tuning 47.18% 6.84% 80.25% 61.85%
-Prefix Tuning 63.94% 32.85% 76.33% 52.67%
-LoRA 57.34% 25.28% 82.44% 64.96%
-$(IA)^{3}$ 69.63% 39.47% 81.32% 65.84%
-200 Full FT 52.48% 21.18% 86.26% 73.93%
-Prompt Tuning 67.05% 34.12% 83.26% 67.26%
-Prefix Tuning 72.47% 46.00% 81.15% 62.49%
-LoRA 66.61% 41.44% 85.20% 71.60%
-$(IA)^{3}$ 71.92% 44.12% 84.05% 68.31%
-500 Full FT 67.78% 43.34% 86.69% 74.50%
-Prompt Tuning 62.44% 27.96% 80.20% 60.40%
-Prefix Tuning 69.92% 41.74% 83.26% 66.72%
-LoRA 71.08% 45.30% 84.92% 69.84%
-$(IA)^{3}$ 67.87% 39.54% 84.25% 69.17%
-1,000 Full FT 75.22% 50.81% 85.49% 73.83%
-Prompt Tuning 69.63% 39.74% 82.58% 65.30%
-Prefix Tuning 73.44% 47.02% 84.16% 68.84%
-LoRA 74.76% 49.54% 86.67% 74.16%
-$(IA)^{3}$ 63.95% 35.41% 84.07% 68.13%
-DeepSeekCoder-1.3B 100 Full FT 44.86% 0.00% 81.48% 68.21%
-Prompt Tuning 47.43% 14.34% 78.45% 56.93%
-Prefix Tuning 59.82% 20.70% 82.35% 67.39%
-LoRA 54.90% 16.98% 83.48% 67.47%
-$(IA)^{3}$ 50.93% 8.63% 74.02% 48.52%
-200 Full FT 72.76% 50.97% 85.25% 72.17%
-Prompt Tuning 60.80% 30.66% 77.63% 55.49%
-Prefix Tuning 69.20% 39.65% 82.52% 67.05%
-LoRA 66.65% 36.72% 83.18% 66.49%
-$(IA)^{3}$ 64.97% 32.56% 77.91% 55.91%
-500 Full FT 66.34% 40.34% 83.32% 69.66%
-Prompt Tuning 62.13% 30.85% 80.12% 60.25%
-Prefix Tuning 71.96% 44.14% 83.91% 68.70%
-LoRA 72.94% 45.89% 83.79% 69.70%
-$(IA)^{3}$ 63.07% 34.91% 79.17% 58.36%
-1,000 Full FT 72.80% 45.63% 82.50% 65.25%
-Prompt Tuning 70.24% 42.31% 82.04% 64.10%
-Prefix Tuning 70.26% 42.97% 84.43% 69.58%
-LoRA 71.84% 44.27% 86.13% 73.00%
-$(IA)^{3}$ 73.33% 46.68% 80.33% 60.65%
-According to the results reported in Table 7, it is evident that when the data is limited to 100 or 200 samples, the effectiveness of both GraphCodeBERT and DeepSeek-Coder-1.3B with PEFT methods in detecting method-level code smells drops dramatically.
+Results of GrahphCodeBERT and DeepSeek-Coder-1.3B with different PEFT methods and full fine-tuning on detecting CC, CM, and DC in low-resource scenarios (bold: highest values per scenario for a model, underline: highest values among all scenarios for a model)
+Model #Training Samples Method CC CM DC
+GraphCodeBERT 50 Full FT 19.23% 54.61% 62.76%
+Prompt Tuning -0.90% 38.19% 59.51%
+Prefix Tuning 6.47% 31.64% 50.79%
+LoRA 7.29% 44.99% 62.51%
+$(IA)^{3}$ 11.84% 45.15% 61.40%
+100 Full FT 24.67% 51.58% 68.28%
+Prompt Tuning 3.91% 39.03% 61.49%
+Prefix Tuning 15.32% 38.87% 54.19%
+LoRA 23.14% 46.98% 68.71%
+$(IA)^{3}$ 23.66% 43.99% 65.62%
+250 Full FT 28.38% 55.36% 64.93%
+Prompt Tuning 13.05% 42.03% 64.69%
+Prefix Tuning 20.02% 47.02% 62.19%
+LoRA 31.11% 51.91% 66.79%
+$(IA)^{3}$ 28.49% 52.45% 64.05%
+500 Full FT 34.85% 57.03% 67.64%
+Prompt Tuning 18.87% 47.50% 63.41%
+Prefix Tuning 25.87% 53.05% 64.63%
+LoRA 36.52% 58.31% 67.83%
+$(IA)^{3}$ 34.14% 56.22% 68.58%
+DeepSeek-Coder-1.3B 50 Full FT 11.38% 42.46% 68.08%
+Prompt Tuning 5.88% 25.99% 48.34%
+Prefix Tuning -3.70% 43.16% 41.93%
+LoRA -1.48% 36.74% 71.84%
+$(IA)^{3}$ 1.54% 17.95% 50.29%
+100 Full FT 12.27% 24.94% 74.57%
+Prompt Tuning 13.05% 33.00% 52.83%
+Prefix Tuning 0.68% 45.19% 52.24%
+LoRA 10.96% 46.15% 74.01%
+$(IA)^{3}$ 6.55% 20.21% 59.53%
+250 Full FT 26.69% 48.02% 74.61%
+Prompt Tuning 22.18% 38.16% 60.06%
+Prefix Tuning 14.87% 45.18% 61.13%
+LoRA 20.53% 55.47% 75.33%
+$(IA)^{3}$ 15.28% 35.25% 65.16%
+500 Full FT 29.33% 58.89% 69.70%
+Prompt Tuning 22.09% 40.15% 65.99%
+Prefix Tuning 17.72% 46.98% 60.89%
+LoRA 32.03% 59.06% 69.57%
+$(IA)^{3}$ 9.86% 43.50% 70.78%
+4.3.2. Results of RQ3
+According to the results reported in Table 8, it is evident that when the data is limited to 50 samples, the effectiveness of both GraphCodeBERT and DeepSeek-Coder-1.3B with PEFT methods in detecting code smells drops, with CC experiencing the most pronounced performance degradation.
 This significant decline in performance highlights the challenges small and large LMs face in low-resource scenarios.
-However, when the number of training samples of CC or CM increases to 1,000, both models show improved performance compared to when only 100 samples are available.
+However, when the number of training samples of CC, CM, or DC increases to 500, both models show improved performance compared to when only 50 samples are available.
 This improvement underscores that a larger dataset enables both small and large LMs using PEFT methods to achieve substantially better performance in code smell detection.
-Moreover, with 1,000 samples, GraphCodeBERT and DeepSeekCoder-1.3B with PEFT methods are able to yield reasonably good results compared to training with the original datasets for both types of method-level code smells.
-For example, the F1-scores of GraphCodeBERT training with 1,000 samples only degrade by no more than 3.26%, and the MCC decrease by no more than 7.22% compared to training with sufficient data for detecting CM, despite the dataset size being reduced by 90.35% ((10,366 - 1,000) / 10,366).
-This highlights the benefit of PEFT in resource-limited settings, as it enables the reduction of not only training parameters but also the amount of required training data, while still maintaining strong effectiveness on code smell detection.
-Notably, when the number of training samples increases to 1,000, GraphCodeBERT and DeepSeek-Coder-1.3B with PEFT methods demonstrate effectiveness in detecting CC and CM that is on par with, or even exceeds, that of full fine-tuning, further highlighting the superiority of PEFT methods in low-resource scenarios.
-Secondly, compared to the results in Table 5 and Table 6, we can find that GraphCodeBERT with full fine-tuning performs better in low-resource scenarios than when there is sufficient data.
-This improvement can be attributed to the fact that, with ample data, full fine-tuning may lead to overfitting, causing the model to memorize rather than generalize the data.
-In contrast, in low-resource scenarios, the model is forced to learn more generalizable patterns, thus improving its performance.
-Finally, in low-resource scenarios, LoRA tends to be the most effective PEFT method.
-As indicated by the results of RQ1 (see Section 4.1.4), for both GraphCodeBERT and DeepSeek-Coder-1.3B, fine-tuning with prefix tuning enables them to achieve their best performance compared to the other three PEFT methods.
-However, when data is scarce, GraphCodeBERT and DeepSeek-Coder-1.3B often attain their highest F1-scores or MCC values through LoRA.
+Moreover, with 500 samples, GraphCodeBERT and DeepSeekCoder-1.3B with PEFT methods are able to yield reasonably good results compared to training with the original datasets for code smell detection.
+When the number of training samples reaches 250 or 500, both models with PEFT techniques are able to outperform their fully fine-tuned counterparts.
+Additionally, in low-resource scenarios, LoRA tends to be the most effective PEFT method.
+As indicated by the results of RQ1 (see Section 4.1.6), for GraphCodeBERT, prefix tuning and $(IA)^{3}$ deliver superior performance, while for DeepSeek-Coder-1.3B, $(IA)^{3}$ consistently yields the best performance across different code smells.
+However, when data is scarce, GraphCodeBERT and DeepSeek-Coder-1.3B often attain their highest MCC values through LoRA.
 It suggests that under different resource settings, the most effective PEFT method may vary.
-Answer to RQ3: In low-resource scenarios, the effectiveness of PEFT methods for code smell detection diminishes considerably when data is limited to 100 or 200 samples.
-However, as the number of training samples increases, the performance of PEFT methods shows significant improvement, and in some cases, even surpasses that of full fine-tuning.
+Answer to RQ3: In low-resource scenarios, the effectiveness of PEFT methods for code smell detection diminishes considerably when data is limited to 50 samples.
+However, as the number of training samples increases, the performance of PEFT methods shows significant improvement.
+Moreover, when the number of training samples reaches 250 or 500, several PEFT techniques are able to outperform full fine-tuning, demonstrating their potential even in moderately low-resource settings.
+Table 9. Prompt 1 for GPT-4o-mini and DeepSeek-v3 in detecting CC, CM, FE, and DC
+Code Smell Prompt Template*
+CC You are a professional Java programmer. Given a piece of Java code, determine whether it contains “Complex Conditional” code smell. If it does, return only “Yes”, else return “No”.
+Definition of “Complex Conditional”: This code smell occurs when a conditional statement is excessively intricate.
+### Input: [Code Sample]
+### Output: [Yes/No]
+CM You are a professional Java programmer. Given a piece of Java code, determine whether it contains “Complex Method” code smell. If it does, return only “Yes”, else return “No”.
+Definition of “Complex Method”: This code smell occurs when a method exhibits high cyclomatic complexity.
+### Input: [Code Sample]
+### Output: [Yes/No]
+FE You are a professional Java programmer. Given a piece of Java code, determine whether it contains “Feature Envy” code smell. If it does, return only “Yes”, else return “No”.
+Definition of “Feature Envy”: This code smell occurs when a method is more interested in a class other than the one it actually is in.
+### Input: [Code Sample]
+### Output: [Yes/No]
+DC You are a professional Java programmer. Given a piece of Java code, determine whether it contains “Data Class” code smell. If it does, return only “Yes”, else return “No”.
+Definition of “Data Class”: This code smell occurs when a class contains only fields and basic methods for accessing them.
+### Input: [Code Sample]
+### Output: [Yes/No]
+  • * Prompt 1 shown in this table was curated by the authors. Prompt 2 and Prompt 3, which were generated by ChatGPT, are available in our replication package (replicationPackage, ).
 4.4. RQ4: What is the effectiveness of PEFT methods compared to state-of-the-art code smell detection approaches?
-Prompt Template for Detecting Complex Conditional with GPT-3.5-turbo You are a professional Java programmer.
-Given a piece of Java code, determine whether it contains “Complex Conditional” code smell.
-If it does, return only “Yes”, else return “No”.
-Definition of “Complex Conditional”: This code smell occurs when a conditional statement is excessively intricate.
-### Input: [Code Sample] ### Output: [Yes/No]
-Prompt Template for Detecting Complex Method with GPT-3.5-turbo You are a professional Java programmer.
-Given a piece of Java code, determine whether it contains “Complex Method” code smell.
-If it does, return only “Yes”, else return “No”.
-Definition of “Complex Method”: This code smell occurs when a method exhibits high cyclomatic complexity.
-### Input: [Code Sample] ### Output: [Yes/No]
-Traditional heuristic-based tools for code smell detection have been widely used in academia (sharma2021qscored, ; shah2023mining, ; barbosa2020revealing, ), while DL has recently gained widespread adoption in research on code smell detection, achieving impressive performance (zhang2024code, ).
-In this RQ, we established both heuristic-based (i.e., DesigniteJava) and best-in-class DL detectors (i.e., DeepSmells (ho2023fusion, ) and AE-Dense (sharma2021code, )), as well as a general-purpose LLM (i.e., GPT-3.5-turobo [^3] https://platform.openai.com/docs/models/o1) which has not been specifically tuned for method-level code smell detection, as baselines.
-To obtain valid responses from GPT-3.5-turbo in a zero-shot setting (kojima2024large, ), we designed the above prompt templates that include both instructions and definitions of CC or CM to guide ChatGPT in annotating the test sets.
-GraphCodeBERT with $(IA)^{3}$ for CC and UnixCoder with LoRA for CM are selected for comparison.
-The results of this evaluation are shown in Table 8.
-From the results reported in Table 8, we can find that GraphCodeBERT with $(IA)^{3}$ and UnixCoder with LoRA exhibit significantly better performance than the state-of-the-art code smell detection approaches and GPT-3.5-turbo in detecting CC and CM, respectively.
-Specifically, for CC, GraphCodeBERT with $(IA)^{3}$ significantly surpasses the effectiveness of the other four baseline methods with an F1-score disparity ranging from 23.41% to 47.94% and an MCC difference ranging from 13.24% to 50.59%.
-For CM, UnixCoder with LoRA achieves an F1-score improvement ranging from 11.22% to 54.32% and an MCC improvement ranging from 10.28% to 72.91%.
-Table 8. Results of GraphCodeBERT with $(IA)^{3}$ and UnixCoder with LoRA compared to state-of-the-art code smell detection approaches on detecting CC and CM (bold: highest values per code smell)
+4.4.1. Experiment Setup
+Figure 7. Results of DeepSeek-v3 and GPT-4o-mini on detecting CC, CM, FE, and DC under ICL
+In this RQ, we established both heuristic-based (i.e., DesigniteJava and PMD) and best-in-class DL detectors (i.e., DeepSmells (ho2023fusion, ) and AE-Dense (sharma2021code, )) as baselines for comparison.
+Additionally, In-Context Learning (ICL) (tom2020language, ) has demonstrated impressive performance in various SE tasks, such as code generation (gao2024what, ; jia2024acecoder, ).
+To evaluate its potential in code smell detection, we explored the effectiveness of two recent and powerful general-purpose LLMs, GPT-4o-mini (hurst2024gpt4o, ) and DeepSeek-v3 (liu2024deepseek, ), under both zero-shot and few-shot settings.
+We selected these two models because they represent the latest advancements in commercial and open-weight general-purpose LLMs, respectively, and have demonstrated strong performance across a range of natural language and programming-related tasks (rasheed2024task, ; bruni2025benchmarking, ; liu2024deepseek, ; jahin2025unveiling, ).
+Specifically, we first designed a prompt template (i.e., prompt 1) for each of the four types of code smells examined in this study, incorporating both instructional text and smell definitions, as shown in Table 9.
+We then used ChatGPT (ChatGPT, ) to generate two additional variants for each template (i.e., prompt 2 and prompt 3), resulting in three distinct prompts per code smell type.
+For the zero-shot setting, these prompts were directly provided to GPT-4o-mini and DeepSeek-v3 without any labeled examples.
+In the few-shot setting, example pairs were randomly sampled from the training set, where each pair consisted of one code sample with a code smell and one without, along with their corresponding labels.
+These examples were combined with the base prompts to form few-shot configurations containing one, two, or three example pairs.
+Due to resource constraints (e.g., limited API usage budget), we restricted the few-shot setting to include no more than three example pairs per prompt.
+As a result, each code smell type yields a total of 12 unique prompt-example configurations, derived from three prompt variants combined with four levels of contextual information (zero-shot, plus one-, two-, and three-pair few-shot settings).
+This ICL setup ensures diversity in both prompt formulations and example composition, allowing for a more comprehensive evaluation of general-purpose LLMs and enabling a fair comparison with our proposed PEFT-based methods in the context of code smell detection.
+Figure 7 presents the performance of GPT-4o-mini and DeepSeek-v3 on code smell detection in both zero- and few-shot settings.
+Table 10 summarizes the final results for RQ4, where we report only the best-performing configurations for both GPT-4o-mini and DeepSeek-v3 under ICL across all prompts and example variants.
+4.4.2. ICL
+It can be observed from Figure 7 that DeepSeek-v3 generally demonstrates better effectiveness in code smell detection compared to GPT-4o-mini.
+Additionally, both models achieve better performance under few-shot settings than in the zero-shot setting.
+However, their performance does not consistently improve as the number of example pairs increases.
+4.4.3. Comparison with State-of-the-Art Approaches
+Table 10. Results of PEFT methods compared to state-of-the-art code smell detection approaches on detecting CC, CM, FE, and DC (bold: highest values per code smell)
 Code Smell Method Precision Recall F1 MCC
-CC DesigniteJava 34.58% 98.74% 51.22% 43.67%
-DeepSmells 21.61% 47.17% 29.64% 6.32%
-AE-Dense 42.90% 53.82% 47.74% 15.20%
-GPT-3.5-turbo 42.86% 73.58% 54.17% 42.66%
-GraphCodeBERT with $(IA)^{3}$ 83.33% 74.30% 77.58% 56.91%
-CM DesigniteJava 61.88% 99.27% 76.23% 65.92%
-DeepSmells 33.82% 56.80% 42.39% 4.65%
-AE-Dense 24.33% 55.76% 33.88% 9.78%
-GPT-3.5-turbo 81.35% 73.06% 76.98% 67.28%
-UnixCoder with LoRA 86.99% 90.66% 88.20% 77.56%
-Answer to RQ4: In comparison with existing code smell detection approaches, LMs tuned with PEFT methods exhibit notable performance improvements on identifying CC and CM, highlighting the effectiveness of PEFT methods for code smell detection.
+CC DesigniteJava 64.01% 76.96% 69.89% 34.41%
+DeepSmells 51.06% 100.00% 67.61% 6.63%
+AE-Dense 51.13% 49.10% 50.10% 2.92%
+GPT-4o-mini 71.48% 64.46% 61.29% 35.24%
+DeepSeek-v3 72.38% 71.74% 71.53% 44.12%
+StarCoderBase-3B with $(IA)^{3}$ 75.18% 74.78% 74.68% 49.96%
+CM DesigniteJava 64.69% 98.61% 78.12% 52.58%
+DeepSmells 63.64% 79.87% 70.83% 33.35%
+AE-Dense 50.54% 52.05% 51.29% 3.05%
+GPT-4o-mini 79.06% 67.25% 63.54% 44.78%
+DeepSeek-v3 79.74% 73.78% 72.40% 53.19%
+DeepSeek-Coder-6.7B with $(IA)^{3}$ 83.52% 83.36% 83.34% 66.88%
+FE DesigniteJava 73.91% 54.84% 62.96% 36.73%
+DeepSmells 97.10% 51.54% 67.34% 7.48%
+AE-Dense 50.82% 100.00% 67.39% 12.80%
+GPT-4o-mini 77.53% 77.42% 77.40% 54.95%
+DeepSeek-v3 74.29% 74.19% 74.17% 48.49%
+DeepSeek-Coder-1.3B with $(IA)^{3}$ 83.87% 83.87% 83.87% 67.74%
+DC PMD 76.29% 89.74% 82.47% 62.83%
+DeepSmells 87.72% 67.68% 76.41% 49.55%
+AE-Dense 52.79% 94.36% 67.70% 16.18%
+GPT-4o-mini 84.87% 82.51% 82.21% 67.35%
+DeepSeek-v3 87.57% 87.57% 87.57% 75.15%
+GraphCodeBERT with prefix tuning 90.95% 90.03% 89.97% 79.26%
+Based on the results presented in Table 10, we can find that LMs fine-tuned with PEFT techniques consistently demonstrate the highest effectiveness among all existing methods for code smell detection.
+Additionally, among conventional baseline approaches, DeepSeek-v3 achieves the most competitive overall performance.
+While it shows a significant performance gap compared to LMs tuned with PEFT methods for CM and FE detection (with MCC differences exceeding 10%), its performance in CC and DC detection remains relatively strong, with only a 4%-5% reduction in MCC values, demonstrating its relative strength in these specific code smell types.
+Furthermore, our results indicate that traditional DL-based methods (DeepSmells and AE-Dense) exhibit substantially inferior performance compared to both PEFT and the other baseline approaches.
+This finding underscores the limitations of conventional DL models in code smell detection, particularly when compared to more advanced techniques like PEFT techniques and ICL.
+Answer to RQ4: In comparison with existing code smell detection approaches, LMs tuned with PEFT methods reveal notable performance improvements, highlighting the effectiveness of PEFT methods for code smell detection.
 5. Discussion
@@
 However, the effectiveness of existing methods for detecting this type of code smell is considerably limited.
-As shown in Table 8, existing methods — whether heuristic-based (e.g., DesigniteJava), DL-based (e.g., DeepSmells and AE-Dense), or GPT-3.5-turbo — demonstrate limited success in detecting CC, with all MCC values below 45%, and some even as low as 6.32%.
-Although our proposed method, GraphCodeBERT with $(IA)^{3}$, improved the performance of CC code smell detection significantly, reaching an F1-score of 77.58% and an MCC value of 56.91%, it still falls short of expectations due to challenges such as the considerable existence of false positives and negatives, which may limit its practical applicability.
+As shown in Table 10, existing methods, including heuristic-based (e.g., DesigniteJava), DL-based (e.g., DeepSmells) and AE-Dense), and general-purpose LLMs under ICL settings (e.g., GPT-4o-mini and DeepSeek-v3), all demonstrate limited success in detecting CC, with all MCC values below 50%, and some even as low as 6.63%.
+Although our proposed method, StarCoderBase-3B with $(IA)^{3}$, improved the performance of CC code smell detection significantly, reaching an F1-score of 74.68% and a MCC value of 49.96%, it still falls short of expectations due to challenges such as the considerable existence of false positives and negatives, which may limit its practical applicability.
 Accurate identification of CC code smells can help developers better refactor the code, ultimately improving the maintainability of the software system.
 Despite the progress made, the detection of CC remains a challenging task that requires further improvements.
-Model size is not the decisive factor contributing to the effectiveness of code smell detection: Previous research has demonstrated that larger models with more parameters tend to deliver better performance on code-related tasks (nijkamp2022codegen, ; roziere2024codellama, ; weyssow2023exploring, ).
-For example, Weyssow et al.
-(weyssow2023exploring, ) suggested a potential shift in the software engineering domain, moving away from the long-established practice of fine-tuning smaller LMs like CodeT5 over the past 3-4 years, and towards embracing larger LLMs.
-However, the results of RQ1 indicate that, for the method-level code smell detection task, model size is not the sole determinant of performance.
-In fact, all the small models, with the exception of CodeT5, generally perform as well as or better than larger LLMs in detecting CC and CM (see Section 4.1.3).
-Furthermore, within the same model family, larger models do not always outperform its smaller-sized counterparts (see Section 4.1.2).
-In contrast, the alignment between the model and the specific task at hand seems to play a crucial role in determining effectiveness for code smell detection.
-While CodeT5, when fine-tuned with PEFT methods, has demonstrated strong performance in tasks such as code summarization and code change prediction (liu2023empirical, ; fan2024exploring, ), it surprisingly underperforms in the task of code smell detection.
-This suggests that models may exhibit varying performance across different downstream tasks, emphasizing the importance of aligning a model’s design and fine-tuning with the specific requirements of each target task.
-Therefore, it is essential to focus on selecting models that are well-suited to the method-level code smell detection task, rather than relying solely on model size as a performance indicator.
-Considering the model used and the availability of resources when selecting the appropriate PEFT method for code smell detection: Based on the results of RQ1 and RQ3, we can find that for method-level code smell detection, the most effective PEFT method can differ significantly depending on the models and the amount of available data.
-From Table 5 and Table 6, it is obvious that each model responds to different PEFT techniques in unique ways, suggesting that the optimal PEFT method is rather model-dependent.
-The effectiveness of PEFT methods can change even for the same model under different resource scenarios.
-For instance, in low-resource settings with limited data, the best PEFT method for models like GraphCodeBERT and DeepSeek-Coder-1.3B shifts from prefix tuning to LoRA.
-Moreover, peak GPU consumption can vary across different models (see Section 4.1.2).
-Generally, $(IA)^{3}$ requires more GPU resources than LoRA for code smell detection, whereas the reverse is true for StarCoderBase-7B.
-These variations underscore that there is no universally best PEFT method for method-level code smell detection; instead, the most suitable method depends on the specific model and resource constraints.
-Therefore, we recommend that practitioners consider both the model’s characteristics and the availability of resources, including data volume and GPU memory, when selecting the appropriate PEFT method for a given task.
-Start model selection for code smell detection with a smaller dataset to optimize resource usage: In low-resource scenarios, when the dataset size was reduced by more than 90%, the models’ effectiveness did not degrade significantly, and they still yielded promising results, as shown in the results of RQ3 (see Section 4.3).
+PEFT methods can enhance the capacity of LMs in class-level code understanding: In this study, we selected four representative types of code smells, i.e., CC, CM, FE, and DC, to evaluate the effectiveness of PEFT methods.
+While only DC is formally categorized as a class-level smell, FE detection also inherently requires class-level analysis as well, as it involves analyzing method-class dependencies (see Section 3.3.1 and Section 3.3.4).
+According to the recent studies (silva2024detecting, ; mesbah2025leveraging, ), LLMs not specifically fine-tuned for code smell detection generally perform worse when tasked with identifying smells that demand class-level understanding.
+This aligns with our intuitive expectation that class-level code understanding is more difficult, as it requires capturing higher-level structural relationships (e.g., field-method interactions, class cohesion, and design-level abstractions).
+Despite this challenge, most existing research on PEFT methods has predominantly concentrated on the their perfomance for method-level code understanding (liu2023empirical, ; wang2023one, ; liu2024delving, ), leaving their effectiveness on class-level tasks largely unexplored.
+Our work bridges this critical gap by evaluating PEFT methods on code smell detection, a task spans both method- and class-level analysis.
+Surprisingly, as discussed in Section 4.1.5, our results reveal that LLMs with PEFT methods not only overcome the anticipated class-level challenges but even outperform method-level detection in some cases.
+These findings suggest that PEFT methods not only improve parameter efficiency but also strengthen the proficiency of LLMs to perform robust code understanding across different levels of granularity.
+While larger LMs with PEFT methods often achieve better effectiveness in code smell detection, SLMs still retain distinct advantages: The results of RQ1 reveal that LLMs generally outperform SLMs in code smell detection, particularly in identifying FE.
+This performance advantage aligns with established neural scaling laws (kaplan2020scaling, ), which posit that model capabilities improve predictably with increased scale.
+However, SLMs are able to achieve comparable performance in detecting CC and CM, and even significantly outperform LLMs in the detection of DC.
+Moreover, as illustrated in Figure 5, SLMs exhibit substantially lower peak GPU memory consumption compared to LLMs, making them a more resource-efficient choice.
+These findings highlight a critical trade-off: while LLMs tend to yield higher absolute performance, SLMs offer a competitive and cost-effective alternative, especially in settings with limited computational resources.
+Considering the model used, the code smell type, and the availability of resources when selecting the appropriate PEFT method: Based on the results of RQ1 and RQ3, we can find that for code smell detection, the most effective PEFT method can differ significantly depending on the models and the amount of available data.
+From Table 6 and Table 7, we observe that although $(IA)^{3}$ generally achieves the best performance for LLMs across most code smell types, this is not universally the case.
+For certain models, particularly SLMs, and specific types of code smells, prefix tuning or even prompt tuning may outperform $(IA)^{3}$.
+As indicated in Figure 5, the relatively high peak GPU memory consumption of $(IA)^{3}$ may limit its practicality in constrained environments, where more lightweight PEFT methods could be more suitable.
+Moreover, in low-resource scenarios where training data is limited, LoRA emerges as the most effective choice (see results of RQ3 in Section 4.3.2).
+These findings emphasize the importance of selecting PEFT strategies holistically, taking into account not only the model size and available resources, including both GPU memory and training data, but also the specific demands of the code smell being detected.
+Start model selection for code smell detection with a smaller dataset to optimize resource usage: In low-resource scenarios, even when the dataset size was drastically reduced, the models’ effectiveness did not degrade significantly, and they still yielded promising results, as shown in the results of RQ3.
 This finding indicates that, even with a significantly smaller dataset, both small and large LMs can maintain robust performance for the code smell detection task.
@@
 By leveraging smaller datasets for initial evaluation, one can make informed decisions about which models are likely to perform well with limited resources, and further fine-tune or scale up as necessary, without wasting valuable time or computational power.
+ICL shows promise as an effective alternative to PEFT methods for certain types of code smells: In RQ4, we investigate the performance of general-purpose LLMs in code smell detection under In-Context Learning (ICL) settings.
+As shown in Table 10, for CC and DC, the optimal ICL configuration of DeepSeek-v3 achieves MCC scores only 4%–5% lower than those attained by the top-performing PEFT-tuned LMs.
+This suggests that, for certain types of code smells, well-crafted prompts and a few in-context examples may enable general-purpose LLMs to produce competitive results.
+Moreover, applying ICL technique requires no model training or parameter updates (zhou2024large, ; tom2020language, ; li2025large, ), which is particularly appealing for resource-constrained scenarios or when model weights cannot be modified (e.g., in commercial LLM APIs).
+However, for CM and FE, ICL lags significantly behind, with performance deficits of up to 13.69% and 29.25% in MCC, respectively, compared to our PEFT-based approaches.
+These results indicate that detecting these smell types likely requires deeper model adaptation to capture subtle semantic patterns and structural nuances that few-shot prompting alone cannot convey.
+Overall, while ICL offers a lightweight and flexible alternative, especially for rapid prototyping or scenarios with limited computational resources, PEFT techniques remain the more robust and reliable solution for achieving high-accuracy code smell detection across a broad range of smell types.
+Future work could explore the use of Retrieval-Augmented Generation (RAG) for code smell detection: According to Figure 7, while models generally perform better in code smell detection when provided with randomly selected examples compared to the zero-shot setting, their performance does not consistently improve as the number of example pairs increases.
+This observation suggests that simply increasing the number of in-context examples without considering their quality, relevance, or diversity may not be sufficient to enhance model effectiveness for the task of code smell detection.
+By contrast, Retrieval-Augmented Generation (RAG) (lewis2020retrieval, ) can dynamically retrieve the most relevant examples from an external knowledge base to augment the prompts.
+By grounding predictions in high-quality and contextually appropriate evidence, RAG has the potential to enhance detection accuracy while avoiding the need for fine-tuning or complex prompt engineering.
+As such, future research could explore whether RAG-based approaches can provide a more effective and scalable solution to PEFT methods for code smell detection.
 6. Threats to Validity
@@
 We first used DesigniteJava (designitejava, ) to detect potential positive and negative samples, then manually reviewed all the candidate data to validate the datasets, ensuring there were no false positives or negatives.
-First, relying solely on DesigniteJava for initial sample detection may introduce bias.
-However, DesigniteJava is a well-established and widely adopted tool in Java code smell detection (chaniotaki2021architecture, ; sharma2021qscored, ; shah2023mining, ), providing a robust foundation for the identification of potential code smells.
+First, relying solely on DesigniteJava and PMD for initial sample detection may introduce bias.
+However, DesigniteJava and PMD are well-established and widely adopted tools in Java code smell detection (chaniotaki2021architecture, ; sharma2021qscored, ; shah2023mining, ; tushar2016does, ; taibi2017how, ), providing a robust foundation for the identification of potential code smells.
 We also acknowledge that future studies could benefit from incorporating additional tools in the initial code smell detection process, which may help mitigate the threat to construct validity by providing more diverse perspectives on the data.
 Second, the manual validation process, while aimed at ensuring dataset accuracy, could introduce some subjectivity due to its reliance on human judgment.
-To minimize this threat, we conducted separate pilot validations for CC and CM before formally annotating all the data samples.
-This pilot phase helped ensure that the two annotators used the same criteria for validation, thus improving consistency and reducing potential bias in the dataset labeling.
-External Validity: In this study, we evaluated the effectiveness of state-of-the-art PEFT methods using 4 small LMs and 6 LLMs.
+To minimize this threat, we conducted separate pilot validations for the four types of code smells before formally annotating all the data samples.
+During this phase, we established and refined clear identification criteria to ensure alignment between annotators.
+This approach not only improved labeling consistency but also reduced potential bias in the dataset.
+Lastly, due to budget constraints for commercial API usage, our evaluation of general-purpose LLMs (GPT-4o-mini and DeepSeek-v3) under ICL settings in RQ4 was limited to a maximum of three example pairs.
+This sample size may not fully represent the upper bounds of ICL performance.
+Future work could expand on this by increasing the number of example pairs and evaluating additional models to better understand the full potential of ICL performance.
+However, prior studies such as UniLog (xu2024unilog, ) have shown that the performance of LLMs on software engineering tasks does not always improve linearly with more examples, which suggests that while increasing the number of examples may be beneficial, performance gains are not guaranteed.
+As such, we believe our current evaluation still provides valuable insights and that this threat to validity is partially mitigated.
+External Validity: In this study, we evaluated the effectiveness of state-of-the-art PEFT methods using 4 small LMs and 5 LLMs.
 However, due to limited computational resources, we did not include larger models with over 7B parameters, which may impact the generalizability of our findings.
-Additionally, we only employed four popular and typical PEFT techniques, i.e., prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$.
-Other PEFT methods, such as adapter tuning (houlsby2019parameter, ), could be explored in our future work.
-Moreover, our study only applied PEFT methods to detect Java code smells.
-Expanding this research to include code smells in other programming languages, such as C# and Python, could provide further insights into the effectiveness and adaptability of PEFT methods across diverse language-specific code smell patterns.
+Furthermore, we only employed four widely used and typical PEFT techniques, i.e., prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$.
+While these methods cover the major categories defined in the PEFT library, the exclusion of other approaches may limit the generalizability of our findings.
+For example, methods such as adapter tuning (houlsby2019parameter, ) could be explored in future work to provide a more comprehensive evaluation.
+Moreover, our study applied PEFT methods exclusively to the detection of Java code smells, focusing on four specific types: three method-level smells (Complex Conditional, Complex Method, and Feature Envy) and one class-level smell (Data Class).
+While this selection provides a certain degree of diversity, it does not fully capture the wide spectrum of code smell types encountered in real-world software systems.
+Expanding future work to include a wider range of smell categories would enhance the comprehensiveness of the evaluation.
+In addition, our current work is limited to Java.
+Applying PEFT techniques to detect code smells in other programming languages, such as C# or Python, could yield further insights into their generalizability and adaptability across different language ecosystems.
 Internal Validity: The settings of hyper-parameters for the four PEFT methods and full fine-tuning play a crucial role in the performance of the models.
@@
 To mitigate this threat, we referred to previous works and conducted pilot experiments to determine the appropriate hyper-parameter values.
+Another potential threat to internal validity stems from conducting only a single experimental run per configuration due to computational resource constraints.
+Although we ensured consistency across all experiments by using a fixed random seed (42), this approach cannot fully account for the inherent stochasticity in the training process.
+To assess the impact of randomness, we conducted supplementary experiments using two alternative seeds (0 and 999) for a representative case (GraphCodeBERT on CC detection).
+The results shown in Table 11 indicate that there are only small variations in the performance metrics across different random seeds.
+These minor differences suggest that our core findings are robust to random initialization effects, with the overall trends remaining consistent.
+While some performance fluctuations are observed, they do not significantly alter the conclusions drawn from the experiments.
+However, we acknowledge that conducting multiple runs with different seeds would yield more statistically reliable conclusions.
+Table 11. Results of GraphCodeBERT with PEFT methods and full fine-tuning on detecting CC across different seeds
+Seed Method Precision Recall F1 MCC
+0 Full FT 73.47% 72.17% 71.79% 45.62%
+Prompt Tuning 64.89% 64.89% 64.89% 29.78%
+Prefix Tuning 71.32% 71.30% 71.30% 42.62%
+LoRA 73.17% 73.04% 73.01% 46.21%
+$(IA)^{3}$ 72.44% 72.17% 72.09% 44.61%
+42 Full FT 74.06% 74.02% 74.01% 48.08%
+Prompt Tuning 65.49% 65.33% 65.24% 30.81%
+Prefix Tuning 70.08% 69.89% 69.82% 39.97%
+LoRA 73.15% 73.15% 73.15% 46.30%
+$(IA)^{3}$ 72.50% 72.50% 72.50% 45.00%
+999 Full FT 73.19% 72.83% 72.72% 46.02%
+Prompt Tuning 64.26% 64.24% 64.22% 28.50%
+Prefix Tuning 71.21% 69.57% 68.96% 40.74%
+LoRA 73.07% 72.93% 72.90% 46.00%
+$(IA)^{3}$ 72.23% 72.17% 72.15% 44.41%
 Additionally, data leakage represents another potential threat to internal validity.
@@
 Fontana et al.
-(arcelli2016comparing, ) evaluated 16 ML algorithms across four types of code smells using 74 software systems, reporting that J48 and Random Forest delivered the highest performance.
+(fontana2016comparing, ) evaluated 16 ML algorithms across four types of code smells using 74 software systems, reporting that J48 and Random Forest delivered the highest performance.
 Liu et al.
@@
 They also anticipated that leveraging CodeBERT would enhance the effectiveness of code smell detection.
-Recently, Wu et al.
+We conduct an extensive analysis of the Machine Learning Code Quality (MLCQ) dataset, focusing on how these LLMs perform when prompted to identify and classify code smells.
+Recently, researchers have started to explore the use of LLMs for code smell detection.
+Mesbah et al.
+(mesbah2025leveraging, ) used the MLCQ dataset to analyze how GPT-4 and LLaMA perform when prompted to identify and classify code smells.
+Silva et al.
+(silva2024detecting, ) employed two types of prompts to query ChatGPT for identifying Blob, Data Class, Feature Envy, and Long Method code smells.
+Sadik et al.
+(sadik2025benchmarking, ) created a dataset for evaluating LLMs on automated code smell detection.
+They benchmarked two state-of-the-art LLMs and found that LLMs are not yet ready to replace deterministic tools in all aspects of code quality assurance.
+Wu et al.
 (wu2024ismell, ) trained a Mixture of Experts (MoE) model, leveraging specialized expert toolsets for code smell detection.
 They subsequently utilized the detection results to prompt LLMs for refactoring.
-Most prevailing ML- or DL-based approaches for code smell detection train models from scratch.
-Compared to their works, we harness the power of LLMs using PEFT methods to detect two types of complex method-level code smells (Complex Conditional and Complex Method), aiming to enhance the models’ ability to capture the nuanced relationships between different types of method-level code smells.
+Most prevailing ML-based or DL-based approaches for code smell detection train models from scratch.
+On the other hand, existing LLM-based approaches typically rely on simple prompting strategies without any form of fine-tuning or adaptation.
+Compared to their works, we harness the power of LLMs using PEFT methods to detect four types of widely studied code smells (Complex Conditional, Complex Method, Feature Envy, and Data Class), aiming to enhance the models’ ability to capture the nuanced relationships between different code smell types.
 7.2. PEFT Methods on Software Engineering Tasks
-Several studies have evaluated PEFT methods on various SE tasks.
+Several studies have evaluated PEFT methods on various Software Engineering (SE) tasks.
 Wang et al.
@@
 (weyssow2023exploring, ) compared PEFT methods, including prompt tuning and others, with In-Context Learning (ICL) in terms of code generation.
-Their investigation reveals the superiority of PEFT over ICL and highlights the potential opportunities of tuning LLMs in diverse software engineering scenarios.
+Their investigation reveals the superiority of PEFT over ICL and highlights the potential opportunities of tuning LLMs in diverse SE scenarios.
 Li et al.
@@
 The results show that $(IA)^{3}$ achieves the highest fixing ability compared to prompt tuning, prefix tuning, and LoRA.
-These studies focus on evaluating PEFT methods across a range of software engineering tasks, such as code summarization and code generation.
-Compared to their works, our work pays specific attention to a binary classification task, aiming to assess the effectiveness of LLMs with PEFT methods in identifying different types of code smells.
+These studies focus on evaluating PEFT methods across a range of SE tasks, primarily at the method level.
+Compared to their works, our work specifically targets the task of code smell detection, which, to the best of our knowledge, has not been studied before.
+Additionally, our work further explores the capacity of LLMs on class-level code understanding, thereby addressing a significant gap in existing research.
 8. Conclusions and Future Work
@@
 We also evaluated how the performance of these PEFT methods varies under different conditions, such as varying hyper-parameter settings and the availability of training data.
-Our analysis reveals the efficiency of PEFT methods compared to full fine-tuning, underscoring the potential of PEFT methods to enhance method-level code smell detection and offer valuable insights for optimizing LMs for other software engineering tasks.
-Additionally, we have constructed high-quality datasets for CC and CM code smells, and we provided the replication package of this work online (replicationPackage, ) to facilitate future research and ensure the reproducibility of our study results.
+Our analysis reveals that PEFT methods are more efficient than full fine-tuning and In-Context Learning (ICL) in most cases, underscoring the potential of PEFT methods to enhance code smell detection and offer valuable insights for optimizing LMs for other Software Engineering (SE) tasks.
+Additionally, we have constructed high-quality datasets for four widely studied code smells, i.e., Complex Conditional, Complex Method, Feature Envy, and Data Class, and we provided the replication package of this work online (replicationPackage, ) to facilitate future research and ensure the reproducibility of our study results.
 In the next step, we plan to further explore how PEFT performs in cross-language scenarios for code smell detection.
-Moreover, we plan to investigate the application of PEFT methods for fine-tuning LMs to assist in code smell refactoring, thereby improving overall code quality and maintainability by providing just-in-time code smell detection and refactoring suggestions.
+Moreover, we plan to investigate the application of PEFT methods for fine-tuning LMs to assist in a wider range of SE tasks, such as bug detection, code completion, and code smell refactoring.
+These directions will help us better understand the generalizability and practical value of PEFT methods in real-world software development settings.
 Data Availability
@@
 Acknowledgements.
-This work has been partially supported by the National Natural Science Foundation of China (NSFC) with Grant No. 62172311.
+This work has been partially supported by the National Natural Science Foundation of China (NSFC) with Grant No.
+62172311.
+The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of Wuhan University.

### v3

@@
 Abstract.
-Code smells, which are suboptimal coding practices that can potentially lead to defects or maintenance issues, can negatively impact the quality of software systems.
-Most existing code smell detection methods rely on heuristics-based or machine learning (ML) and deep learning (DL)-based techniques.
-However, these techniques have several drawbacks (e.g., unsatisfactory performance).
-Large Language Models (LLMs) have garnered significant attention in the Software Engineering (SE) field, achieving state-of-the-art performance across a wide range of SE tasks.
-Parameter-Efficient Fine-Tuning (PEFT) methods, which are commonly used to adapt LLMs to specific tasks with fewer parameters and reduced computational resources, have emerged as a promising approach for enhancing the performance of LLMs in various SE tasks.
-However, LLMs have not yet been explored for code smell detection, and their effectiveness for this task remains unclear.
-Furthermore, no comprehensive investigation has been conducted on the efficiency of PEFT methods for code smell detection.
-In this regard, we systematically evaluate the effectiveness of state-of-the-art PEFT methods on both small and large Language Models (LMs) for code smell detection.
-To begin, we constructed high-quality Java code smell datasets sourced from GitHub.
-We then fine-tuned four small LMs (SLMs) and five LLMs using various PEFT techniques, including prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$, for code smell detection.
-Our comparison against full fine-tuning revealed that PEFT methods not only achieve comparable or better effectiveness but also consume less peak GPU memory.
-Our analysis further explored the performance of SLMs versus LLMs in the context of code smell detection.
-Generally, LLMs outperform SLMs on detecting Complex Conditional, Complex Method, and Feature Envy.
-However, SLMs demonstrate superior performance on detecting Data Class.
-We also investigated the impact of varying hyper-parameter settings and the availability of training data.
-Our results show that increasing the number of trainable parameters does not necessarily improve the effectiveness of PEFT methods for code smell detection.
-In contrast, we observed significant performance gains when the training dataset size was increased.
-Finally, our results show that LMs with PEFT methods outperform traditional heuristic-based detectors, advanced DL-based approaches, and general-purpose LLMs under In-Context Learning in the task of code smell detection.
-Specifically, we observe improvements in MCC ranging from 4.11% to 19.25% across the four types of code smells studied, demonstrating a more effective solution for this task.
-In conclusion, our work provides a comprehensive evaluation of PEFT methods on both small and large LMs for code smell detection.
-The results highlight the efficiency and effectiveness of PEFT methods in this domain, offering valuable insights for optimizing LMs for code smell detection.
-We suggest that the choice of PEFT method for code smell detection should be guided by the model used, data availability, and computational resources.
-Our replication package is available at (replicationPackage, ), enabling others to reproduce and extend our findings.
+Code smells are indicators of suboptimal design that negatively impact software quality.
+However, automated code smell detection remains a persistent challenge: heuristics-based tools suffer from high sensitivity to threshold selection and inherent subjectivity, while Machine Learning (ML) and Deep Learning (DL) models yield unsatisfactory performance.
+Although Large Language Models (LLMs) offer a promising solution to these limitations, their adaptation is impeded by the prohibitive cost of full fine-tuning and the lack of “LM-ready” benchmarks, as existing code smell datasets largely rely on software metrics or contain noisy, unverified labels.
+To address these gaps and enable a rigorous assessment of efficient LLM adaptation for code smell detection, we present a comprehensive study comprising two synergistic contributions.
+First, we constructed a high-quality, source-code-centric benchmark covering four non-trivial code smells—Complex Conditional, Complex Method, Feature Envy, and Data Class—and ensured its reliability through a meticulous two-stage manual review.
+Second, leveraging this benchmark, we systematically evaluated the effectiveness of four Parameter-Efficient Fine-Tuning (PEFT) methods (i.e., prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$) across four Small LMs (SLMs) and five LLMs.
+We benchmarked these PEFT-tuned LMs against a broad suite of baselines, including traditional heuristics-based detectors (DesigniteJava and PMD), DL-based approaches (DeepSmells and AE-Dense), and cutting-edge general-purpose LLMs (GPT-4o-mini and DeepSeek-v3) with In-Context Learning.
+Our results demonstrate that PEFT methods match or exceed full fine-tuning performance while significantly reducing peak GPU memory usage, and outperform all baselines with MCC improvements ranging from 0.33% to 13.69%, particularly for some code smells.
+Furthermore, we provide actionable insights into the choice of PEFT method for code smell detection, which should be guided by the model used, data availability, and computational resources.
+Our replication package is available at (replicationPackage) to support reproducibility and future extensions.
 Code Smell Detection, Large Language Model, Parameter-Efficient Fine-Tuning
@@
 1. Introduction
-Code smells are indicators of potential issues in code or design (fowler1999refactoring, ), which could adversely affect software quality, particularly in terms of maintainability (palomba2018maintainability, ), code readability (abbes2011empirical, ), and testability (chatzigeorgiou2010investigating, ).
-Detecting code smells can be used to identify software design issues that lead to potential problems in future development and maintenance (fowler1999refactoring, ; lanza2007object, ; moha2010decor, ).
-Manual identification of code smells can be time-consuming and costly (prokic2024prescriptive, ), prompting many researchers to explore automated methods for more efficient code smell detection.
-Existing automated code smell detection methods are typically heuristics-based (marinescu2004detection, ; morales2017on, ; moha2010decor, ).
-Traditional heuristics-based detectors for code smells rely on a set of manually curated features and thresholds to discriminate between smelly and non-smelly classes (azeem2019machine, ).
-While it has been shown that these static detectors provide reasonably accurate results, they are proven to have limitations that hinder their practical application (fernandes2016review, ; zhang2011code, ), particularly because their performance is highly dependent on the selection of thresholds.
-Besides, code smells identified by such heuristics-based methods can be subjectively interpreted by developers, leading to a significant gap between detection results and developers’ agreement (azeem2019machine, ; fontana2016antipattern, ; mantyla2006subjective, ; schumacher2010building, ).
-To address these issues, many researchers have incorporated machine learning (ML) and deep learning (DL) algorithms to mitigate the performance instability in code smell detection that arises from the manual threshold-setting process in traditional detectors (fontana2013code, ; fontana2016comparing, ; sharma2021code, ; liu2021deep, ).
-By learning to discern between smelly and non-smelly code elements, ML and DL techniques offer a solution to the subjectivity problem inherent in heuristics-based methods (nucci2018detecting, ).
-However, despite these advancements, ML and DL methods still face significant challenges in achieving considerable performance for code smell detection (pereira2022code, ; alazba2023deep, ).
-In recent years, in the field of Natural Language Processing (NLP), increased computational power, advanced machine learning techniques, and access to very large-scale data have led to the emergence of Large Language Models (LLMs).
-The exceptional performance of LLMs in NLP tasks (e.g., text classification, generation, and summarization) has garnered significant attention and interest from the Software Engineering (SE) domain.
-Researchers have begun employed LLMs, particularly code-specific LLMs like CodeBERT-Base (feng2020codebert, ) and CodeT5 (wang2021codet5, ), to tackle various SE tasks, such as code review (lu2023llamareviewer, ) and vulnerability repair (le2024study, ).
-However, directly using LLMs without any task-specific fine-tuning for code smell detection yields suboptimal results (mesbah2025leveraging, ; silva2024detecting, ; sadik2025benchmarking, ), as these models lack adaptation to the unique demands of the task.
-Fine-tuning enables the adaptation of Language Models (LMs) to specific downstream tasks by updating model parameters, allowing them to better capture task-relevant features and improve performance.
-Among all the fine-tuning methods, Parameter-Efficient Fine-Tuning (PEFT) has emerged as a promising approach, offering an efficient alternative to full fine-tuning, particularly when computational resources are limited.
-PEFT methods have shown potential in adapting LLMs to various downstream tasks with fewer parameters and reduced resource requirements, thus expanding the practical applicability of LLMs in diverse SE contexts.
-Previous studies have explored the efficiency of PEFT methods for LMs on automated program repair (li2024comprehensive, ), code generation (liu2023empirical, ), code search (wang2023one, ), etc.
-However, to the best of our knowledge, no prior research has focused on improving the performance of LLMs for code smell detection, nor has any in-depth study been conducted analyzing the application of PEFT methods in this area.
-Additionally, identifying the code smells selected in this study requires both method- and class-level code understanding.
-While existing research on PEFT methods has predominantly focused on code-related tasks at the method level (li2024comprehensive, ; liu2023empirical, ; wang2023one, ; liu2024delving, ; weyssow2023exploring, ), a significant gap remains in assessing their effectiveness at the class level.
-By leveraging the task of code smell detection, our study aims to fill this gap by systematically exploring the capacities of LMs with PEFT methods in class-level code understanding.
-Furthermore, utilizing PEFT methods for code smell detection can facilitate just-in-time code smell detection.
-Traditional heuristic-based methods for code smell detection are often applied after the software systems have been completed (designitejava, ; pmd, ; marinescu2005iplasma, ).
-As a result, code smells are typically discovered at later stages of development, requiring refactoring that involves modifying all relevant parts of the system, thereby increasing the cost and complexity of addressing these code smells.
-In contrast, incorporating LLMs with PEFT techniques allows for real-time detection of code smells, enabling developers to identify and address potential code smells as they write code.
-For instance, upon completing a function, developers can immediately receive feedback on whether it contains a code smell.
-This proactive approach contributes to maintaining high code quality throughout the development life cycle, leading to more efficient and effective code maintenance and improvement (ardimento2021temporal, ; ardimento2021transfer, ).
-To this end, we conduct a series of experiments on fine-tuning both small and large LMs with PEFT methods on the code smell detection task.
-In this paper, we define small LMs (SLMs) as those with fewer than 1B parameters, while LMs with 1B parameters or more are considered LLMs.
-Specifically, we construct high-quality Java code smell datasets comprising four types of widely studied and sophisticated code smells at both the method and class level, i.e., Complex Conditional (CC), Complex Method (CM), Feature Envy (FE), and Data Class (DC).
-These datasets are developed to address the lack of suitable existing resources, as most available datasets either do not conform to the required data formats or are based on outdated codebases.
-We then apply various PEFT methods, including prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$, to fine-tune four popular SLMs and five cutting-edge LLMs with parameter sizes ranging from 125M to 7B.
-The effectiveness of these PEFT methods is assessed for code smell detection, and their performance is compared to that of full fine-tuning in terms of both effectiveness and peak GPU memory usage.
-Additionally, we investigate the performance of PEFT techniques under different hyper-parameter settings and evaluate their effectiveness in low-resource scenarios with varying numbers of training samples on the code smell detection task.
-Finally, we employ a heuristics-based static analysis tool alongside two state-of-the-art DL-based code smell detection tools to compare the performance of PEFT methods with existing approaches.
-The main contributions of this work are as follows:
-  • We mine Java repositories from GitHub and employ traditional heuristic-based detectors, along with a two-stage manual review, to create high-quality and balanced datasets for four types of code smells, i.e., Complex Conditional, Complex Method, Feature Envy, and Data Class.
-  • We conduct a comprehensive evaluation of multiple PEFT methods on both small and large LMs for code smell detection.
-Our results indicate that LLMs, when fine-tuned, do not always outperform SLMs.
-Moreover, PEFT methods achieve better or comparable effectiveness compared to full fine-tuning.
-Overall, LMs with PEFT methods exhibit promising capabilities in method- and class-level code understanding.
-  • We explore the performance of PEFT methods under different hyper-parameter settings and in low-resource scenarios.
-Our results reveal that adjusting hyper-parameters to increase trainable parameters does not necessarily lead to better effectiveness in detecting code smells.
-However, the performance of PEFT methods improves with more training samples.
-  • We compare a heuristic-based detector, two best-in-class DL-based detection tools, and two general-purpose LLMs using In-Context Learning (ICL) against our PEFT-based method on code smell detection.
-Our results demonstrate that LMs with PEFT methods consistently outperform all baselines across multiple code smell types, highlighting the superiority of PEFT methods.
-  • We provide actionable suggestions and insights for both developers and researchers for using PEFT methods on code smell detection.
-The remainder of this paper is organized as follows: Section 2 presents the background information.
-Section 3 outlines the research questions and the process of dataset construction, while the study results are provided in Section 4.
-Section 5 discusses the insights for practitioners.
-Potential threats to validity are addressed in Section 6, with related work reviewed in Section 7.
-Finally, Section 8 concludes the paper and suggests future directions for research.
-2. Background
-2.1. The Need of Applying Language Models on Code Smell Detection
-Existing heuristic-based code smell detectors rely on pre-defined rules and metrics, while ML- or DL-based approaches depend on algorithms to identify the optimal threshold for code smell classification.
-However, these tools often fail to consider the inherent context of the code, which is crucial for accurately detecting and understanding code smells (sharma2021code, ).
-For example, the Java method shown in Figure 1 contains a switch-case statement with seven cases.
-From the perspective of a Java developer, this method simply maps the numbers 1 through 7 to the days of the week, from Monday to Sunday.
-It is straight-forward and easy to understand, as all the case statements follow the same structure.
-However, when scanned by an existing heuristic-based code smell detector, such as DesigniteJava (designitejava, ), the tool flags the method with the Complex Method code smell.
-This decision is based solely on the method’s cyclomatic complexity, which exceeds a pre-defined threshold, without considering the method’s context.
-In contrast, Language Models (LMs), such as ChatGPT (achiam2023gpt, ), have the advantage of considering the inherent context of the code (wang2021codet5, ; roziere2024codellama, ).
-The contextual understanding allows the LM to accurately determine that the method does not exhibit Complex Method, as it can perform a role similar to a skilled developer assessing the code.
-The LM can evaluate not only traditional metrics, such as cyclomatic complexity, but also analyze additional factors, including readability, modularity, and overall design.
-Therefore, compared to traditional heuristic-based static analysis tools and certain ML- or DL-based detectors, LMs have the potential to detect code smells more accurately by considering both structural and contextual factors, rather than relying solely on numerical thresholds or pre-defined rules.
-Figure 1. An example of a Java method that is considered to contain Complex Method code smell by the heuristic-based tool
-2.2. Full Fine-tuning
-Language Models (LMs) which are pre-trained on massive general-purpose corpus via self-supervised method, can be applied to various downstream NLP tasks by adapting their rich possession of learned knowledge through full fine-tuning (qiu2020pre, ).
+Code smells are indicators of underlying design or implementation issues that may degrade software quality (fowler1999refactoring), particularly in terms of maintainability (palomba2018maintainability), readability (abbes2011empirical), and testability (chatzigeorgiou2010investigating).
+Detecting code smells helps to uncover software design flaws that lead to potential problems in future development and maintenance (fowler1999refactoring; lanza2007object; moha2010decor).
+However, manual identification of code smells is often time-consuming and cost-prohibitive (prokic2024prescriptive), which has prompted extensive research into automated detection techniques.
+Early work predominantly relied on heuristics-based detectors (marinescu2004detection; morales2017on; moha2010decor).
+These methods apply manually curated features and thresholds to discriminate between smelly and non-smelly instances (azeem2019machine).
+While effective in some settings, such static tools exhibit well-documented shortcomings (fernandes2016review; zhang2011code): their performance is highly sensitive to threshold selection, they offer limited semantic understanding, and their results often diverge from human judgment due to the inherent subjectivity of code smells (azeem2019machine; fontana2016antipattern; mantyla2006subjective; schumacher2010building).
+To overcome these limitations, researchers have increasingly incorporated Machine Learning (ML) and Deep Learning (DL) algorithms for code smell detection.
+These approaches aim to mitigate the performance instability inherent in the manual threshold-setting of heuristics-based detectors (fontana2013code; fontana2016comparing; sharma2021code; liu2021deep).
+By reducing reliance on rigid thresholds, ML and DL models can partially learn semantic cues (nucci2018detecting).
+Nonetheless, recent studies indicate that these models still struggle to achieve considerable performance (pereira2022code; alazba2023deep), as they frequently fail to capture the deeper semantic and structural relationships within source code.
+Recently, Large Language Models (LLMs) have emerged as a transformative force in Software Engineering (SE), achieving state-of-the-art performance across a wide range of tasks (lu2023llamareviewer; le2024study).
+Despite their impressive general capabilities, directly employing off-the-shelf LLMs for code smell detection without any task-specific fine-tuning yields suboptimal results (mesbah2025leveraging; silva2024detecting; sadik2025benchmarking).
+While effective adaptation is essential, two primary challenges currently impede progress:
+Challenge 1: Lack of High-Quality, LM-Ready Datasets.
+Most existing code smell datasets comprise calculated software metrics rather than the raw source code required by LMs.
+Moreover, they typically focus on relatively simple smells that do not demand complex semantic reasoning.
+Furthermore, datasets that do include code often rely on noisy labels generated directly by heuristics-based detectors without human verification, leading to the “Oracle Problem”.
+This incompatibility renders existing benchmarks unsuitable for the effective training or evaluation of LMs in code smell detection.
+Consequently, without high-quality, source-code-centric benchmarks covering non-trivial code smells, a fair investigation into the effectiveness of LMs for code smell detection is infeasible.
+Challenge 2: Unclear Effectiveness of LM Adaptation Methods.
+Although Parameter-Efficient Fine-Tuning (PEFT) methods have shown promise across various SE tasks by enabling task-specific adaptation with only a small number of trainable parameters (wang2023one; liu2023empirical; he2022towards; liu2024delving; weyssow2023exploring; li2024comprehensive), their effectiveness for code smell detection remains largely unexplored.
+It is still unclear whether PEFT approaches can match the performance of full fine-tuning or outperform existing baselines in code smell detection.
+Investigating PEFT methods in the context of code smell detection also fills a broader research gap.
+Prior studies on PEFT have primarily concentrated on method-level tasks (li2024comprehensive; liu2023empirical; wang2023one; liu2024delving; weyssow2023exploring), leaving their effectiveness in class-level contexts largely unexplored.
+Our study addresses this gap by exploring the capabilities of PEFT-adapted LMs in capturing class-level semantics through the detection of code smells, such as Feature Envy and Data Class, which inherently necessitate class-level code understanding.
+Furthermore, utilizing PEFT techniques facilitates just-in-time code smell detection.
+Traditional heuristics-based detectors are often applied post-development (designitejava; pmd; marinescu2005iplasma), discovering issues only after software systems are fully implemented and thereby increasing refactoring overhead.
+In contrast, PEFT-tuned LMs enable real-time identification of code smells, allowing developers to receive immediate feedback upon completing a code block.
+This proactive approach supports high code quality throughout the development lifecycle, leading to more efficient software maintenance (ardimento2021temporal; ardimento2021transfer).
+To this end, we systematically evaluate the effectiveness of state-of-the-art PEFT methods on both Small LMs (SLMs, <<1B parameters) and Large LMs (LLMs, $\geq$1B parameters) for code smell detection.
+To facilitate this evaluation and resolve the aforementioned data bottleneck, we first curated a high-quality Java code smell dataset covering four types of widely studied and sophisticated code smells at both method and class levels: Complex Conditional (CC), Complex Method (CM), Feature Envy (FE), and Data Class (DC).
+Crucially, we subjected this dataset to a rigorous two-stage manual review to guarantee its reliability.
+With a robust benchmark established, we then addressed the adaptation challenge by fine-tuning four SLMs and five LLMs using various PEFT techniques (i.e., prompt tuning, prefix tuning, LoRA, and (IA)3).
+We compared these PEFT-tuned LMs against a comprehensive suite of baselines, encompassing heuristics-based detectors, DL-based approaches, and general purpose LLMs with In-Context Learning (ICL).
+The main contributions of this work are summarized as follows:
+  • LM-Ready Benchmark for Code Smell Detection: We construct a high-quality, source-code-centric dataset covering four non-trivial code smells—Complex Conditional, Complex Method, Feature Envy, and Data Class.
+By mining recent Java repositories, applying heuristics-based detectors, and performing a stringent two-stage manual review, we address the limitations of existing code smell datasets (e.g, reliance on code metrics, focusing on simple smells, and unverified labels), ensuring high data quality and a balanced distribution.
+This establishes a reliable ground truth necessary for training and evaluating LMs for code smell detection, serving as the foundation for our subsequent experiments.
+  • Systematic Evaluation of PEFT Methods: Leveraging our curated dataset, we conduct the first extensive evaluation of PEFT methods for code smell detection on both SLMs and LLMs.
+We compare multiple PEFT techniques, including prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$, against full fine-tuning.
+Our analysis comprehensively assesses their effectiveness, peak GPU memory usage, and performance under different hyper-parameter configurations and low-resource scenarios.
+  • Comprehensive Comparison with SOTA Baselines: To validate the utility of PEFT-tuned LMs for code smell detection, we benchmark these techniques against a wide range of baselines.
+These baselines span traditional heuristics-based tools, advanced DL-based detectors, and general-purpose LLMs with In-Context Learning under zero-shot, few-shot, and criteria-informed settings.
+Our results demonstrate that PEFT methods consistently achieve comparable or superior effectiveness while reducing computational cost, highlighting their practical value for the code smell detection.
+  • Practical Insights for Developers and Researchers: Based on our extensive experiments, we distill actionable recommendations for applying PEFT methods to code smell detection.
+We offer specific guidance on model selection, hyper-parameter tuning strategies, and the trade-offs between efficiency and performance, providing a useful reference for future research and practice in this domain.
+The remainder of this paper is organized as follows: Section 2 introduces the background and motivation.
+Section 3 details the dataset construction process, while Section 4 outlines the research questions and experimental setup.
+The study results are presented in Section 5, followed by a discussion in Section 6.
+Threats to validity are addressed in Section 7, with the related work reviewed in Section 8.
+Finally, Section 9 concludes the paper.
+2. Background and Motivation
+2.1. The Case for Applying Language Models on Code Smell Detection
+Existing heuristics-based code smell detectors rely on predefined rules and metrics, while ML- or DL-based approaches depend on algorithms to identify the optimal threshold for code smell classification.
+However, these tools often overlook the semantic and contextual information of code, which is crucial for accurately detecting and understanding code smells in real-world software development (sharma2021code).
+For instance, the Java method shown in Figure 1 contains a switch-case statement with seven cases.
+From the perspective of a Java developer, this method simply maps the numbers 1 through 7 to the days of the week.
+It is straightforward and easy to understand, as all the case statements follow the same structure.
+While the method could be refactored to reduce repetition, it has a single responsibility, and its logic is immediately clear to a reader.
+Therefore, the method does not exhibit a Complex Method (CM) code smell in practice.
+However, when scanned by the heuristics-based code smell detector, DesigniteJava (designitejava), the method is flagged as CM solely because its cyclomatic complexity exceeds a predefined threshold, while its semantic clarity is ignored.
+Figure 1. An example of Complex Method (CM) code smell detection by DesigniteJava, ChatGPT, and human developers
+public static String format2Duration(long ms) {
+long days = MILLISECONDS.toDays(ms);
+long hours = MILLISECONDS.toDurationHours(ms);
+long minutes = MILLISECONDS.toDurationMinutes(ms);
+long seconds = MILLISECONDS.toDurationSeconds(ms);
+if (days == 0 && hours == 0 && minutes == 0 && seconds == 0) {
+seconds = 1;
+StringBuilder strBuilder = new StringBuilder();
+if (days > 0) strBuilder.append(days).append("d ");
+if (hours > 0) strBuilder.append(hours).append("h ");
+if (minutes > 0) strBuilder.append(minutes).append("m ");
+if (seconds > 0) strBuilder.append(seconds).append("s");
+return strBuilder.toString();
+(a) False positive: identified as CC-positive by DesigniteJava but clean labeled by developers
+public boolean onActionItemClicked(ActionMode mode, MenuItem menuItem) {
+int id = menuItem.getItemId();
+if (id == R.id.delete_entry) {
+confirmDelete(mode);
+return true;
+} else if (id == R.id.select_all) {
+if (mSelected.size() >= mAdapter.getCount())
+mode.finish();
+else {
+for (int i = 0; i < mAdapter.getCount(); i++) {
+if (!mListView.isItemChecked(i))
+mListView.setItemChecked(i, true);
+return true;
+} else {
+return false;
+(b) False negative: identified as clean by DesigniteJava but contains developer-labeled CC code smell
+Figure 2. Examples of Complex Conditional (CC) code smell detection by DesigniteJava, highlighting a false positive (a) and a false negative (b)
+Figure 2 illustrates additional limitations in rule-based detection for Complex Conditional (CC) code smells.
+Example (a) is identified as CC-positive by DesigniteJava but is considered clean by human developers: the conditional if (days == 0 && hours == 0 && minutes == 0 && seconds == 0) exceeds the operand threshold predefined by DesigniteJava, yet the logic is simple and readable, resulting in a false positive.
+Example (b) is labeled as clean by DesigniteJava but contains a developer-identified CC: nested loops and interacting conditionals make the overall logic difficult to follow, even though no single conditional exceeds structural thresholds, resulting in a false negative.
+In contrast, Language Models (LMs) such as ChatGPT (achiam2023gpt) can leverage semantic and contextual understanding (wang2021codet5; roziere2024codellama), enabling detection of both CM and CC code smells consistent with human judgment.
+For CM, LMs can recognize that a method with high cyclomatic complexity or many lines of code may still be clear and maintainable if its logic is transparent and responsibilities are well-defined.
+For CC, LMs can evaluate not only structural metrics, such as nesting depth or the number of boolean operands, but also semantic aspects such as readability, logical flow, and how the conditions interact, allowing them to correctly identify both false positives and false negatives that rule-based tools might misclassify.
+By integrating structural and semantic reasoning, LMs offer a more accurate and nuanced approach than traditional heuristics or purely threshold-based ML or DL tools.
+However, applying LMs on code smell detection effectively requires high-quality data and efficient training strategies, which motivates our work.
+2.2. The Gap in Code Smell Detection Datasets
+Despite the growing use of Large Language Models (LLMs) in software engineering research, existing code smell datasets remain a major bottleneck for effective model training.
+We identify three critical limitations that make current public datasets inadequate for fine-tuning LMs for code smell detection.
+1.
+Predominance of Metric-Based Data: Traditional code smell detection treated the task as tabular classification, leading to widely used datasets containing only software metrics rather than actual source code (fontana2016comparing; binh2022mlcodesmell; madeyski2023detecting).
+While suitable for training classical classifiers like Random Forests or Logistic Regression, these metric-based datasets are incompatible with LMs, which require tokenized source code to capture semantic and syntactic patterns.
+2.
+Limited Scope of Code Smell Types: Existing datasets often lack diversity in the types of smells covered (fontana2016comparing; madeyski2020mlcq; fontana2017code; liu2021deep; moha2009decor).
+There is a disproportionate focus on easily detectable, metric-based smells (e.g., Long Method).
+Conversely, there is a scarcity of high-quality samples for structurally intricate smells, such as Complex Conditional and Complex Method, which necessitate semantic reasoning beyond simple metric thresholds.
+3.
+The “Oracle Problem” and Label Reliability: A significant number of large-scale datasets rely on heuristics-based detectors (e.g., PMD, DECOR, and iPlasma) as the “oracle” to automatically generate labels, typically bypassing manual verification (moha2009decor; lenarduzzi2019technical; sae2016context; yang2015classification; khomh2012exploratory).
+However, the inherent inaccuracies of these tools introduce noise into the ground truth (palomba2018diffuseness; de2017towards).
+Training an LM on such datasets risks merely teaching the model to emulate the flaws and biases of heuristics-based detectors rather than identifying genuine design issues.
+In summary, current code smell datasets are not “LM-ready”.
+To effectively fine-tune an LM for code smell detection, there is an urgent need for a dataset that provides raw, compilable source code, covers complex semantic smells, and relies on high-quality, human-validated labels.
+This motivates the construction of our proposed dataset.
+2.3. Full Fine-Tuning and Parameter-Efficient Fine-Tuning (PEFT)
+2.3.1. Full Fine-Tuning
+Language Models (LMs) which are pre-trained on massive general-purpose corpus via self-supervised method, can be applied to various downstream NLP tasks by adapting their rich possession of learned knowledge through full fine-tuning (qiu2020pre).
 Full fine-tuning involves updating all parameters of an LM, including multi-head attention (MHA), feed-forward neural network (FFN), layer normalization, and so on.
-However, as LMs scale up, full fine-tuning a large-sized LM, i.e., LLM, can be excessively costly and ultimately make it practically unfeasible (hu2024survey, ; ding2023parameter, ).
-2.3. Parameter-Efficient Fine-Tuning (PEFT) Methods
+However, as LMs scale up, full fine-tuning a large-sized LM, i.e., LLM, can be excessively costly and ultimately make it practically unfeasible (hu2024survey; ding2023parameter).
+2.3.2. Parameter-Efficient Fine-Tuning (PEFT) Methods
 Instead of full fine-tuning, pre-trained LMs could be tailored to downstream tasks by Parameter-Efficient Fine-Tuning (PEFT) methods.
 PEFT methods involve optimizing only a subset of model parameters or incorporating external modules for new tasks while keeping the majority of parameters frozen.
-Compared to full fine-tuning, PEFT methods can reduce the substantial computational and data demands of LLMs and achieve comparable or even better results (han2024parameter, ).
-The right side of Figure 2 illustrates several state-of-the-art PEFT methods employed in this study, including prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$.
+Compared to full fine-tuning, PEFT methods can reduce the substantial computational and data demands of LLMs and achieve comparable or even better results (han2024parameter).
+In this study, we employed the Hugging Face PEFT library [^1] https://github.com/huggingface/peft to implement and evaluate all PEFT methods.
+This library, developed by the Hugging Face team, offers a standardized and efficient framework that supports several widely adopted PEFT techniques and has been extensively used in recent research (li2024comprehensive; weyssow2023exploring; wang2025beyond; storhaug2024parameter).
+In addition, it is fully compatible with the Transformers library, allowing seamless integration with open-source LMs available on the Hugging Face platform.
+This unified setup ensures reproducibility, simplifies implementation, and enables fair comparisons across different PEFT strategies.
+Therefore, in this work, we focus exclusively on the PEFT methods provided by the PEFT library.
+The right side of Figure 5 illustrates the four state-of-the-art PEFT techniques adopted in our study: prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$.
 These methods were selected based on several considerations.
-First, they are all supported by the widely used Hugging Face PEFT library [^1] https://github.com/huggingface/peft, which provides a standardized and efficient framework for applying PEFT to LMs.
-Second, according to the PEFT library’s documentation [^2] https://huggingface.co/docs/peft/index, PEFT methods are categorized into three main types: Prompt-based methods, LoRA methods, and $(IA)^{3}$.
+First, they are all natively supported by the PEFT library, allowing stable fine-tuning of LMs while avoiding out-of-memory errors.
+Second, according to the official documentation [^2] https://huggingface.co/docs/peft/index, PEFT methods are categorized into three main types: Prompt-based methods, LoRA methods, and $(IA)^{3}$.
 The four selected methods span these categories, making them representative of the major approaches currently available.
-Finally, these methods are not only recent but also widely adopted in contemporary research (weyssow2023exploring, ; li2024comprehensive, ; mannisto2025comparative, ; zhou2024empirical, ; taylor2022galactica, ; esmaeili2024empirical, ; wistuba2024choice, ).
+Finally, these methods are not only recent but also widely adopted in contemporary research (weyssow2023exploring; li2024comprehensive; mannisto2025comparative; zhou2024empirical; taylor2022galactica; esmaeili2024empirical; wistuba2024choice).
 Their demonstrated effectiveness across various tasks and their popularity in both the NLP and SE communities make them ideal candidates for this study.
 By selecting these four methods, we ensure that our experiments cover a range of the most promising and commonly adopted techniques in PEFT.
-Prompt tuning (lester2021power, ) directly appends learnable vectors, known as soft prompts, to the initial input embedding layer.
+Prompt tuning (lester2021power) directly appends learnable vectors, known as soft prompts, to the initial input embedding layer.
 When fine-tuning PLMs with prompt tuning, only the soft prompts are updated, while the remaining model weights and architecture are kept fixed (not updated).
-Prefix tuning (li2021prefix, ) introduces adjustable prefix (i.e., a sequence of task-specific vectors) across all Transformer layers.
+Prefix tuning (li2021prefix) introduces adjustable prefix (i.e., a sequence of task-specific vectors) across all Transformer layers.
 Specifically, these trainable prefix vectors are prepended to the key $K$ and value $V$ of all the MHA mechanisms.
-Prefix tuning adopts a reparameterization strategy, utilizing a multilayer perceptron (MLP) layer to generate the added vectors instead of optimizing them in a straightforward manner (han2024parameter, ).
-LoRA (Low-Rank Adaptation) (hu2022lora, ) injects trainable low-rank decomposition matrices into the query $W_{q}$, key $W_{k}$, and value $W_{v}$ projection matrices of the MHA.
+Prefix tuning adopts a reparameterization strategy, utilizing a multilayer perceptron (MLP) layer to generate the added vectors instead of optimizing them in a straightforward manner (han2024parameter).
+LoRA (Low-Rank Adaptation) (hu2022lora) injects trainable low-rank decomposition matrices into the query $W_{q}$, key $W_{k}$, and value $W_{v}$ projection matrices of the MHA.
 For a pre-trained weight matrix $W_{0}\in\{W_{q},W_{k},W_{v}\}$, an additive operation $W_{0}+\Delta W=W_{0}+BA$ is performed, where the rank of matrices $B$ and $A$ is considerably smaller than that of $W_{0}$, consequently effectively reducing the number of trainable parameters while preserving the essential characteristics of the original matrices.
-$\mathbf{(IA)^{3}}$ (liu2022few, ) (Infused Adapter by Inhibiting and Amplifying Inner Activation) incorporates three learnable rescaling vectors, $l_{k}$, $l_{v}$, and $l_{ff}$, to rescale the key $K$ and value $V$ in MHA and the inner activations in FFN.
+$\mathbf{(IA)^{3}}$ (liu2022few) (Infused Adapter by Inhibiting and Amplifying Inner Activation) incorporates three learnable rescaling vectors, $l_{k}$, $l_{v}$, and $l_{ff}$, to rescale the key $K$ and value $V$ in MHA and the inner activations in FFN.
 The operations conducted within the MHA can be outlined as follows:
-[…]
+(1) $\text{MHA(x)}=Softmax\Biggl(\frac{Q(l_{k}\odot K^{T})}{\sqrt{d_{k}}}\Biggr)(l_{v}\odot V)$
 where $\odot$ denotes the element-wise multiplication and $d_{k}$ refers to the dimension of key $K$.
 The modifications in the FFN can be defined as:
-[…]
+(2) $\text{FFN(x)}=W_{up}(l_{ff}\odot\sigma(W_{down}x))$
 where $\sigma$ is the activation function within FFN, and $W_{up}$ and $W_{down}$ indicate the weight matrices of the FFN.
-3. Methodology
-Figure 2. Left: architecture of our proposed method for code smell detection. Right: a detailed illustration of the PEFT module structure for the four PEFT methods.
-3.1. Research Goal
-Figure 2 illustrates the architecture of our LLM-based code smell detector using PEFT methods.
-We frame the code smell detection task as a binary classification problem, where the input is a piece of code.
-A sequence classification head is added on top of the LLM to handle the classification task.
-For model fine-tuning, we use four distinct PEFT methods.
-The model then outputs either 00 or $1$, representing the absence or presence of the specific type of code smell.
-3.2. Research Questions
-This study aims to evaluate the effectiveness of state-of-the-art PEFT methods (prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$) in SLMs and LLMs for code smell detection.
-To achieve this goal, we conducted a set of experiments to answer the following Research Questions (RQs):
-RQ1: What is the effectiveness of different PEFT methods and full fine-tuning on various SLMs and LLMs?
-Previous studies have shown remarkable results using PEFT methods to fine-tune SLMs for software engineering tasks like code clone detection and defect detection (liu2023empirical, ; liu2024delving, ; li2024comprehensive, ; wang2023one, ; weyssow2023exploring, ).
-However, these studies have not paid attention to the effectiveness of PEFT methods for code smell detection.
-With this RQ, we aim to compare and validate the effectiveness of PEFT methods across multiple LMs of varying parameter sizes (see Section 3.4) specifically for code smell detection.
-RQ2: What is the effectiveness of PEFT methods with different hyper-parameters?
-When fine-tuning LMs for downstream tasks, the configuration of hyper-parameters plays a crucial role in determining the final results by affecting the number of trainable parameters.
-In prompt tuning and prefix tuning, the number of virtual tokens serves as a critical hyper-parameter (lester2021power, ; li2021prefix, ).
-For LoRA, the key hyper-parameter is the rank $r$ (hu2022lora, ), while in $(IA)^{3}$, it is the selection of activated modules (liu2022few, ).
-Due to our limited computational resources and the popularity of LoRA compared to other PEFT methods (gupta2024beyond, ; zhang2024star, ), we decided to employ LoRA to fine-tune a SLM and an LLM with different $r$ values in a range of {8, 32, 64, 640, 1,280, 2,560} to explore the effect of hyper-parameter settings on PEFT methods for code smell detection.
-RQ3: What is the effectiveness of PEFT methods in low-resource scenarios?
-Low-resource scenarios, characterized by constraints on the amount of training data, pose significant challenges to fine-tuning LMs for specific tasks.
-Due to the scarcity of real-world datasets (zakeri2023systematic, ), these scenarios are common in tasks such as code smell detection.
-Therefore, it is crucial to investigate how the number of training samples influences the effectiveness of PEFT methods for code smell detection.
-RQ4: What is the effectiveness of PEFT methods compared to state-of-the-art code smell detection approaches?
-Existing code smell detectors include traditional heuristic-based and DL-based methods.
-Additionally, the use of general-purpose LLMs with In-Context Learning (ICL) has become increasingly popular in SE tasks (zhou2024large, ; shin2023prompt, ; li2025large, ; xu2024unilog, ), and such models can also be leveraged for code smell detection.
-With this RQ, we aim to assess how LMs utilizing PEFT methods compare against heuristic-based code smell detectors, established DL techniques, and general-purpose LLMs under zero- and few-shot settings for the task of code smell detection.
-3.3. Dataset Construction
-In this section, we elaborate on the process with the following steps of constructing the high-quality and balanced datasets for four types of code smells used to fine-tune both small and large LMs for code smell detection.
-We created our own datasets because most existing datasets only offer code metrics related to the code smells instead of code snippets, which are not usable for our purposes (sharma2021qscored, ; binh2022mlcodesmell, ; madeyski2023detecting, ).
-While some code smell datasets provide valid code samples, they do not cover the specific types of code smells we focus on (e.g., Complex Conditional and Complex Method) (madeyski2020mlcq, ).
-To better align with current software development practices, we constructed new datasets based on trending GitHub projects created within the past six years.
-Furthermore, since data imbalance may negatively impact model performance on code smell detection (khleel2024improving, ; nandini2025improving, ; fontana2016comparing, ; nunes2024tuning, ), we ensured that our constructed datasets contain an equal number of positive and negative samples.
-Figure 3 illustrates the overview of the dataset construction process.
+3. Dataset Construction
+In this section, we elaborate on the process of constructing high-quality and balanced datasets for four types of code smells, rendering them LM-ready and tailored specifically for fine-tuning both small and large LMs.
+As discussed in Section 2.2, existing code smell datasets suffer from several critical limitations, including the absence of raw source code, insufficient coverage of complex smells, and label noise.
+To address these issues and ensure reproducibility, we do not rely on any existing datasets.
+Instead, we create our benchmark from scratch by mining recent Java repositories from GitHub, providing a solid foundation for the subsequent fine-tuning experiments on code smell detection.
+Our construction pipeline is guided by three key objectives: (1) Source-Code Centricity: ensuring all samples contain compilable raw code tokens; (2) Diversity: covering four non-trivial code smells (i.e., Complex Conditional, Complex Method, Feature Envy, Data Class); and (3) Reliability: incorporating a rigorous two-stage manual review to alleviate the “Oracle Problem”.
+Figure 3 provides an overview of the dataset construction process.
 Figure 3. Overview of dataset construction
-3.3.1. Selected Code Smells
+3.1. Selected Code Smells
 In this study, we intend to use LMs with PEFT methods to detect four types of code smells, comprising three at the method level (i.e., Complex Conditional, Complex Method, and Feature Envy) and one at the class level (i.e., Data Class).
 The definitions of these four code smell types are provided below:
-We selected CC, CM, FE, and DC because they are among the most extensively studied code smells in prior research on ML- and DL-based code smell detection (e.g., (sharma2021code, ; ho2023fusion, ; zhang2024code, ; alazba2023deep, ; fontana2016comparing, ; barbez2020machine, ; kim2017finding, ; liu2023deep, ; liu2018deep, ; skipina2024automatic, ; yang2024feature, ; fontana2013code, ; madeyski2020mlcq, )).
-Moreover, these code smells are frequently observed in real-world software projects and are known to significantly impact code quality, maintainability, and the likelihood of introducing defects (fowler1999refactoring, ).
-For example, CC and CM can lead to convoluted control flows that hinder code readability and increase cognitive load on developers, while FE and DC often indicate bad modularization and poor class design, which complicate maintenance and evolution (fowler1999refactoring, ; sjoberg2012questioning, ; khomh2009exploratory, ).
-Another reason for choosing these specific smell types is the availability of detection tools, such as DesigniteJava (designitejava, ) and PMD (pmd, ), which facilitates the collection of candidate code samples potentially exhibiting these smells.
+  • Complex Conditional (CC): This code smell occurs when a conditional statement is excessively intricate (sharma2017house).
+  • Complex Method (CM): This code smell occurs when a method exhibits high cyclomatic complexity (sharma2017house).
+  • Feature Envy (FE): This code smell occurs when a method is more interested in a class other than the one it actually is in (fowler1999refactoring; fontana2016comparing).
+  • Data Class (DC): This code smell occurs when a class contains only fields and basic methods for accessing them (fowler1999refactoring; fontana2016comparing).
+We selected CC, CM, FE, and DC because they are among the most extensively studied code smells in prior research on ML- and DL-based code smell detection (e.g., (sharma2021code; ho2023fusion; zhang2024code; alazba2023deep; fontana2016comparing; barbez2020machine; kim2017finding; liu2023deep; liu2018deep; skipina2024automatic; yang2024feature; fontana2013code; madeyski2020mlcq)).
+Moreover, these code smells are frequently observed in real-world software projects and are known to significantly impact code quality, maintainability, and the likelihood of introducing defects (fowler1999refactoring).
+For example, CC and CM can lead to convoluted control flows that hinder code readability and increase cognitive load on developers, while FE and DC often indicate bad modularization and poor class design, which complicate maintenance and evolution (fowler1999refactoring; sjoberg2012questioning; khomh2009exploratory).
+Another reason for choosing these specific smell types is the availability of detection tools, such as DesigniteJava (designitejava) and PMD (pmd), which facilitates the collection of candidate code samples potentially exhibiting these smells.
 We excluded overly simple smells (e.g., Long Method) that can be easily identified using basic heuristics.
-In contrast, detecting the four selected code smells requires LLMs to capture subtle features, consider aspects such as cohesion and coupling, and analyze the overall structure of a code sample (sharma2021code, ), which expands the ambition and enhances the applicability of this study.
+In contrast, detecting the four selected code smells requires LLMs to capture subtle features, consider aspects such as cohesion and coupling, and analyze the overall structure of a code sample (sharma2021code), which expands the ambition and enhances the applicability of this study.
 At the same time, identifying these selected code smells allows us to assess the effectiveness of LMs fine-tuned with PEFT methods in both method-level and class-level code understanding.
 CC, CM, and FE are generally considered method-level code smells, while DC represents a class-level code smell.
-However, detecting FE requires comprehension of not only the focal method but also its surrounding class context and inter-class dependencies (fowler1999refactoring, ; tsantalis2009indentification, ; liu2023deep, ; alkharabsheh2021analysing, ).
+However, detecting FE requires comprehension of not only the focal method but also its surrounding class context and inter-class dependencies (fowler1999refactoring; tsantalis2009indentification; liu2023deep; alkharabsheh2021analysing).
 Therefore, it also demands class-level code understanding.
-The selection of these four code smells allows our study to address a key gap in existing research related to PEFT methods (li2024comprehensive, ; liu2023empirical, ; wang2023one, ; liu2024delving, ; weyssow2023exploring, ), which has primarily concentrated on the effectiveness of PEFT techniques on method-level code-related tasks.
-3.3.2. Java Repository Selection
+The selection of these four code smells allows our study to address a key gap in existing research related to PEFT methods (li2024comprehensive; liu2023empirical; wang2023one; liu2024delving; weyssow2023exploring), which has primarily concentrated on the effectiveness of PEFT techniques on method-level code-related tasks.
+3.2. Java Repository Selection
 To construct a code smell dataset that accurately represents modern coding practices, we selected Java repositories created in the past six years from GitHub.
@@
 By employing these stringent filtering criteria, we obtained a list of 401 GitHub repositories coded in Java.
-3.3.3. Potential Code Smell Detection
-DesigniteJava is an open-source tool capable of detecting 18 design smells and 10 implementation smells, whereas PMD is a rule-based static analysis tool that identifies common programming issues, including code smells.
-Both of them have been extensively used in prior studies (sharma2021qscored, ; shah2023mining, ; barbosa2020revealing, ; fernandes2016review, ; aniche2018code, ; wu2024ismell, ).
+3.3. Potential Code Smell Detection
+DesigniteJava (designitejava) is an open-source tool capable of detecting 18 design smells and 10 implementation smells, whereas PMD (pmd) is a rule-based static analysis tool that identifies common programming issues, including code smells.
+Both of them have been extensively used in prior studies (sharma2021qscored; shah2023mining; barbosa2020revealing; fernandes2016review; aniche2018code; wu2024ismell).
 In this paper, we leveraged DesigniteJava to detect potential instances of CC, CM, and FE, and used PMD to detect potential instances of DC.
@@
 These inaccuracies stem from the rule-based and heuristic nature of DesigniteJava and PMD, which relies on predefined criteria that may not capture all nuances of the code.
-3.3.4. Data Preparation
-CodeSplitJava (codesplitjava, ) is a utility program designed to extract methods or classes from Java source code and organize them into separate files.
-We employed CodeSplitJava to parse the 401 Java repositories collected in Section 3.3.2, generating individual code files for each method or class.
+3.4. Data Preparation
+CodeSplitJava (codesplitjava) is a utility program designed to extract methods or classes from Java source code and organize them into separate files.
+We employed CodeSplitJava to parse the 401 Java repositories collected in Section 3.2, generating individual code files for each method or class.
 The file paths reflect the hierarchical structure of the original code (i.e., namespaces and packages are converted into nested folders).
-By aligning the list of detected code smells obtained in Section 3.3.3 with the paths of the code files generated by CodeSplitJava, we were able to collect code snippets that were labeled by heuristic-based detectors as either containing code smells or being free of them.
+By aligning the list of detected code smells obtained in Section 3.3 with the paths of the code files generated by CodeSplitJava, we were able to collect code snippets that were labeled by heuristics-based detectors as either containing code smells or being free of them.
 Table 1.
@@
 Note that although FE is a method-level code smell, we retained the entire class that contains the method as the sample.
-This decision aligns with the definition of FE (see Section 3.3.1), which emphasizes the method’s interaction with its enclosing class.
+This decision aligns with the definition of FE (see Section 3.1), which emphasizes the method’s interaction with its enclosing class.
 Providing the full class context ensures that the code sample contains sufficient structural and semantic information for FE detection.
-3.3.5. Manual Review
+3.5. Manual Review
 After the above steps, we obtained two sets of samples for each code smell: a potential positive set and a potential negative set, with the latter being significantly larger in size.
@@
 Figure 4. The process of two-stage manual review in dataset construction
-Table 2. Number of positive and negative samples after manual review
-Code Smell #Positive #Negative Total
-Complex Conditional 4,604 4,604 9,208
-Complex Method 5,748 5,748 11,496
-Feature Envy 307 307 614
-Data Class 6,922 6,922 13,844
+Table 2. Number of positive and negative samples at each stage of manual review
+Code Smell Stage 1 Stage 2 Total
+#Positive #Negative #Positive #Negative #Positive #Negative
+Complex Conditional 3,715 3,186 889 1,418 4,604 4,604
+Complex Method 5,699 4,120 49 1,628 5,748 5,748
+Feature Envy 184 153 123 154 307 307
+Data Class 5,904 1,217 1,018 5,705 6,922 6,922
 In order to ensure the consistency of data annotation, a pilot review was performed by the first and the fourth authors before conducting the formal manual review process.
-For each code smell type (i.e., CC, CM, FE, and DC), 30 potential positive samples and 30 potential negative samples labeled by heuristic-based tools were randomly selected.
+For each code smell type (i.e., CC, CM, FE, and DC), 30 potential positive samples and 30 potential negative samples labeled by heuristics-based tools were randomly selected.
 The two authors independently examined these samples and then compared their labeling results.
-Cohen’s Kappa coefficient (jacob1960coefficient, ) was used to assess inter-rater agreement between the two authors.
+Cohen’s Kappa coefficient (jacob1960coefficient) was used to assess inter-rater agreement between the two authors.
 The Kappa coefficient for CC was 0.83, for CM was 0.80, for FE was 0.77, and for DC was 0.87.
@@
 Following the pilot review, the two authors held a discussion to reconcile discrepancies and formalize the labeling criteria below:
-  • CC: Labeled as positive if the method contains at least one conditional statement that either (i) has a nesting depth of three or more, making the method hard to read, or (ii) includes compound boolean expressions involving three or more operands, exhibiting complex logic that makes the condition challenging to evaluate and understand.
+  • CC: Labeled as positive if the method contains at least one conditional statement that satisfies (i) a nesting depth of three or more or (ii) compound boolean expressions involving three or more operands, and simultaneously (iii) exhibits semantically complex logic that makes the condition difficult to evaluate and understand.
   • CM: Labeled as positive if the method simultaneously satisfies both (i) three or more execution paths with complex control flow, and (ii) an implementation that is difficult to comprehend or follow.
@@
   • DC: Labeled as positive if the class contains (i) primarily fields accompanied by trivial getters/setters (at least 80% of its fields), and (ii) lacks methods that implement significant logic or domain behavior, indicating its role as a passive data holder.
-These criteria were grounded in established definitions from prior studies (sharma2017house, ; fowler1999refactoring, ; fontana2016comparing, ; pietrzak2006leveraging, ) and refined based on insights from our pilot study.
+These criteria were grounded in established definitions from prior studies (sharma2017house; fowler1999refactoring; fontana2016comparing; pietrzak2006leveraging) and refined based on insights from our pilot study.
 During the formal manual review, the two authors collaboratively labeled all samples (i.e., each sample was assigned to one of the two authors), which took approximately 300 person-hours in total.
@@
 In such cases, all three authors discussed the sample until consensus was reached, ensuring annotation reliability and consistency.
-The final count of positive and negative samples, following the manual review, is presented in Table 2.
-3.3.6. Data Split
-In this step, we split the code samples of CC, CM, FE, and DC into training, validation, and test sets following an 8:1:1 ratio, as employed in previous studies (yang2024federated, ; liu2023benchmarking, ).
+Table 2 summarizes the numbers of positive and negative samples collected at each stage of the manual review process.
+3.6. Data Split
+In this step, we split the code samples of CC, CM, FE, and DC into training, validation, and test sets following an 8:1:1 ratio, as employed in previous studies (yang2024federated; liu2023benchmarking).
 For each type of code smell, we ensured that the number of positive and negative samples remains equal across the training, validation, and test sets, thereby maintaining class balance throughout the entire dataset.
@@
 #Negative 5,538 692 692 6,922
+4. Experimental Setup
+4.1. Task Formulation
+We formulate code smell detection as a binary classification task.
+Given a code snippet $\mathcal{C}={c_{1},c_{2},\dots,c_{n}}$ consisting of a sequence of tokens, the goal is to learn a mapping function $f:\mathcal{C}\to\mathcal{Y}$, where $\mathcal{Y}\in{0,1}$.
+Here, $1$ indicates the presence of a specific code smell, and 0 denotes its absence.
+As illustrated in Figure 5, we employ a Language Model (LM) as the backbone for this function $f$.
+To adapt the LM for this specific classification task, we append a sequence classification head on top of the model and optimize it using four Parameter-Efficient Fine-Tuning (PEFT) techniques.
+4.2. Research Questions
+This study aims to evaluate the effectiveness of state-of-the-art PEFT methods (prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$) in SLMs and LLMs for code smell detection.
+To achieve this goal, we conducted a set of experiments to answer the following Research Questions (RQs):
+RQ1: What is the effectiveness of different PEFT methods and full fine-tuning on various SLMs and LLMs?
+Previous studies have shown remarkable results using PEFT methods to fine-tune SLMs for software engineering tasks like code clone detection and defect detection (liu2023empirical; liu2024delving; li2024comprehensive; wang2023one; weyssow2023exploring).
+However, these studies have not paid attention to the effectiveness of PEFT methods for code smell detection.
+With this RQ, we aim to compare and validate the effectiveness of PEFT methods across multiple LMs of varying parameter sizes (see Section 4.3) specifically for code smell detection.
+RQ2: What is the effectiveness of PEFT methods with different hyper-parameters?
+When fine-tuning LMs for downstream tasks, the configuration of hyper-parameters plays a crucial role in determining the final results by affecting the number of trainable parameters.
+In prompt tuning and prefix tuning, the number of virtual tokens serves as a critical hyper-parameter (lester2021power; li2021prefix).
+For LoRA, the key hyper-parameter is the rank $r$ (hu2022lora), while in $(IA)^{3}$, it is the selection of activated modules (liu2022few).
+Due to our limited computational resources and the popularity of LoRA compared to other PEFT methods (gupta2024beyond; zhang2024star), we decided to employ LoRA to fine-tune a SLM and an LLM with different $r$ values in a range of {8, 32, 64, 640, 1,280, 2,560} to explore the effect of hyper-parameter settings on PEFT methods for code smell detection.
+RQ3: What is the effectiveness of PEFT methods in low-resource scenarios?
+Low-resource scenarios, characterized by constraints on the amount of training data, pose significant challenges to fine-tuning LMs for specific tasks.
+Due to the scarcity of real-world datasets (zakeri2023systematic), these scenarios are common in tasks such as code smell detection.
+Therefore, it is crucial to investigate how the number of training samples influences the effectiveness of PEFT methods for code smell detection.
+RQ4: What is the effectiveness of PEFT methods compared to state-of-the-art code smell detection approaches?
+Existing code smell detectors include traditional heuristics-based and DL-based methods.
+Additionally, the use of general-purpose LLMs with In-Context Learning (ICL) has become increasingly popular in SE tasks (zhou2024large; shin2023prompt; li2025large; xu2024unilog), and such models can also be leveraged for code smell detection.
+With this RQ, we aim to assess how LMs utilizing PEFT methods compare against heuristics-based code smell detectors, established DL techniques, and general-purpose LLMs under zero- and few-shot settings for the task of code smell detection.
+Figure 5. Left: architecture of our proposed method for code smell detection. Right: a detailed illustration of the PEFT module structure for the four PEFT methods.
 4.3. Base Model Selection
@@
 Consistent with the definition adopted by Weyssow et al.
-(weyssow2023exploring, ), we classified models with 1B or more parameters as LLMs, and those with fewer parameters as SLMs.
-Table 4. Selected LMs and their corresponding parameter size and architecture
-Model #Params Architecture
-SLMs CodeBERT 125M Encoder-only
-GraphCodeBERT 125M
-UnixCoder 126M
-CodeT5 220M Encoder-decoder
-LLMs StarCoderBase-1B 1.1B Decoder-only
-StarCoderBase-3B 3B
-DeepSeek-Coder-1.3B 1.3B
-DeepSeek-Coder-6.7B 6.7B
-CodeLlama-7B 7B
+(weyssow2023exploring), we classified models with 1B or more parameters as LLMs, and those with fewer parameters as SLMs.
+Table 4. Selected LMs and their corresponding parameter sizes, context window limits, and architecture types
+Model #Params Context Window Architecture
+SLMs CodeBERT 125M 512 Encoder-only
+GraphCodeBERT 125M 512
+UnixCoder 125M 1,024
+CodeT5 220M 512 Encoder-decoder
+LLMs StarCoderBase-1B 1.1B 1,024 Decoder-only
+StarCoderBase-3B 3B 1,024
+DeepSeek-Coder-1.3B 1.3B 1,024
+DeepSeek-Coder-6.7B 6.7B 1,024
+CodeLlama-7B 6.7B 1,024
 First, the selected LMs should be open-source as PEFT methods require access to their parameters.
@@
 Fourth, the selected models should either be popular SLMs or recent advanced LLMs to ensure they represent state-of-the-art capabilities.
-Based on the above criteria, we employed 4 SLMs (i.e., CodeBERT (feng2020codebert, ), GraphCodeBERT (guo2021graphcodebert, ), UnixCoder (guo2022unixcoder, ), and CodeT5 (wang2021codet5, )) and 5 LLMs (StarCoderBase-1B (li2023starcoder, ), StarCoderBase-3B (li2023starcoder, ), DeepSeek-Coder-1.3B (guo2024deepseekcoder, ), DeepSeek-Coder-6.7B (guo2024deepseekcoder, ), and CodeLlama-7B (roziere2024codellama, )) for our experiments.
+Based on the above criteria, we employed 4 SLMs (i.e., CodeBERT (feng2020codebert), GraphCodeBERT (guo2021graphcodebert), UnixCoder (guo2022unixcoder), and CodeT5 (wang2021codet5)) and 5 LLMs (StarCoderBase-1B (li2023starcoder), StarCoderBase-3B (li2023starcoder), DeepSeek-Coder-1.3B (guo2024deepseekcoder), DeepSeek-Coder-6.7B (guo2024deepseekcoder), and CodeLlama-7B (roziere2024codellama)) for our experiments.
 To ensure a comprehensive analysis, we deliberately included models representing all three major Transformer architectures: encoder-only, encoder-decoder, and decoder-only.
-Although decoder-only models were originally designed for generative tasks (vaswani2017attention, ; radford2019language, ), prior studies have shown their competitive performance in sequence classification tasks as well (wolf2020transformers, ; kariyappa2024progressive, ).
+Although decoder-only models were originally designed for generative tasks (vaswani2017attention; radford2019language), prior studies have shown their competitive performance in sequence classification tasks as well (wolf2020transformers; kariyappa2024progressive).
 Therefore, we also included decoder-only models as base models in our study.
 Additionally, all the models we selected have demonstrated encouraging results across various code understanding tasks like code clone detection, defect detection, etc.
-(liu2024vul, ; karmakar2021what, ; lu2021codexglue, ; li2022automating, ; tang2023csgvd, ; zeng2022an, ; chen2023diversevul, ; wang2023one, ; fu2022vulrepair, ; nguyen2024empirical, ; su2024distilled, ).
-Table 4 presents the details of the 9 base models, including their parameter sizes and architectural types.
+(liu2024vul; karmakar2021what; lu2021codexglue; li2022automating; tang2023csgvd; zeng2022an; chen2023diversevul; wang2023one; fu2022vulrepair; nguyen2024empirical; su2024distilled).
+Table 4 presents the details of the 9 base models, including their parameter sizes, context window limits, and architectural types.
+The input sequences of CodeBERT, GraphCodeBERT, and CodeT5 were truncated at 512 tokens, which corresponds to their maximum supported input length.
+For all other models, inputs were uniformly truncated at 1,024 tokens.
+4.4. Baselines
+In RQ4, to comprehensively evaluate the effectiveness of PEFT-tuned LMs for code smell detection, we compare them against three distinct categories of baselines: traditional heuristics-based detectors, existing Deep Learning (DL)-based approaches, and state-of-the-art general-purpose LLMs using In-Context Learning (ICL).
+4.4.1. Heuristics-based detectors
+We select two prominent heuristics-based detectors that rely on predefined metrics and rule-based algorithms to identify code smells:
+  • DesigniteJava (designitejava): An industrial-strength tool that detects architecture, design, and implementation smells by analyzing source code structure and calculating metrics (e.g., cyclomatic complexity and class coupling).
+In this study, we deploy DesigniteJava to detect CC, CM, and FE.
+  • PMD (pmd): A widely adopted static analyzer that flags potential inefficiencies and bad practices in Java code through rigorous rule matching.
+In this study, we specifically task PMD with identifying DC instances.
+These tools serve as foundational reference points in our study and are extensively cited in prior literature (sharma2021qscored; shah2023mining; barbosa2020revealing; fernandes2016review; aniche2018code; wu2024ismell).
+Furthermore, since these tools were utilized in our initial code smell collection phase (see Section 3.3), comparing against them provides critical insight into how much improvement our PEFT-based approach offers over the original labeling sources.
+4.4.2. DL-based Approaches
+We incorporate two representative DL models as baselines to evaluate the performance of established neural architectures:
+  • DeepSmells (ho2023fusion): A model that utilizes a 1D-Convolutional Neural Network (1D-CNN) followed by a Long Short-Term Memory (LSTM) network to learn representations from tokenized source code sequences.
+  • AE-Dense (sharma2021code): An Autoencoder-based model designed to learn dense vector representations of code instances in an unsupervised manner.
+Collectively, these comparative DL models represent the state-of-the-art in leveraging sequential patterns and latent feature learning for code smell detection, respectively.
+4.4.3. LLMs with In-Context Learning (ICL)
+To benchmark the performance of our fine-tuned models against the “prompting” paradigm without parameter updates for code smell detection, we leverage two powerful general-purpose LLMs.
+These models exemplify the latest advances in commercial and open-weight architectures under the ICL paradigm:
+  • GPT-4o-mini (hurst2024gpt4o): One of the most advanced proprietary LLMs developed by OpenAI, known for its superior reasoning capabilities and code understanding in lightweight deployments.
+  • DeepSeek-v3 (liu2024deepseek): A current cutting-edge open-source mixture-of-experts (MoE) model (671B parameters), which serves as a strong reference for the performance of large-scale open-weight models.
+Both models have demonstrated exceptional capabilities across natural language and programming tasks (rasheed2024task; bruni2025benchmarking; liu2024deepseek; jahin2025unveiling).
+We assess their effectiveness in code smell detection under three ICL settings: zero-shot, few-shot, and a criteria-informed setting that provides explicit labeling rules (refer to Section 5.4 for details).
 4.5. Metrics
-We use metrics that align with previous research (sharma2021code, ; ho2023fusion, ; pecorelli2019comparing, ) to maintain consistency in evaluating the effectiveness of PEFT methods for code smell detection on LLMs.
+We use metrics that align with previous research (sharma2021code; ho2023fusion; pecorelli2019comparing) to maintain consistency in evaluating the effectiveness of PEFT methods for code smell detection on LLMs.
 Specifically, we utilize precision, recall, F1-score, and MCC (Matthews Correlation Coefficient) for assessing the performance of detecting CC and CM.
-We include MCC because, unlike other metrics like F1-score, MCC accounts for true negative instances (yao2020assessing, ).
+We include MCC because, unlike other metrics like F1-score, MCC accounts for true negative instances (yao2020assessing).
 These metrics are calculated using the following formulas:
@@
 (4) $\text{Recall}=\frac{TP}{TP+FN}$
-[…]
+(5) $\text{F1-Score}=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$
 (6) $\text{MCC}=\frac{TP\times TN-FP\times FN}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}}$
-where $TP$, $TN$, $FP$, $FN$ represent true positives, true negatives, false positives, and false negatives, respectively.
+where $TP$, TNTN, FPFP, FNFN represent true positives, true negatives, false positives, and false negatives, respectively.
 4.6. Implementation Details
 The code for our experiments is implemented by the PEFT library [^4] https://huggingface.co/docs/peft/index from Hugging Face.
-After reviewing prior studies (liu2023empirical, ; liu2024delving, ; li2024comprehensive, ; wang2023one, ; weyssow2023exploring, ) and observing the results of pilot experiments, we set the learning rate to $3e^{-4}$ for all PEFT methods and $1e^{-5}$ for full fine-tuning the SLMs.
-Throughout all the fine-tuning experiments, we utilized the AdamW optimizer (loshchilov2017decoupled, ) and fixed the Transformer input window size at 1,024.
+After reviewing prior studies (liu2023empirical; liu2024delving; li2024comprehensive; wang2023one; weyssow2023exploring) and observing the results of pilot experiments, we set the learning rate to $3e^{-4}$ for all PEFT methods and $1e^{-5}$ for full fine-tuning the SLMs.
+Throughout all the fine-tuning experiments, we utilized the AdamW optimizer (loshchilov2017decoupled) and fixed the Transformer input window size at 1,024.
 The choice to limit the input length to 1,024 tokens was driven by hardware constraints.
-As the GPU memory consumption of transformer-based models grows roughly quadratically with input length (tay2022efficient, ), this setting enables all selected base models to train within available GPU memory capacity.
+As the GPU memory consumption of transformer-based models grows roughly quadratically with input length (tay2022efficient), this setting enables all selected base models to train within available GPU memory capacity.
 All the models were trained for 10 epochs, and we selected the epochs at which models exhibit minimum loss on the validation sets for inference.
-Table 5 shows the hyperparameter settings of different PEFT methods.
-We adopted a batch size of 1 for all experiments to avoid GPU out-of-memory issues.
+Table 5 shows the hyper-parameter settings of different PEFT methods.
+Due to resource constraints, the batch size was set to 1 for LLMs with large parameter sizes to avoid out-of-memory issues.
+To ensure consistency and comparability, we adopted a batch size of 1 for all experiments.
 We used a fixed random seed (42) across all runs to ensure experimental reproducibility.
 For prompt tuning and prefix tuning, we used 20 virtual tokens, which are learnable vectors prepended to the input and optimized during training.
-This setting is widely used in the literature and offers a balance between representational capacity and computational cost (li2021prefix, ; weyssow2023exploring, ).
-As for LoRA, we set the rank $r$ to 8, the scaling factor $\alpha$ to 16, and the dropout rate to 0.1 to prevent overfitting, following configurations reported in (liu2024delving, ).
+This setting is widely used in the literature and offers a balance between representational capacity and computational cost (li2021prefix; weyssow2023exploring).
+As for LoRA, we set the rank $r$ to 8, the scaling factor $\alpha$ to 16, and the dropout rate to 0.1 to prevent overfitting, following configurations reported in (liu2024delving).
 For both LoRA and ${IA}^{3}$, we applied parameter-efficient adaptation to the core components of the Transformer architecture, including the query ($Q$), key ($K$), and value ($V$) projections in the attention layers.
-In the case of $(\text{IA})^{3}$, we also included the Feed-Forward Network (FFN) for adaptation (see Section 2.3).
+In the case of $(\text{IA})^{3}$, we also included the Feed-Forward Network (FFN) for adaptation (see Section 2.3.2).
 We executed all models on a single Linux GPU server equipped with a 40GB NVIDIA A100 GPU.
-Additionally, we have released our datasets and source code online for replication purpose (replicationPackage, ) where more details can be found.
-Table 5. Hyperparameter settings of different PEFT methods and full fine-tuning
-PEFT Method Hyperparameter Setting Batch Size Seed
+Additionally, we have released our datasets and source code online for replication purpose (replicationPackage) where more details can be found.
+Table 5. Hyper-parameter settings of different PEFT methods and full fine-tuning
+PEFT Method Hyper-parameter Setting Batch Size Seed
 Prompt Tuning Number of virtual tokens: 20 1 42
@@
 LoRA Rank $r$: 8
-$alpha$: 16
-$dropout$: 0.1
+alphaalpha: 16
+dropoutdropout: 0.1
 Target modules: $Q$, $K$, and $V$
 $(IA)^{3}$ Target modules: $Q$, $K$, $V$, and FFN
-Full Fine-tuing None
+Full Fine-tuning None
 5. Results
-In this section, we first illustrate the experiment setup for each RQ formulated in Section 3, and we then present the results and findings for the four RQs.
+In this section, we present the results and findings for the four RQs formulated in Section 4.2.
+For each RQ, we first specify the particular experimental configuration and execution strategy based on the general setup in Section 4, followed by a detailed analysis of the results.
 5.1. RQ1: What is the effectiveness of different PEFT methods and full fine-tuning on various SLMs and LLMs?
-4.1.1. Experiment Setup
 Since our objective is to demonstrate the effectiveness of PEFT methods on both small and large LMs for code smell detection, we conducted a series of experiments involving four PEFT methods (i.e., prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$) and a group of models with different parameter sizes.
@@
 Table 6. Results of small LMs and LLMs with different PEFT methods on detecting CC and CM (bold: highest values per model, underline: highest values among all models)
-Model Method CC CM
+Model Method #Params $C$ CM
 F1 MCC F1 MCC
-SLMs CodeBERT Full FT 70.54% 44.41% 80.57% 61.19%
-Prompt Tuning 62.38% 25.90% 76.13% 54.00%
-Prefix Tuning 63.88% 27.87% 81.33% 62.94%
-LoRA 73.04% 46.10% 71.00% 49.79%
-$(IA)^{3}$ 69.44% 40.49% 81.32% 62.96%
-GraphCodeBERT Full FT 74.01% 48.08% 80.66% 61.33%
-Prompt Tuning 65.24% 30.81% 78.90% 58.60%
-Prefix Tuning 69.82% 39.97% 79.18% 60.46%
-LoRA 73.15% 46.30% 80.03% 60.93%
-$(IA)^{3}$ 72.50% 45.00% 80.00% 60.45%
-UnixCoder Full FT 73.04% 46.10% 82.39% 64.89%
-Prompt Tuning 69.35% 38.70% 78.66% 57.95%
-Prefix Tuning 71.63% 43.26% 82.71% 65.84%
-LoRA 72.19% 46.04% 73.18% 51.52%
-$(IA)^{3}$ 73.47% 46.98% 82.23% 64.47%
-CodeT5 Full FT 73.01% 46.19% 81.13% 62.74%
-LoRA 33.33% 0.00% 83.29% -2.95%
-$(IA)^{3}$ 43.38% -4.05% 40.32% 16.70%
-LLMs StarCoderBase-1B Prompt Tuning 71.07% 42.21% 80.05% 60.11%
-Prefix Tuning 70.10% 40.23% 76.88% 54.01%
-LoRA 41.41% -1.62% 45.41% 7.01%
-$(IA)^{3}$ 74.27% 49.01% 80.90% 62.01%
-StarCoderBase-3B Prompt Tuning 69.12% 38.29% 81.87% 63.85%
-Prefix Tuning 67.81% 35.69% 78.25% 56.92%
-LoRA 40.13% 7.93% 68.41% 46.00%
-$(IA)^{3}$ 74.68% 49.96% 82.07% 64.81%
-Deepseek-Coder-1.3B Prompt Tuning 70.21% 40.44% 81.77% 63.75%
-Prefix Tuning 69.38% 39.62% 79.06% 58.36%
-LoRA 69.44% 40.49% 78.42% 57.96%
-$(IA)^{3}$ 72.50% 45.01% 81.87% 63.82%
-Deepseek-Coder-6.7B Prompt Tuning 67.93% 35.87% 81.19% 63.11%
-Prefix Tuning 68.80% 37.62% 80.09% 60.57%
-LoRA 69.98% 40.05% 81.34% 63.63%
-$(IA)^{3}$ 70.84% 42.36% 83.34% 66.88%
-CodeLlama-7B Prompt Tuning 68.68% 37.92% 81.25% 63.43%
-Prefix Tuning 53.41% 15.95% 80.77% 62.08%
-LoRA 70.12% 40.70% 80.70% 62.59%
-$(IA)^{3}$ 70.48% 42.32% 83.00% 66.17%
+SLMs CodeBERT Full FT 125M 70.54% 44.41% 80.57% 61.19%
+Prompt Tuning 0.61M 62.38% 25.90% 76.13% 54.00%
+Prefix Tuning 0.96M 63.88% 27.87% 81.33% 62.94%
+LoRA 1.03M 73.04% 46.10% 71.00% 49.79%
+$(IA)^{3}$ 0.67M 69.44% 40.49% 81.32% 62.96%
+GraphCodeBERT Full FT 125M 74.01% 48.08% 80.66% 61.33%
+Prompt Tuning 0.61M 65.24% 30.81% 78.90% 58.60%
+Prefix Tuning 0.96M 69.82% 39.97% 79.18% 60.46%
+LoRA 1.03M 73.15% 46.30% 80.03% 60.93%
+$(IA)^{3}$ 0.67M 72.50% 45.00% 80.00% 60.45%
+UnixCoder Full FT 125M 73.04% 46.10% 82.39% 64.89%
+Prompt Tuning 0.61M 69.35% 38.70% 78.66% 57.95%
+Prefix Tuning 0.96M 71.63% 43.26% 82.71% 65.84%
+LoRA 1.03M 72.19% 46.04% 73.18% 51.52%
+$(IA)^{3}$ 0.67M 73.47% 46.98% 82.23% 64.47%
+CodeT5 Full FT 220M 73.01% 46.19% 81.13% 62.74%
+LoRA 1.33M 33.33% 0.00% 83.29% -2.95%
+$(IA)^{3}$ 0.16M 43.38% -4.05% 40.32% 16.70%
+LLMs StarCoderBase-1B Prompt Tuning 0.05M 71.07% 42.21% 80.05% 60.11%
+Prefix Tuning 1.97M 70.10% 40.23% 76.88% 54.01%
+LoRA 3.59M 41.41% -1.62% 45.41% 7.01%
+$(IA)^{3}$ 0.31M 74.27% 49.01% 80.90% 62.01%
+StarCoderBase-3B Prompt Tuning 0.06M 69.12% 38.29% 81.87% 63.85%
+Prefix Tuning 4.06M 67.81% 35.69% 78.25% 56.92%
+LoRA 7.38M 40.13% 7.93% 68.41% 46.00%
+$(IA)^{3}$ 0.62M 74.68% 49.96% 82.07% 64.81%
+Deepseek-Coder-1.3B Prompt Tuning 0.05M 70.21% 40.44% 81.77% 63.75%
+Prefix Tuning 1.97M 69.38% 39.62% 79.06% 58.36%
+LoRA 2.36M 69.44% 40.49% 78.42% 57.96%
+$(IA)^{3}$ 0.28M 72.50% 45.01% 81.87% 63.82%
+Deepseek-Coder-6.7B Prompt Tuning 0.09M 67.93% 35.87% 81.19% 63.11%
+Prefix Tuning 5.25M 68.80% 37.62% 80.09% 60.57%
+LoRA 6.30M 69.98% 40.05% 81.34% 63.63%
+$(IA)^{3}$ 0.75M 70.84% 42.36% 83.34% 66.88%
+CodeLlama-7B Prompt Tuning 0.09M 68.68% 37.92% 81.25% 63.43%
+Prefix Tuning 5.25M 53.41% 15.95% 80.77% 62.08%
+LoRA 6.30M 70.12% 40.70% 80.70% 62.59%
+$(IA)^{3}$ 0.75M 70.48% 42.32% 83.00% 66.17%
 Table 7. Results of small LMs and LLMs with different PEFT methods on detecting FE and DC (bold: highest values per model, underline: highest values among all models)
-Model Method FE DC
+Model Method #Params FE DC
 F1 MCC F1 MCC
-SLMs CodeBERT Full FT 74.09% 48.80% 81.98% 65.60%
-Prompt Tuning 56.17% 13.07% 81.56% 64.57%
-Prefix Tuning 62.82% 25.93% 81.81% 65.43%
-LoRA 67.44% 36.17% 83.17% 68.48%
-$(IA)^{3}$ 66.49% 38.48% 83.30% 68.18%
-GraphCodeBERT Full FT 77.40% 54.95% 82.66% 67.42%
-Prompt Tuning 66.88% 37.49% 83.75% 68.95%
-Prefix Tuning 74.19% 48.39% 89.26% 79.26%
-LoRA 72.23% 46.36% 85.85% 72.40%
-$(IA)^{3}$ 69.84% 45.48% 85.58% 72.56%
-UnixCoder Full FT 77.04% 56.76% 83.51% 69.49%
-Prompt Tuning 77.32% 55.30% 83.68% 68.74%
-Prefix Tuning 75.80% 51.64% 83.89% 69.28%
-LoRA 70.97% 41.94% 83.60% 68.65%
-$(IA)^{3}$ 73.95% 49.32% 83.96% 69.44%
-CodeT5 Full FT 66.49% 38.48% 83.46% 69.12%
-LoRA 54.07% 10.02% 70.03% 49.06%
-$(IA)^{3}$ 54.65% 9.76% 46.79% -2.66%
-LLMs StarCoderBase-1B Prompt Tuning 72.23% 46.36% 84.66% 71.21%
-Prefix Tuning 75.65% 52.30% 85.44% 72.19%
-LoRA 78.58% 60.68% 79.30% 58.85%
-$(IA)^{3}$ 73.95% 49.32% 84.96% 71.71%
-StarCoderBase-3B Prompt Tuning 68.96% 39.74% 83.63% 68.47%
-Prefix Tuning 67.61% 35.78% 86.56% 74.06%
-LoRA 70.85% 42.29% 83.70% 68.64%
-$(IA)^{3}$ 69.35% 38.73% 84.90% 71.46%
-DeepSeek-Coder-1.3B Prompt Tuning 78.98% 58.34% 84.17% 69.99%
-Prefix Tuning 66.88% 37.49% 82.44% 64.90%
-LoRA 78.98% 58.34% 82.72% 67.68%
-$(IA)^{3}$ 83.87% 67.74% 85.84% 72.57%
-DeepSeek-Coder-6.7B Prompt Tuning 78.89% 58.83% 85.56% 72.72%
-Prefix Tuning 80.56% 61.81% 83.03% 68.14%
-LoRA 77.32% 55.30% 83.43% 68.64%
-$(IA)^{3}$ 77.32% 55.30% 85.34% 72.34%
-CodeLlama-7B Prompt Tuning 72.52% 45.37% 85.49% 72.55%
-Prefix Tuning 63.57% 30.67% 82.30% 64.61%
-LoRA 77.32% 55.30% 83.10% 68.31%
-$(IA)^{3}$ 74.19% 48.39% 83.54% 68.45%
+SLMs CodeBERT Full FT 125M 74.09% 48.80% 81.98% 65.60%
+Prompt Tuning 0.61M 56.17% 13.07% 81.56% 64.57%
+Prefix Tuning 0.96M 62.82% 25.93% 81.81% 65.43%
+LoRA 1.03M 67.44% 36.17% 83.17% 68.48%
+$(IA)^{3}$ 0.67M 66.49% 38.48% 83.30% 68.18%
+GraphCodeBERT Full FT 125M 77.40% 54.95% 82.66% 67.42%
+Prompt Tuning 0.61M 66.88% 37.49% 83.75% 68.95%
+Prefix Tuning 0.96M 74.19% 48.39% 89.26% 79.26%
+LoRA 1.03M 72.23% 46.36% 85.85% 72.40%
+$(IA)^{3}$ 0.67M 69.84% 45.48% 85.58% 72.56%
+UnixCoder Full FT 125M 77.04% 56.76% 83.51% 69.49%
+Prompt Tuning 0.61M 77.32% 55.30% 83.68% 68.74%
+Prefix Tuning 0.96M 75.80% 51.64% 83.89% 69.28%
+LoRA 1.03M 70.97% 41.94% 83.60% 68.65%
+$(IA)^{3}$ 0.67M 73.95% 49.32% 83.96% 69.44%
+CodeT5 Full FT 220M 66.49% 38.48% 83.46% 69.12%
+LoRA 1.33M 54.07% 10.02% 70.03% 49.06%
+$(IA)^{3}$ 0.16M 54.65% 9.76% 46.79% -2.66%
+LLMs StarCoderBase-1B Prompt Tuning 0.05M 72.23% 46.36% 84.66% 71.21%
+Prefix Tuning 1.97M 75.65% 52.30% 85.44% 72.19%
+LoRA 3.59M 78.58% 60.68% 79.30% 58.85%
+$(IA)^{3}$ 0.31M 73.95% 49.32% 84.96% 71.71%
+StarCoderBase-3B Prompt Tuning 0.06M 68.96% 39.74% 83.63% 68.47%
+Prefix Tuning 4.06M 67.61% 35.78% 86.56% 74.06%
+LoRA 7.38M 70.85% 42.29% 83.70% 68.64%
+$(IA)^{3}$ 0.62M 69.35% 38.73% 84.90% 71.46%
+DeepSeek-Coder-1.3B Prompt Tuning 0.05M 78.98% 58.34% 84.17% 69.99%
+Prefix Tuning 1.97M 66.88% 37.49% 82.44% 64.90%
+LoRA 2.36M 78.98% 58.34% 82.72% 67.68%
+$(IA)^{3}$ 0.28M 83.87% 67.74% 85.84% 72.57%
+DeepSeek-Coder-6.7B Prompt Tuning 0.09M 78.89% 58.83% 85.56% 72.72%
+Prefix Tuning 5.25M 80.56% 61.81% 83.03% 68.14%
+LoRA 6.30M 77.32% 55.30% 83.43% 68.64%
+$(IA)^{3}$ 0.75M 77.32% 55.30% 85.34% 72.34%
+CodeLlama-7B Prompt Tuning 0.09M 72.52% 45.37% 85.49% 72.55%
+Prefix Tuning 5.25M 63.57% 30.67% 82.30% 64.61%
+LoRA 6.30M 77.32% 55.30% 83.10% 68.31%
+$(IA)^{3}$ 0.75M 74.19% 48.39% 83.54% 68.45%
 5.1.1. SLMs
@@
 For example, in CC detection, whether using full fine-tuning or PEFT methods, CodeBERT consistently achieved lower F1-scores and MCC values compared to GraphCodeBERT, with differences ranging from 0.11% to 5.93% and 0.21% to 12.10%, respectively.
-This is reasonable because GraphCodeBERT is built upon CodeBERT and further incorporates code structure information (i.e., data flow) during the pre-training phase (guo2021graphcodebert, ).
+This is reasonable because GraphCodeBERT is built upon CodeBERT and further incorporates code structure information (i.e., data flow) during the pre-training phase (guo2021graphcodebert).
 Across all SLMs, CodeT5 with PEFT methods significantly lags behind the other three models with PEFT methods in the code smell detection task, which requires both method- and class-level code understanding.
@@
 In contrast, for FE, LLMs fine-tuned with PEFT methods substantially outperform SLMs, likely due to the more complex and semantic nature of this smell type, which benefits from the richer contextual representations and broader knowledge captured by larger models.
-Figure 5. Peak GPU memory usage for detecting CC: (a) SLMs with PEFT methods and Full Fine-Tuning; (b) LLMs with PEFT methods
+Figure 6. Peak GPU memory usage for detecting $C$: (a) SLMs with PEFT methods and Full Fine-Tuning; (b) LLMs with PEFT methods
 Additionally, Figure 6 summarizes the peak GPU memory usage of both SLMs and LLMs with different PEFT methods on CC detection.
@@
 5.1.4. Method-level vs. Class-level Code Understanding
-The detection of CC and CM requires method-level code understanding, whereas identifying FE and DC necessitates class-level analysis (see Section 3.3.1).
+The detection of CC and CM requires method-level code understanding, whereas identifying FE and DC necessitates class-level analysis (see Section 3.1).
 By comparing the detection results of CC and CM with those of FE and DC, we observe that LMs demonstrate comparable effectiveness at both levels of code understanding.
@@
 5.1.5. Best PEFT Method
-Generally, for SLMs, both $(IA)^{3}$ and prefix tuning demonstrate strong performance across different models.
-For LLMs, $(IA)^{3}$ consistently delivers the best results on the detection of all four types of code smells.
-To be specific, the most effective PEFT method for SLMs on the task of code smell detection varies depending on the model: for instance, CodeBERT performs optimally with $(IA)^{3}$, whereas UnixCoder tends to achieve better results with prefix tuning.
-However, for LLMs, $(IA)^{3}$ always outperforms other PEFT methods across all studied code smells.
-Whether for StarCoderBase Models, Deepseek-Coder Models, or CodeLlama-7B, they all achieve superior performance when fine-tuned using $(IA)^{3}$.
-Additionally, for LLMs, prompt tuning can be employed as an alternative option.
+Generally, for SLMs, both $(IA)^{3}$ and prefix tuning achieve strong performance across different models, while for LLMs, $(IA)^{3}$ consistently delivers the best results across all four types of code smells.
+To be specific, the optimal PEFT method for SLMs on the task of code smell detection depends on the specific model: for instance, CodeBERT performs best with $(IA)^{3}$, whereas UnixCoder tends to favor prefix tuning.
+In contrast, $(IA)^{3}$ outperforms other PEFT techniques for all LLMs studied, including StarCoderBase Models, Deepseek-Coder Models, and CodeLlama-7B.
+Additionally, prompt tuning can serve as an alternative for LLMs, though its effectiveness is generally lower.
 Notably, despite achieving promising results on certain models and code smell types, LoRA exhibits instability in performance and can yield markedly poor effectiveness in some cases.
-An example of this is using LoRA on StarCoderBase-1B for CM detection.
-A case in point is the use of LoRA on StarCoderBase-1B for detecting CM.
-While StarCoderBase-1B achieves a minimum MCC of 54.01% with the other three PEFT methods (i.e., prompt tuning, prefix tuning, and $(IA)^{3}$), its performance drops sharply to 7.01% with LoRA, resulting in a disparity of over 40%.
+For example, on StarCoderBase-1B for CM detection, LoRA’s MCC drops to 7.01%, while the other three PEFT methods achieve a minimum MCC of 54.01%, resulting in a disparity of over 40%.
+As shown in Tables 6 and 7, the additional parameters injected by all PEFT methods are negligible relative to the total model size.
+Prefix tuning and LoRA generally introduce more trainable parameters than prompt tuning and $(IA)^{3}$, with LoRA adding the most and prompt tuning the least.
+For SLMs, the absolute differences in the number of added parameters remain modest due to their limited model size, whereas for LLMs the gaps become substantially larger.
+Overall, $(IA)^{3}$ introduces a relatively small amount of trainable parameters while delivering strong performance gains for both SLMs and LLMs in code smell detection.
+For SLMs, prefix tuning still remains a practical alternative; however, for LLMs it brings a considerably larger parameter overhead without providing comparable performance benefits.
 5.1.6. PEFT Methods vs. Full Fine-tuning
-According to Table 6 and Table 7, we can find that all the four PEFT methods achieve better or at least comparable effectiveness than full fine-tuning on most small models for code smell detection.
-For small LMs (excluding CodeT5), prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$ generally achieve varying degrees of improvement in F1-scores and MCC values compared to full fine-tuning for most code smell types.
+According to Tables 6 and 7, all four PEFT methods achieve better or at least comparable effectiveness than full fine-tuning on most small models for code smell detection.
+For SLMs (excluding CodeT5), prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$ generally achieve varying degrees of improvement in F1-scores and MCC values compared to full fine-tuning across most code smell types.
 However, for FE, all PEFT methods perform significantly worse than full fine-tuning.
-These results suggest that, in most cases, PEFT methods can match or surpass the performance of standard full fine-tuning for code smell detection while adjusting fewer number of trainable parameters.
-PEFT methods focus on enhancing the effectiveness of LMs to levels similar to full fine-tuning, while aiming to minimize the use of computational resources, such as peak GPU memory consumption.
-Taking CC detection as an example, Figure 5 (a) depicts the peak GPU memory consumption during the fine-tuning of the four SLMs.
-It can be observed that juxtaposed with full fine-tuning, all of the four PEFT methods lead to a substantial reduction in peak GPU memory usage, which further validates the efficiency of PEFT methods for code smell detection relative to full fine-tuning.
+The “#Params” columns in Tables 6 and 7 further show that all PEFT methods introduce a relatively small number of trainable parameters compared to full fine-tuning, typically ranging from tens to hundreds of millions.
+This demonstrates that, in most cases, PEFT approaches can match or even surpass the performance of standard full fine-tuning for code smell detection while updating far fewer parameters, highlighting their efficiency for the task of code smell detection.
+In addition to reducing trainable parameters, PEFT methods also lower computational overhead, particularly peak GPU memory usage, without sacrificing performance.
+Taking the detection of CC as an example, Figure 6(a) depicts the peak GPU memory consumption during the fine-tuning of the four SLMs.
+It can be observed that, juxtaposed with full fine-tuning, all PEFT methods lead to a substantial reduction in peak GPU memory usage, which further validates the efficiency of PEFT methods for code smell detection relative to full fine-tuning.
 Answer to RQ1: Overall, LLMs perform slightly better than SLMs in code smell detection, except for DC.
-Among different PEFT methods, both $(IA)^{3}$ and prefix tuning are most effective for SLMs, while $(IA)^{3}$ consistently yields the best results for LLMs.
-All PEFT techniques outperform or match the performance of full fine-tuning for several code smell types, while also reducing peak GPU memory usage, demonstrating their efficiency.
-However, for the detection of FE, full fine-tuning remains more effective.
+Among the PEFT methods, both $(IA)^{3}$ and prefix tuning are highly effective for SLMs, whereas $(IA)^{3}$ consistently delivers the best results for LLMs.
+All PEFT techniques either exceed or match full fine-tuning while requiring far fewer trainable parameters and significantly reducing peak GPU memory usage, underscoring their efficiency for code smell detection.
 5.2. RQ2: What is the effectiveness of PEFT methods with different hyper-parameters?
-4.2.1. Experiment Setup
 Given the restrictions of our computational resources and popularity of LoRA, we focused on a single PEFT method, LoRA, to investigate how SLMs and LLMs perform under different hyper-parameter settings.
@@
 The experiment results are presented in Figure 7.
-Figure 6. MCC results of GraphCodeBERT and DeepSeek-Coder-1.3B using LoRA for code smell detection under different rank settings: (a) CC; (b) CM; (c) FE; (d) DC
+Figure 7. MCC results of GraphCodeBERT and DeepSeek-Coder-1.3B using LoRA for code smell detection under different rank settings: (a) $C$; (b) CM; (c) FE; (d) DC
 5.2.1. Results of RQ2
@@
 5.3. RQ3: What is the effectiveness of PEFT methods in low-resource scenarios?
-4.3.1. Experiment Setup
-To simulate low-resource scenarios, following previous studies (wang2023one, ; liu2023empirical, ), we randomly selected 50, 100, 250, and 500 samples for CC, CM, and DC, respectively, based on the size of our dataset.
+To simulate low-resource scenarios, following previous studies (wang2023one; liu2023empirical), we randomly selected 50, 100, 250, and 500 samples for CC, CM, and DC, respectively, based on the size of our dataset.
 Note that each subset is balanced, containing an equal number of positive and negative samples.
-We excluded FE from this RQ, as the total number of available FE samples in our training set is already below 500 (see Section 3.3.6), making it inherently a low-resource task.
+We excluded FE from this RQ, as the total number of available FE samples in our training set is already below 500 (see Section 3.6), making it inherently a low-resource task.
 The experiment results are shown in Table 8.
@@
 Results of GrahphCodeBERT and DeepSeek-Coder-1.3B with different PEFT methods and full fine-tuning on detecting CC, CM, and DC in low-resource scenarios (bold: highest values per scenario for a model, underline: highest values among all scenarios for a model)
-Model #Training Samples Method CC CM DC
+Model #Training Samples Method $C$ CM DC
 GraphCodeBERT 50 Full FT 19.23% 54.61% 62.76%
@@
 Moreover, when the number of training samples reaches 250 or 500, several PEFT techniques are able to outperform full fine-tuning, demonstrating their potential even in moderately low-resource settings.
+5.4. RQ4: What is the effectiveness of PEFT methods compared to state-of-the-art code smell detection approaches?
 Table 9. Prompt 1 for GPT-4o-mini and DeepSeek-v3 in detecting CC, CM, FE, and DC
 Code Smell Prompt Template*
-CC You are a professional Java programmer. Given a piece of Java code, determine whether it contains “Complex Conditional” code smell. If it does, return only “Yes”, else return “No”.
+$C$ You are a professional Java programmer. Given a piece of Java code, determine whether it contains “Complex Conditional” code smell. If it does, return only “Yes”, else return “No”.
 Definition of “Complex Conditional”: This code smell occurs when a conditional statement is excessively intricate.
@@
 ### Output: [Yes/No]
-  • * Prompt 1 shown in this table was curated by the authors. Prompt 2 and Prompt 3, which were generated by ChatGPT, are available in our replication package (replicationPackage, ).
-4.4. RQ4: What is the effectiveness of PEFT methods compared to state-of-the-art code smell detection approaches?
-4.4.1. Experiment Setup
+  • * Prompt 1 shown in this table was curated by the authors. Prompt 2 and Prompt 3, which were generated by ChatGPT, are available in our replication package (replicationPackage).
 Figure 8. Results of DeepSeek-v3 and GPT-4o-mini on detecting CC, CM, FE, and DC under ICL
-In this RQ, we established both heuristic-based (i.e., DesigniteJava and PMD) and best-in-class DL detectors (i.e., DeepSmells (ho2023fusion, ) and AE-Dense (sharma2021code, )) as baselines for comparison.
-Additionally, In-Context Learning (ICL) (tom2020language, ) has demonstrated impressive performance in various SE tasks, such as code generation (gao2024what, ; jia2024acecoder, ).
-To evaluate its potential in code smell detection, we explored the effectiveness of two recent and powerful general-purpose LLMs, GPT-4o-mini (hurst2024gpt4o, ) and DeepSeek-v3 (liu2024deepseek, ), under both zero-shot and few-shot settings.
-We selected these two models because they represent the latest advancements in commercial and open-weight general-purpose LLMs, respectively, and have demonstrated strong performance across a range of natural language and programming-related tasks (rasheed2024task, ; bruni2025benchmarking, ; liu2024deepseek, ; jahin2025unveiling, ).
-Specifically, we first designed a prompt template (i.e., prompt 1) for each of the four types of code smells examined in this study, incorporating both instructional text and smell definitions, as shown in Table 9.
-We then used ChatGPT (ChatGPT, ) to generate two additional variants for each template (i.e., prompt 2 and prompt 3), resulting in three distinct prompts per code smell type.
-For the zero-shot setting, these prompts were directly provided to GPT-4o-mini and DeepSeek-v3 without any labeled examples.
-In the few-shot setting, example pairs were randomly sampled from the training set, where each pair consisted of one code sample with a code smell and one without, along with their corresponding labels.
-These examples were combined with the base prompts to form few-shot configurations containing one, two, or three example pairs.
-Due to resource constraints (e.g., limited API usage budget), we restricted the few-shot setting to include no more than three example pairs per prompt.
-As a result, each code smell type yields a total of 12 unique prompt-example configurations, derived from three prompt variants combined with four levels of contextual information (zero-shot, plus one-, two-, and three-pair few-shot settings).
-This ICL setup ensures diversity in both prompt formulations and example composition, allowing for a more comprehensive evaluation of general-purpose LLMs and enabling a fair comparison with our proposed PEFT-based methods in the context of code smell detection.
-Figure 7 presents the performance of GPT-4o-mini and DeepSeek-v3 on code smell detection in both zero- and few-shot settings.
-Table 10 summarizes the final results for RQ4, where we report only the best-performing configurations for both GPT-4o-mini and DeepSeek-v3 under ICL across all prompts and example variants.
-4.4.2. ICL
-It can be observed from Figure 7 that DeepSeek-v3 generally demonstrates better effectiveness in code smell detection compared to GPT-4o-mini.
-Additionally, both models achieve better performance under few-shot settings than in the zero-shot setting.
-However, their performance does not consistently improve as the number of example pairs increases.
-4.4.3. Comparison with State-of-the-Art Approaches
+Figure 9. Results of DeepSeek-v3 and GPT-4o-mini on detecting CC, CM, FE, and DC with and without labeling criteria
+To answer this RQ, we benchmark the performance of our PEFT-based methods against the three categories of baselines established in Section 4.4: heuristics-based detectors (DesigniteJava, PMD), DL-based approaches (DeepSmells, AE-Dense), and general-purpose LLMs (GPT-4o-mini, DeepSeek-v3) with In-Context Learning (ICL).
+For the general-purpose LLMs, we evaluated their effectiveness under the ICL paradigm.
+To ensure the robustness of our evaluation against prompt sensitivity, we designed three prompt formulations for each code smell type: a base prompt containing instructional text and definitions (Prompt 1, as shown in Table 9), and two additional variants (Prompts 2 and 3) generated using ChatGPT (ChatGPT) to introduce linguistic diversity.
+These prompts were applied in three ICL settings:
+  1. (1)
+    Zero-shot setting: prompts used directly without examples;
+  2. (2)
+    Few-shot setting: prompts augmented with one, two, or three randomly sampled example pairs (consisting of one smelly and one clean code sample) to provide contextual guidance;
+  3. (3)
+    Criteria-informed setting: prompts enriched with the explicit labeling rules derived from our manual review process (Section 3.5).
+This multi-faceted design ensures diversity in both prompt formulation and contextual information, thereby supporting a comprehensive evaluation of general-purpose LLMs and enabling a fair comparison with our PEFT-based methods in the context of code smell detection.
+In total, this yielded 15 unique configurations per smell type (combinations of three prompt variants and five varying contextual levels).
+Figure 8 visualizes the performance trends of GPT-4o-mini and DeepSeek-v3 across the zero- and few-shot settings, while Figure 9 illustrates the impact of incorporating labeling criteria.
+Table 10 summarizes the final comparative results, reporting the best-performing configuration of each LLM alongside our PEFT methods and other baselines.
+5.4.1. Performance of General-Purpose LLMs with ICL
+It can be observed from Figures 8 and 9 that DeepSeek-v3 generally demonstrates better effectiveness in code smell detection compared to GPT-4o-mini across all ICL settings.
+Both models perform better under few-shot settings than in the zero-shot setting; however, their performance does not consistently improve as the number of example pairs increases.
+Moreover, GPT-4o-mini shows a noticeable performance gain under the criteria-informed setting, whereas DeepSeek-v3 performs better when labeling criteria are not provided.
+This difference may stem from model characteristics: GPT-4o-mini, a smaller commercial model with stronger instruction alignment, benefits more from explicit and structured guidance such as labeling criteria, which compensates for its limited parameter capacity and contextual reasoning ability.
+In contrast, DeepSeek-v3, with its larger scale and broader training on code-related corpora, can generalize more effectively from implicit semantics and example patterns, exhibiting stronger reasoning ability even without explicit rules.
+5.4.2. Overall Comparison with State-of-the-Art Approaches
 Table 10. Results of PEFT methods compared to state-of-the-art code smell detection approaches on detecting CC, CM, FE, and DC (bold: highest values per code smell)
 Code Smell Method Precision Recall F1 MCC
-CC DesigniteJava 64.01% 76.96% 69.89% 34.41%
+$C$ DesigniteJava 64.01% 76.96% 69.89% 34.41%
 DeepSmells 51.06% 100.00% 67.61% 6.63%
 AE-Dense 51.13% 49.10% 50.10% 2.92%
-GPT-4o-mini 71.48% 64.46% 61.29% 35.24%
+GPT-4o-mini 72.13% 70.11% 69.41% 42.19%
 DeepSeek-v3 72.38% 71.74% 71.53% 44.12%
@@
 GPT-4o-mini 77.53% 77.42% 77.40% 54.95%
-DeepSeek-v3 74.29% 74.19% 74.17% 48.49%
+DeepSeek-v3 77.00% 74.19% 73.50% 51.12%
 DeepSeek-Coder-1.3B with $(IA)^{3}$ 83.87% 83.87% 83.87% 67.74%
@@
 GPT-4o-mini 84.87% 82.51% 82.21% 67.35%
-DeepSeek-v3 87.57% 87.57% 87.57% 75.15%
+DeepSeek-v3 89.48% 89.45% 89.45% 78.93%
 GraphCodeBERT with prefix tuning 90.95% 90.03% 89.97% 79.26%
 Based on the results presented in Table 10, we can find that LMs fine-tuned with PEFT techniques consistently demonstrate the highest effectiveness among all existing methods for code smell detection.
-Additionally, among conventional baseline approaches, DeepSeek-v3 achieves the most competitive overall performance.
-While it shows a significant performance gap compared to LMs tuned with PEFT methods for CM and FE detection (with MCC differences exceeding 10%), its performance in CC and DC detection remains relatively strong, with only a 4%-5% reduction in MCC values, demonstrating its relative strength in these specific code smell types.
+Specifically, compared to the best-performing baseline, PEFT-tuned LMs achieve MCC improvements of 5.85%, 13.69%, 12.79%, and 0.33% for CC, CM, FE, and DC, respectively.
+Among various baseline approaches, DeepSeek-v3 achieves the most competitive overall performance.
+However, despite its relative strength in detecting CC and DC, a significant performance gap remains: DeepSeek-v3 lags considerably behind PEFT-tuned LMs in CM and FE detection, where the performance differences exceed 10%.
 Furthermore, our results indicate that traditional DL-based methods (DeepSmells and AE-Dense) exhibit substantially inferior performance compared to both PEFT and the other baseline approaches.
-This finding underscores the limitations of conventional DL models in code smell detection, particularly when compared to more advanced techniques like PEFT techniques and ICL.
+This finding underscores the limitations of conventional DL models in code smell detection, particularly when compared to more advanced techniques like PEFT and ICL.
 Answer to RQ4: In comparison with existing code smell detection approaches, LMs tuned with PEFT methods reveal notable performance improvements, highlighting the effectiveness of PEFT methods for code smell detection.
 6. Discussion
-Approaches with improved effectiveness for CC code smell detection are needed: According to Fowler (fowler1999refactoring, ), four factors contribute to making programs difficult to work with, one of which is Complex Conditional (CC) logic that complicates modification.
-The presence of CC in source code hinders testing and maintenance (fowler1999refactoring, ; liu2014automated, ; alhefdhi2022framework, ).
-Thus, identifying CC code smells in software can significantly improve code quality (liu2014automated, ).
+Approaches with improved effectiveness for CC code smell detection are needed: According to Fowler (fowler1999refactoring), four factors contribute to making programs difficult to work with, one of which is Complex Conditional (CC) logic that complicates modification.
+The presence of CC in source code hinders testing and maintenance (fowler1999refactoring; liu2014automated; alhefdhi2022framework).
+Thus, identifying CC code smells in software can significantly improve code quality (liu2014automated).
 However, the effectiveness of existing methods for detecting this type of code smell is considerably limited.
-As shown in Table 10, existing methods, including heuristic-based (e.g., DesigniteJava), DL-based (e.g., DeepSmells) and AE-Dense), and general-purpose LLMs under ICL settings (e.g., GPT-4o-mini and DeepSeek-v3), all demonstrate limited success in detecting CC, with all MCC values below 50%, and some even as low as 6.63%.
+As shown in Table 10, existing methods, including heuristics-based (e.g., DesigniteJava), DL-based (e.g., DeepSmells) and AE-Dense), and general-purpose LLMs under ICL settings (e.g., GPT-4o-mini and DeepSeek-v3), all demonstrate limited success in detecting CC, with all MCC values below 50%, and some even as low as 6.63%.
 Although our proposed method, StarCoderBase-3B with $(IA)^{3}$, improved the performance of CC code smell detection significantly, reaching an F1-score of 74.68% and a MCC value of 49.96%, it still falls short of expectations due to challenges such as the considerable existence of false positives and negatives, which may limit its practical applicability.
@@
 PEFT methods can enhance the capacity of LMs in class-level code understanding: In this study, we selected four representative types of code smells, i.e., CC, CM, FE, and DC, to evaluate the effectiveness of PEFT methods.
-While only DC is formally categorized as a class-level smell, FE detection also inherently requires class-level analysis as well, as it involves analyzing method-class dependencies (see Section 3.3.1 and Section 3.3.4).
-According to the recent studies (silva2024detecting, ; mesbah2025leveraging, ), LLMs not specifically fine-tuned for code smell detection generally perform worse when tasked with identifying smells that demand class-level understanding.
+While only DC is formally categorized as a class-level smell, FE detection also inherently requires class-level analysis as well, as it involves analyzing method-class dependencies (see Section 3.1 and Section 3.4).
+According to the recent studies (silva2024detecting; mesbah2025leveraging), LLMs not specifically fine-tuned for code smell detection generally perform worse when tasked with identifying smells that demand class-level understanding.
 This aligns with our intuitive expectation that class-level code understanding is more difficult, as it requires capturing higher-level structural relationships (e.g., field-method interactions, class cohesion, and design-level abstractions).
-Despite this challenge, most existing research on PEFT methods has predominantly concentrated on the their perfomance for method-level code understanding (liu2023empirical, ; wang2023one, ; liu2024delving, ), leaving their effectiveness on class-level tasks largely unexplored.
+Despite this challenge, most existing research on PEFT methods has predominantly concentrated on the their perfomance for method-level code understanding (liu2023empirical; wang2023one; liu2024delving), leaving their effectiveness on class-level tasks largely unexplored.
 Our work bridges this critical gap by evaluating PEFT methods on code smell detection, a task spans both method- and class-level analysis.
@@
 These findings suggest that PEFT methods not only improve parameter efficiency but also strengthen the proficiency of LLMs to perform robust code understanding across different levels of granularity.
+When fine-tuning LMs for code smell detection, more trainable parameters do not guarantee better results: Our findings indicate that the effectiveness of fine-tuning is not primarily determined by the number of parameters being updated.
+Although full fine-tuning adjusts all model parameters, it does not consistently outperform PEFT methods in code smell detection.
+For many SLMs and most code smell types, PEFT approaches achieve comparable or even superior performance while adding only a small fraction of parameters.
+This indicates that large-scale parameter updates are not necessary for effective model adaptation in this task.
+A similar pattern emerges when comparing individual PEFT techniques.
+LoRA and prefix tuning introduce far more trainable parameters than prompt tuning or $(IA)^{3}$, yet their additional capacity does not reliably translate into better performance.
+In some cases, LoRA even performs markedly worse despite being the most parameter-intensive PEFT method.
+In contrast, $(IA)^{3}$ frequently provides the best overall performance while relying on relatively few added parameters.
+Taken together, these observations confirm that increasing the number of trainable parameters alone does not guarantee better outcomes.
+Instead, the effectiveness of fine-tuning depends on how efficiently the model’s internal representations are adapted and on the strategic placement of parameter updates for code smell detection.
 While larger LMs with PEFT methods often achieve better effectiveness in code smell detection, SLMs still retain distinct advantages: The results of RQ1 reveal that LLMs generally outperform SLMs in code smell detection, particularly in identifying FE.
-This performance advantage aligns with established neural scaling laws (kaplan2020scaling, ), which posit that model capabilities improve predictably with increased scale.
+This performance advantage aligns with established neural scaling laws (kaplan2020scaling), which posit that model capabilities improve predictably with increased scale.
 However, SLMs are able to achieve comparable performance in detecting CC and CM, and even significantly outperform LLMs in the detection of DC.
@@
 Consequently, beginning fine-tuning experiments with a smaller-size dataset can be an effective strategy to optimize resource usage.
-This approach allows researchers and practitioners to initially evaluate a wide range of LMs in a shorter time and wither fewer computational resources, saving substantial effort.
+This approach allows researchers and practitioners to initially evaluate a wide range of LMs in a shorter time and with fewer computational resources, saving substantial effort.
 It enables them to identify the most suitable model for the specific task through an initial evaluation before committing to the more resource-intensive process of training on larger datasets.
@@
 By leveraging smaller datasets for initial evaluation, one can make informed decisions about which models are likely to perform well with limited resources, and further fine-tune or scale up as necessary, without wasting valuable time or computational power.
-ICL shows promise as an effective alternative to PEFT methods for certain types of code smells: In RQ4, we investigate the performance of general-purpose LLMs in code smell detection under In-Context Learning (ICL) settings.
-As shown in Table 10, for CC and DC, the optimal ICL configuration of DeepSeek-v3 achieves MCC scores only 4%–5% lower than those attained by the top-performing PEFT-tuned LMs.
-This suggests that, for certain types of code smells, well-crafted prompts and a few in-context examples may enable general-purpose LLMs to produce competitive results.
-Moreover, applying ICL technique requires no model training or parameter updates (zhou2024large, ; tom2020language, ; li2025large, ), which is particularly appealing for resource-constrained scenarios or when model weights cannot be modified (e.g., in commercial LLM APIs).
-However, for CM and FE, ICL lags significantly behind, with performance deficits of up to 13.69% and 29.25% in MCC, respectively, compared to our PEFT-based approaches.
+ICL shows promise as an effective alternative to PEFT methods for certain types of code smells: In RQ4, we investigated the performance of general-purpose LLMs in code smell detection under In-Context Learning (ICL) settings, including zero-shot, few-shot, and criteria-informed setups.
+As shown in Table 10, for CC and DC, the optimal ICL configuration of DeepSeek-v3 achieves MCC scores only 1%–5% lower than those attained by the top-performing PEFT-tuned LMs.
+This suggests that, for certain types of code smells, well-crafted prompts with ICL may enable general-purpose LLMs to produce competitive results.
+Moreover, applying ICL technique requires no model training or parameter updates (zhou2024large; tom2020language; li2025large), which is particularly appealing for resource-constrained scenarios or when model weights cannot be modified (e.g., in commercial LLM APIs).
+However, for CM and FE, ICL lags significantly behind, with performance deficits of up to 13.69% and 16.62% in MCC, respectively, compared to our PEFT-based approaches.
 These results indicate that detecting these smell types likely requires deeper model adaptation to capture subtle semantic patterns and structural nuances that few-shot prompting alone cannot convey.
@@
 This observation suggests that simply increasing the number of in-context examples without considering their quality, relevance, or diversity may not be sufficient to enhance model effectiveness for the task of code smell detection.
-By contrast, Retrieval-Augmented Generation (RAG) (lewis2020retrieval, ) can dynamically retrieve the most relevant examples from an external knowledge base to augment the prompts.
+By contrast, Retrieval-Augmented Generation (RAG) (lewis2020retrieval) can dynamically retrieve the most relevant examples from an external knowledge base to augment the prompts.
 By grounding predictions in high-quality and contextually appropriate evidence, RAG has the potential to enhance detection accuracy while avoiding the need for fine-tuning or complex prompt engineering.
@@
 Construct Validity: One primary threat to the construct validity of this study is the construction of the datasets.
-We first used DesigniteJava (designitejava, ) to detect potential positive and negative samples, then manually reviewed all the candidate data to validate the datasets, ensuring there were no false positives or negatives.
+We first used DesigniteJava (designitejava) to detect potential positive and negative samples, then manually reviewed all the candidate data to validate the datasets, ensuring there were no false positives or negatives.
 First, relying solely on DesigniteJava and PMD for initial sample detection may introduce bias.
-However, DesigniteJava and PMD are well-established and widely adopted tools in Java code smell detection (chaniotaki2021architecture, ; sharma2021qscored, ; shah2023mining, ; tushar2016does, ; taibi2017how, ), providing a robust foundation for the identification of potential code smells.
+However, DesigniteJava and PMD are well-established and widely adopted tools in Java code smell detection (chaniotaki2021architecture; sharma2021qscored; shah2023mining; tushar2016does; taibi2017how), providing a robust foundation for the identification of potential code smells.
 We also acknowledge that future studies could benefit from incorporating additional tools in the initial code smell detection process, which may help mitigate the threat to construct validity by providing more diverse perspectives on the data.
@@
 Future work could expand on this by increasing the number of example pairs and evaluating additional models to better understand the full potential of ICL performance.
-However, prior studies such as UniLog (xu2024unilog, ) have shown that the performance of LLMs on software engineering tasks does not always improve linearly with more examples, which suggests that while increasing the number of examples may be beneficial, performance gains are not guaranteed.
+However, prior studies such as UniLog (xu2024unilog) have shown that the performance of LLMs on software engineering tasks does not always improve linearly with more examples, which suggests that while increasing the number of examples may be beneficial, performance gains are not guaranteed.
 As such, we believe our current evaluation still provides valuable insights and that this threat to validity is partially mitigated.
@@
 However, due to limited computational resources, we did not include larger models with over 7B parameters, which may impact the generalizability of our findings.
-Furthermore, we only employed four widely used and typical PEFT techniques, i.e., prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$.
-While these methods cover the major categories defined in the PEFT library, the exclusion of other approaches may limit the generalizability of our findings.
-For example, methods such as adapter tuning (houlsby2019parameter, ) could be explored in future work to provide a more comprehensive evaluation.
+Furthermore, our study focused exclusively on four widely used and representative PEFT methods (i.e., prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$) implemented within the Hugging Face PEFT library.
+Techniques such as adapter tuning (houlsby2019parameter) were excluded because they are not yet well supported by the library and would require additional external implementations, which could compromise the consistency of our experimental setup.
+As a result, our findings may not fully generalize to PEFT methods beyond the four examined ones.
+To mitigate this limitation, we carefully selected diverse PEFT methods covering the major families (prompt-based, LoRA-based, and $(IA)^{3}$) to ensure representativeness across categories.
+In future work, we plan to incorporate adapter tuning and other emerging PEFT techniques as they become more maturely integrated into the Hugging Face ecosystem, thereby enabling a broader and more comprehensive comparison.
 Moreover, our study applied PEFT methods exclusively to the detection of Java code smells, focusing on four specific types: three method-level smells (Complex Conditional, Complex Method, and Feature Envy) and one class-level smell (Data Class).
@@
 To mitigate this threat, we referred to previous works and conducted pilot experiments to determine the appropriate hyper-parameter values.
-Another potential threat to internal validity stems from conducting only a single experimental run per configuration due to computational resource constraints.
-Although we ensured consistency across all experiments by using a fixed random seed (42), this approach cannot fully account for the inherent stochasticity in the training process.
-To assess the impact of randomness, we conducted supplementary experiments using two alternative seeds (0 and 999) for a representative case (GraphCodeBERT on CC detection).
-The results shown in Table 11 indicate that there are only small variations in the performance metrics across different random seeds.
-These minor differences suggest that our core findings are robust to random initialization effects, with the overall trends remaining consistent.
-While some performance fluctuations are observed, they do not significantly alter the conclusions drawn from the experiments.
-However, we acknowledge that conducting multiple runs with different seeds would yield more statistically reliable conclusions.
+In our experiments, we used a batch size of 1 for all models due to computational resource constraints, which ensured consistency but could potentially affect convergence and performance.
+To evaluate this, we conducted supplementary experiments with different batch sizes, reported in Appendix A.1, which show that batch size has a limited impact on overall detection performance across PEFT methods, suggesting the main conclusions remain robust.
+Similarly, each configuration was run only once using a fixed random seed (42), which does not fully capture the stochasticity inherent in training.
+Additional runs with two alternative seeds (0 and 999) for a representative case (GraphCodeBERT on CC detection), reported in Appendix A.2, indicate minor performance variations, confirming robustness to random initialization.
+Nevertheless, we acknowledge that more comprehensive experiments with larger batch sizes and multiple seeds across more models could provide further insights and increase statistical reliability.
+CodeBERT, GraphCodeBERT, and CodeT5 have a maximum input length of 512 tokens, requiring longer code samples to be truncated.
+This could potentially omit semantically important content relevant for code smell detection.
+We analyzed token length distributions and manually reviewed a subset of long samples (50 positives and 50 negatives per smell type) in Appendix B.
+The review showed that essential indicators of each code smell typically occur within the first 512 tokens, and truncated segments rarely contain critical information.
+Thus, while context window limitations represent a potential threat, their impact on overall evaluation results is likely minimal.
+Additionally, data leakage represents another potential threat to internal validity.
+First, the poor zero-shot performance of GPT-4o-mini in detecting CC and CM code smells suggests that the model was not exposed to the test data (see Section 5.4), as we would expect stronger zero-shot effectiveness if GPT-4o-mini had memorized the data.
+Second, the significant improvements in performance for both small and large LMs following the application of PEFT techniques further suggest that memorization is unlikely to be an issue.
+Such substantial performance gains would not occur if the models had already memorized the test sets.
+Furthermore, Cao et al.
+(cao2024concerned) highlight that code LMs do not consistently exhibit reduced performance on data extending beyond their training cut-off dates.
+This careful management of data, along with the observed performance improvements after fine-tuning, ensures that data leakage was not a major threat, thereby maintaining the internal validity of our findings.
+8. Related Work
+8.1. Code Smell Detection
+Existing approaches to code smell detection predominantly rely on heuristics-based methods, which utilize predefined rules to identify problematic code patterns (pmd; designitejava; campbell2013sonarqube; fokaefs2007jdeodorant; vidal2015jspirit; moha2010decor).
+These methods aim to flag potential smells by analyzing the source code against established standards.
+For instance, PMD (pmd), a widely used static code analysis tool, supports 16 programming languages and applies rule sets to detect code smells.
+These rule sets can be tailored to suit specific project requirements or design standards, making the tool adaptable for different contexts.
+Similarly, Sharma (designitejava) introduced DesigniteJava, an advanced tool capable of identifying code smells at multiple levels — ranging from architecture and design to implementation.
+Traditional heuristics-based code smell detectors face inherent challenges, such as difficulty in defining effective thresholds for detection.
+In response, Machine Learning (ML) and Deep Learning (DL) techniques have gained traction as promising alternatives.
+Maiga et al.
+(maiga2012support) implemented SVMDetect based on an ML technique — support vector machines.
+Fontana et al.
+(fontana2016comparing) evaluated 16 ML algorithms across four types of code smells using 74 software systems, reporting that J48 and Random Forest delivered the highest performance.
+Liu et al.
+(liu2021deep) utilized deep neural networks and advanced DL techniques to perform useful feature selection from source code and map these features to labels (smelly or not).
+Sharma et al.
+(sharma2021code) trained Convolution Neural Network (CNN), Recurrent Neural Network (RNN), and autoencoder DL models in different configurations to detect code smells.
+Their work confirms the feasibility of applying transfer-learning for code smell detection.
+Ho et al.
+(ho2023fusion) developed DeepSmells based on a CNN and long short-term memory (LSTM) networks for detecting code smells.
+They conducted an empirical study to compare DeepSmells with other well-established detection tools.
+Their results demonstrate that DeepSmells outperforms the baselines, achieving superior performance.
+They also anticipated that leveraging CodeBERT would enhance the effectiveness of code smell detection.
+We conduct an extensive analysis of the Machine Learning Code Quality (MLCQ) dataset, focusing on how these LLMs perform when prompted to identify and classify code smells.
+Recently, researchers have started to explore the use of LLMs for code smell detection.
+Mesbah et al.
+(mesbah2025leveraging) used the MLCQ dataset to analyze how GPT-4 and LLaMA perform when prompted to identify and classify code smells.
+Silva et al.
+(silva2024detecting) employed two types of prompts to query ChatGPT for identifying Blob, Data Class, Feature Envy, and Long Method code smells.
+Sadik et al.
+(sadik2025benchmarking) created a dataset for evaluating LLMs on automated code smell detection.
+They benchmarked two state-of-the-art LLMs and found that LLMs are not yet ready to replace deterministic tools in all aspects of code quality assurance.
+Wu et al.
+(wu2024ismell) trained a Mixture of Experts (MoE) model, leveraging specialized expert toolsets for code smell detection.
+They subsequently utilized the detection results to prompt LLMs for refactoring.
+Most prevailing ML-based or DL-based approaches for code smell detection train models from scratch.
+On the other hand, existing LLM-based approaches typically rely on simple prompting strategies without any form of fine-tuning or adaptation.
+Compared to their works, we harness the power of LLMs using PEFT methods to detect four types of widely studied code smells (Complex Conditional, Complex Method, Feature Envy, and Data Class), aiming to enhance the models’ ability to capture the nuanced relationships between different code smell types.
+8.2. PEFT Methods on Software Engineering Tasks
+Several studies have evaluated PEFT methods on various Software Engineering (SE) tasks.
+Wang et al.
+(wang2023one) fine-tuned UniXcoder and CodeT5 by inserting the parameter-efficient structure adapter for code search and summarization.
+According to their results, adapter tuning has shown effectiveness in cross-lingual and low-resource scenarios, surpassing full fine-tuning and effectively mitigating catastrophic forgetting.
+Liu et al.
+(liu2023empirical) carried out experiments by fine-tuning code models with adapter, prefix tuning, LoRA, and MHM (he2022towards) on both the code understanding task and code generation task.
+Their findings indicate that adapter tuning and LoRA can achieve promising performance on these two SE tasks with reduced training costs and memory requirements.
+Liu et al.
+(liu2024delving) examined whether adapter tuning and LoRA can perform well on code-change-related tasks, specially just-in-time defect prediction and commit message generation.
+Their study demonstrates that PEFT methods can serve as a powerful alternative to full fine-tuning for dynamic code comprehension.
+Weyssow et al.
+(weyssow2023exploring) compared PEFT methods, including prompt tuning and others, with In-Context Learning (ICL) in terms of code generation.
+Their investigation reveals the superiority of PEFT over ICL and highlights the potential opportunities of tuning LLMs in diverse SE scenarios.
+Li et al.
+(li2024comprehensive) created an instruction dataset of Automated Program Repair (APR) using prompt engineering.
+They employed this dataset to fine-tune four LLMs using four different PEFT methods.
+The results show that $(IA)^{3}$ achieves the highest fixing ability compared to prompt tuning, prefix tuning, and LoRA.
+These studies focus on evaluating PEFT methods across a range of SE tasks, primarily at the method level.
+Compared to their works, our work specifically targets the task of code smell detection, which, to the best of our knowledge, has not been studied before.
+Additionally, our work further explores the capacity of LLMs on class-level code understanding, thereby addressing a significant gap in existing research.
+9. Conclusions and Future Work
+In this work, we investigated the effectiveness of various PEFT methods on different LMs for code smell detection.
+We also evaluated how the performance of these PEFT methods varies under different conditions, such as varying hyper-parameter settings and the availability of training data.
+Our analysis reveals that PEFT methods are more efficient than full fine-tuning and In-Context Learning (ICL) in most cases, underscoring the potential of PEFT methods to enhance code smell detection and offer valuable insights for optimizing LMs for other Software Engineering (SE) tasks.
+Additionally, we have constructed high-quality datasets for four widely studied code smells, i.e., Complex Conditional, Complex Method, Feature Envy, and Data Class, and we provided the replication package of this work online (replicationPackage) to facilitate future research and ensure the reproducibility of our study results.
+In the next step, we plan to further explore how PEFT performs in cross-language scenarios for code smell detection.
+Moreover, we plan to investigate the application of PEFT methods for fine-tuning LMs to assist in a wider range of SE tasks, such as bug detection, code completion, and code smell refactoring.
+These directions will help us better understand the generalizability and practical value of PEFT methods in real-world software development settings.
+Data Availability
+The replication package for this work has been made available at https://github.com/MabelQi/PEFT4CSD.
+Acknowledgements.
+This work has been partially supported by the National Natural Science Foundation of China (NSFC) with Grant No.
+62172311.
+The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of Wuhan University.
+Appendix
+Appendix A The Impact of Other Hyper-Parameters on PEFT Methods
+In addition to the primary experimental settings discussed in the main text, we investigate the effects of two additional hyper-parameters on PEFT methods: batch size and random seed selection.
+Understanding the sensitivity of PEFT methods to these factors is crucial for assessing the stability and reproducibility of our results.
+Specifically, we examine how varying batch sizes during training influences model efficiency and prediction accuracy.
+We also evaluate performance variations across multiple random seeds to ensure that our findings are not driven by a particular initialization.
+A.1. Batch Size
+Table 11. Results of GraphCodeBERT with PEFT methods and full fine-tuning on detecting CC across different batch sizes
+Batch Size Method Precision Recall F1 MCC
+1 Full FT 74.06% 74.02% 74.01% 48.08%
+Prompt Tuning 65.49% 65.33% 65.24% 30.81%
+Prefix Tuning 70.08% 69.89% 69.82% 39.97%
+LoRA 73.15% 73.15% 73.15% 46.30%
+$(IA)^{3}$ 72.50% 72.50% 72.50% 45.00%
+8 Full FT 72.10% 72.07% 72.06% 44.16%
+Prompt Tuning 63.20% 63.15% 63.12% 26.36%
+Prefix Tuning 71.24% 70.98% 70.89% 42.22%
+LoRA 73.76% 73.59% 73.54% 47.34%
+$(IA)^{3}$ 72.01% 71.85% 71.80% 43.85%
+16 Full FT 73.19% 72.83% 72.72% 46.02%
+Prompt Tuning 63.48% 63.48% 63.47% 26.96%
+Prefix Tuning 69.66% 69.02% 68.77% 38.68%
+LoRA 73.59% 73.59% 73.59% 47.18%
+$(IA)^{3}$ 71.65% 71.30% 71.19% 42.95%
+32 Full FT 73.08% 72.61% 72.47% 45.69%
+Prompt Tuning 63.27% 63.26% 63.25% 26.53%
+Prefix Tuning 69.32% 68.48% 68.13% 37.79%
+LoRA 73.46% 73.15% 73.06% 46.61%
+$(IA)^{3}$ 70.66% 69.35% 68.85% 39.99%
+Table 11 presents the performance of GraphCodeBERT with different PEFT methods (prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$), as well as full fine-tuning, for detecting the Complex Conditional (CC) code smell across various batch sizes (1, 8, 16, and 32).
+From Table 11, we observe that batch size has a limited impact on overall detection performance.
+Full fine-tuning and LoRA consistently achieve the highest MCC values across all batch sizes, while $(IA)^{3}$ also maintains strong performance.
+Prompt tuning generally yields lower scores, and prefix tuning falls in between.
+Minor fluctuations are present; for example, LoRA attains slightly higher F1 and MCC scores at batch size 8 compared to batch size 1.
+Overall, the performance trends remain stable, suggesting that PEFT methods are relatively robust to batch size variations and that the main experimental conclusions are unlikely to be materially affected.
 Table 12. Results of GraphCodeBERT with PEFT methods and full fine-tuning on detecting CC across different seeds
@@
 $(IA)^{3}$ 72.23% 72.17% 72.15% 44.41%
-Additionally, data leakage represents another potential threat to internal validity.
-First, the poor zero-shot performance of GPT-3.5-turbo in detecting CC and CM code smells suggests that the model was not exposed to the test data (see Section 4.4), as we would expect stronger zero-shot effectiveness if GPT-3.5-turbo had memorized the data.
-Second, the significant improvements in performance for both small and large LMs following the application of PEFT techniques further suggest that memorization is unlikely to be an issue.
-Such substantial performance gains would not occur if the models had already memorized the test sets.
-Furthermore, Cao et al.
-(cao2024concerned, ) highlight that code LMs do not consistently exhibit reduced performance on data extending beyond their training cut-off dates.
-This careful management of data, along with the observed performance improvements after fine-tuning, ensures that data leakage was not a major threat, thereby maintaining the internal validity of our findings.
-7. Related Work
-7.1. Code Smell Detection
-Existing approaches to code smell detection predominantly rely on heuristic-based methods, which utilize predefined rules to identify problematic code patterns (pmd, ; designitejava, ; campbell2013sonarqube, ; fokaefs2007jdeodorant, ; vidal2015jspirit, ; moha2010decor, ).
-These methods aim to flag potential smells by analyzing the source code against established standards.
-For instance, PMD (pmd, ), a widely used static code analysis tool, supports 16 programming languages and applies rule sets to detect code smells.
-These rule sets can be tailored to suit specific project requirements or design standards, making the tool adaptable for different contexts.
-Similarly, Sharma (designitejava, ) introduced DesigniteJava, an advanced tool capable of identifying code smells at multiple levels — ranging from architecture and design to implementation.
-Traditional heuristic-based code smell detectors face inherent challenges, such as difficulty in defining effective thresholds for detection.
-In response, Machine Learning (ML) and Deep Learning (DL) techniques have gained traction as promising alternatives.
-Maiga et al.
-(maiga2012support, ) implemented SVMDetect based on an ML technique — support vector machines.
-Fontana et al.
-(fontana2016comparing, ) evaluated 16 ML algorithms across four types of code smells using 74 software systems, reporting that J48 and Random Forest delivered the highest performance.
-Liu et al.
-(liu2021deep, ) utilized deep neural networks and advanced DL techniques to perform useful feature selection from source code and map these features to labels (smelly or not).
-Sharma et al.
-(sharma2021code, ) trained Convolution Neural Network (CNN), Recurrent Neural Network (RNN), and autoencoder DL models in different configurations to detect code smells.
-Their work confirms the feasibility of applying transfer-learning for code smell detection.
-Ho et al.
-(ho2023fusion, ) developed DeepSmells based on a CNN and long short-term memory (LSTM) networks for detecting code smells.
-They conducted an empirical study to compare DeepSmells with other well-established detection tools.
-Their results demonstrate that DeepSmells outperforms the baselines, achieving superior performance.
-They also anticipated that leveraging CodeBERT would enhance the effectiveness of code smell detection.
-We conduct an extensive analysis of the Machine Learning Code Quality (MLCQ) dataset, focusing on how these LLMs perform when prompted to identify and classify code smells.
-Recently, researchers have started to explore the use of LLMs for code smell detection.
-Mesbah et al.
-(mesbah2025leveraging, ) used the MLCQ dataset to analyze how GPT-4 and LLaMA perform when prompted to identify and classify code smells.
-Silva et al.
-(silva2024detecting, ) employed two types of prompts to query ChatGPT for identifying Blob, Data Class, Feature Envy, and Long Method code smells.
-Sadik et al.
-(sadik2025benchmarking, ) created a dataset for evaluating LLMs on automated code smell detection.
-They benchmarked two state-of-the-art LLMs and found that LLMs are not yet ready to replace deterministic tools in all aspects of code quality assurance.
-Wu et al.
-(wu2024ismell, ) trained a Mixture of Experts (MoE) model, leveraging specialized expert toolsets for code smell detection.
-They subsequently utilized the detection results to prompt LLMs for refactoring.
-Most prevailing ML-based or DL-based approaches for code smell detection train models from scratch.
-On the other hand, existing LLM-based approaches typically rely on simple prompting strategies without any form of fine-tuning or adaptation.
-Compared to their works, we harness the power of LLMs using PEFT methods to detect four types of widely studied code smells (Complex Conditional, Complex Method, Feature Envy, and Data Class), aiming to enhance the models’ ability to capture the nuanced relationships between different code smell types.
-7.2. PEFT Methods on Software Engineering Tasks
-Several studies have evaluated PEFT methods on various Software Engineering (SE) tasks.
-Wang et al.
-(wang2023one, ) fine-tuned UniXcoder and CodeT5 by inserting the parameter-efficient structure adapter for code search and summarization.
-According to their results, adapter tuning has shown effectiveness in cross-lingual and low-resource scenarios, surpassing full fine-tuning and effectively mitigating catastrophic forgetting.
-Liu et al.
-(liu2023empirical, ) carried out experiments by fine-tuning code models with adapter, prefix tuning, LoRA, and MHM (he2022towards, ) on both the code understanding task and code generation task.
-Their findings indicate that adapter tuning and LoRA can achieve promising performance on these two SE tasks with reduced training costs and memory requirements.
-Liu et al.
-(liu2024delving, ) examined whether adapter tuning and LoRA can perform well on code-change-related tasks, specially just-in-time defect prediction and commit message generation.
-Their study demonstrates that PEFT methods can serve as a powerful alternative to full fine-tuning for dynamic code comprehension.
-Weyssow et al.
-(weyssow2023exploring, ) compared PEFT methods, including prompt tuning and others, with In-Context Learning (ICL) in terms of code generation.
-Their investigation reveals the superiority of PEFT over ICL and highlights the potential opportunities of tuning LLMs in diverse SE scenarios.
-Li et al.
-(li2024comprehensive, ) created an instruction dataset of Automated Program Repair (APR) using prompt engineering.
-They employed this dataset to fine-tune four LLMs using four different PEFT methods.
-The results show that $(IA)^{3}$ achieves the highest fixing ability compared to prompt tuning, prefix tuning, and LoRA.
-These studies focus on evaluating PEFT methods across a range of SE tasks, primarily at the method level.
-Compared to their works, our work specifically targets the task of code smell detection, which, to the best of our knowledge, has not been studied before.
-Additionally, our work further explores the capacity of LLMs on class-level code understanding, thereby addressing a significant gap in existing research.
-8. Conclusions and Future Work
-In this work, we investigated the effectiveness of various PEFT methods on different LMs for code smell detection.
-We also evaluated how the performance of these PEFT methods varies under different conditions, such as varying hyper-parameter settings and the availability of training data.
-Our analysis reveals that PEFT methods are more efficient than full fine-tuning and In-Context Learning (ICL) in most cases, underscoring the potential of PEFT methods to enhance code smell detection and offer valuable insights for optimizing LMs for other Software Engineering (SE) tasks.
-Additionally, we have constructed high-quality datasets for four widely studied code smells, i.e., Complex Conditional, Complex Method, Feature Envy, and Data Class, and we provided the replication package of this work online (replicationPackage, ) to facilitate future research and ensure the reproducibility of our study results.
-In the next step, we plan to further explore how PEFT performs in cross-language scenarios for code smell detection.
-Moreover, we plan to investigate the application of PEFT methods for fine-tuning LMs to assist in a wider range of SE tasks, such as bug detection, code completion, and code smell refactoring.
-These directions will help us better understand the generalizability and practical value of PEFT methods in real-world software development settings.
-Data Availability
-The replication package for this work has been made available at https://github.com/MabelQi/PEFT4CSD.
-Acknowledgements.
-This work has been partially supported by the National Natural Science Foundation of China (NSFC) with Grant No.
-62172311.
-The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of Wuhan University.
+A.2. Random Seed
+Table 12 presents the results of GraphCodeBERT with different PEFT methods (i.e., prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$), as well as full fine-tuning, for detecting the Complex Conditional (CC) code smell across three different random seeds (0, 42, and 999).
+As shown in Table 12, the variations in performance metrics across different random seeds are minor.
+These small differences indicate that our core findings are robust to the effects of random initialization, with overall trends remaining consistent.
+While some fluctuations in performance are observed, they do not materially alter the conclusions drawn from the experiments.
+Nevertheless, we acknowledge that conducting multiple runs with different seeds would yield more statistically reliable results.
+Appendix B The Impact of Context Window Limit to SLMs
+In the process of data construction, considering our GPU memory limitations, we filtered out any samples that exceed 1,024 tokens.
+However, several SLMs used in our study, including CodeBERT, GraphCodeBERT, and CodeT5, can only accept input sequences up to 512 tokens (see Section 4.3).
+Consequently, any code sample longer than 512 tokens must be truncated before being fed into these models, which may lead to the loss of semantically important information and potentially affect the accuracy of code smell detection.
+To quantify this effect, Figure 10 presents the token length distributions for CodeBERT and CodeT5 across different code smell categories.
+GraphCodeBERT shares the same tokenizer as CodeBERT and therefore exhibits the same distribution, so only the results for CodeBERT are shown.
+The figure indicates that the majority of samples contain fewer than 512 tokens for all smell types.
+Although some samples slightly exceed the 512-token boundary, almost all remain below 600 tokens.
+This suggests that truncation beyond 512 tokens affects only a small portion of the dataset.
+Figure 10. Token length distribution for CodeBERT and CodeT5 across CC, CM, FE, and DC
+To further evaluate whether the truncated segments contain information that is critical for code smell detection, we conducted an additional manual analysis using CodeBERT as a representative model.
+Before the main review, we performed a pilot annotation: for each code smell type, we randomly selected 10 positive and 10 negative samples, which were independently labeled by the first and fourth authors.
+Across all smell types, the two reviewers agreed on 85-95% of cases, indicating high consistency and supporting the reliability of our labeling.
+Following this pilot, we proceeded with the main manual review.
+For each code smell type, we randomly sampled 50 positive and 50 negative instances whose lengths exceed the 512-token limit.
+These samples were manually reviewed to determine whether the truncated portion included code elements that was essential for correct classification.
+Table 13 summarizes the results.
+Table 13. Number of truncated samples in which the truncated portion does not affect the detection of CC, CM, FE, and DC
+$C$ CM FE DC
+#Positives #Negatives #Positives #Negatives #Positives #Negatives #Positives #Negatives
+49 49 50 47 50 49 50 48
+The results show that for nearly all sampled instances, truncation beyond 512 tokens does not remove information critical for code smell detection.
+Only a few cases (1–3 samples per category) potentially lose relevant information.
+Most essential indicators appear within the first 512 tokens, while the truncated segments typically contain auxiliary logic or repeated structures that do not affect classification.
+Overall, these findings, combined with the token length statistics in Figure 10, suggest that context window limitations of CodeBERT, GraphCodeBERT, and CodeT5 have minimal impact on evaluation results.
+Nonetheless, we acknowledge that this review was conducted on a limited subset; for exceptionally long or complex methods, truncation could still obscure semantically important content.
+Future work could extend this analysis to a larger portion of the dataset to further quantify any residual effects more comprehensively.